A Complete Bibliography of Publications in Distributed Computing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 August 2024 Version 1.09

Title word cross-reference

(1+ ∈, β) [EZ06]. (n, k) [BRS18]. 2^{m} [HL99]. ^{2} [GGT18]. Δ [GHM21, GHS17]. ℓ [Tau14]. F [SA95]. k [AGT10, AM97, CMQR23, GKS08, HKR19, HP05, PSS08]. log n [AE18]. m [MT94]. n − k + 1 [BRS18]. o(m) [MK21]. Ω [Zie10]. π [LV10]. r [DT01]. t < n/8 [ZIP95]. Θ(n^{2}) [CDG^{+}24].

-calculus [LV10]. -channels [SA95]. -clock [HL99]. -coloring [GHM21]. -exclusion [AM97, Tau14]. -majority [CMQR23]. -operators [DT01]. -register [MT94]. -service [HKR19]. -set [BRS18, HP05]. -shot [GKS08]. -simultaneous [AGT10]. -spanners [EZ06].

07 [Pel08].

2-Choices [CNS21]. 2-edge-connectivity [CHD20]. 2003 [Fic05]. 2005 [Fra07]. 2007 [Wat09]. 2008 [PS10b, Tau10]. 2009 [Alv11, Kei11]. 2010 [Gue12, LS13]. 2011 [Fra13, Pel13]. 2012 [Agu14, Pan15]. 2013 [Tau15]. 2014 [Att17]. 2015 [Att19, Mos18]. 2016 [Att19]. 2017 [Att20]. 2018 [Att21a]. 2019 [Att21b, Att23]. 2020 [Att22a, Att22b]. 2021 [Att24]. 20th [DS08].

Ability [FG01]. abstract [BKT97, Jon94, KLN11]. abstraction [Gra99]. abstractions [Ksh04]. Access [Baz01, BKK16, CKR09, EFT16]. accessed [AMW05]. achieve [CCW95]. achievement [CT96]. Achieving [BR91a, BNP97, WS09]. Acknowledgement [Ano21]. Action [KS03, HF89, MDH86]. actions [KS92b]. Activating [Pel07]. Active [CM05]. actively [MSS22]. ad [AMZ12, CGR02, DGW05, GNZ17, HWY19, KP05, Pel07]. ad-hoc [AMZ12]. adapting [ABS16]. Adaptive [AT10, AB02, CS94, KA07, AST02, AD07, AADM23, AFG02, AKW06, BEW11, CCW95, DH07, KA12, RSZ13]. Adding [DGK20]. addition [ALK15]. additive [CHHM15, CHKY18, EGN23, Par17]. addressing [GGH05]. adjacent [CM10]. adjusting [AKT14, ZB92]. advantage [BT18]. Adversarial [PST15, CKR09, HST12, ORZ18, SV19]. adversaries [HR13, JYZ20, KKM19, LL15]. adversary [Asp15, AB05, DGFT11a]. advice [DGFK15, FGP09]. against [KKM19, LL15, SV19]. agent [KKS17]. agents [Bro87, DP14, DFS06, FOV08, KK15]. aggregate [ZYC16]. aggregation [ABM17, Yu11]. Agreement [HKK08, PCR14, AWH92, AHR95, BRS18, CJM17, CHT20, Coa93, Cri91, DGFTT13, Fek90, FHS22, GP21, Had87, HP05, LL15, MHG15, MRT08, PT92, SAAAA95, ZIP95, ZB92, Zie10, ACD^{+}23]. agreement-oriented [MRT08]. Algebraic [Bro92, CHKS19, HH87, ABL13]. algorithm [AT12, AFG02, BE10, BGR01, BBP19, CLM17, CSC11, CSZ10, Ciu94, DGHP23, FM04, FZ15, FO23, Ger99, GK93, GGP97, GHV01, HH87, Hua06, IRW13, JRS02, KC99, KGM99, KP08a, MRZ11, NdM16, PT14, Pix88, PSL00, Rhe98, SS17, SAAAA95, SW10, Waw15, WL93, YA95, ZB92]. Algorithms [Mat87, AFS11, AD07, ACR19, AP91, AT96, AM97, ACY21, BC95, BNNN90, BND91, BE14, BBEiKS23, BG87, BCD13, BEW11, CHFV19, CHPS20, CM86b, CBM21, CSZ10, CPS17, CDN18, DRS23, EFO22, EKS16, EZ06, Fek90, GM87, HP15, JSX16, KS92a, KKS10, KKT12, KSV13, KY11, Lam90, LW05, MVP07, MS14, PR01, PS10a, Pet10, Sin91, ST87, TS11, VJ94]. allocation [Jou10, Rhe98, YEFH03]. allow [FGM01]. Allowing [BKT15]. Almost [ALH^{+}23, BBS21]. alternative [KKS17]. alternator [GH07]. Alternators [HHH03]. altruists [KFH12]. amnesiac [HT23]. amoebot [DRS23]. among [BKT97]. amortized [BHMT23]. Analysis [FHS22, Pac91, BR91b, BKM96, CP88, DFN18, HR05, HP10, LM91]. analyzing [LM95]. AND-requests [TL95]. anniversary [DS08]. Announcement [Att12]. Anonymous [BR13, BRS18, GR07, AFR06, CJM17, CGM12, DP14, DFY20b, DN17, FHK17, FPP01, Gel18, GMP23, KS01, Pel07, Waw15]. Anti [Zie10]. Anti- [Zie10]. ANTS [FK17b]. AODV [vGHT16]. application [DN17, HR94]. applications [AST02, AW09, Asp12, AFG02, BEF16, DFV22, KP08b, MG05]. applied [LV10]. Appraising [AFK88, FM03]. approach [ABS16, ABF88, Asp12, BBP19, CH01, CDN19a, DNN17, DP15, Ebe91, GS98, JPZ99, Jos88, LMS08, LS92]. Approximate [PSS08, AAE08b, BFM21, CHDL21, Fek90, HRZ20, TLV24]. Approximating [FOV08]. approximation [CHD20, GS14, HKR19, KP08a, KKT12, KW05, Waw15]. approximations [LPW13]. arbiters [Bla86]. Arbitrary [CDN19a, AA95, BGJ07, CGM12, CKP12, DJV00, DP14, DFS06, GVB21, TLV24]. Arbitration [Lam03]. Arbitration-free [Lam03]. architecture [HMN14]. architectures [CLM17, HU08]. arrays [AP91, Lis91, Myo93, Raj89]. arrivals [ACY21]. Arrowhead [Lam92]. aspect [KS03]. aspect-oriented [KS03]. aspects [TT19]. Aspnes [PSL00]. assertional [Hes94]. asset [KPPT23]. Assigning [FPP01]. assignment [HKR19]. assuming [DIM93]. assumption [DFP04]. assumptions [ADGT08, DNN17, KSV13]. assurance [DS03]. Asymmetric [ACTZ24]. Asymptotically [Fek90]. Asynchronous [BHK^{+}24, BBP19, Bro90, Jou00, KT22, PCR14, ADS22, AJM^{+}23, Asp03, AB05, Baz01, BR13, BMIY08, CHGH19, CBMT96, CDN19b, DP14, GMT18, GHV01, Gue02, HR99, Jon94, KT12, KPPT23, Lam06b, LL15, Liv93, LM95, MK21, MHN18, PT92, SAAAA95, Sch97a, Sch97b, SB95, XTB08, CDN19a]. asynchrony [KKN^{+}24]. at-most-once [CCW95]. ATM [FNP01]. Atomic [AHR95, MT94, ADL90, AHS21, BRS18, CLM17, DGW05, GLS10, Gue02, HHT22, Zie08]. atomicity [DIM93, Ksh98, Lam90]. attack [HST12]. attacks [SV19]. authenticated [BM05, ST87]. authentication [Gon95]. automata [Mic18, San12]. automated [ALK15, BBS12]. Automatic [MM00]. Automation [LKJ19]. autonomous [KKN^{+}24, Koz12, PPV15]. availability [YG09]. available [Yu11]. averaged [BGKO23]. award [Att12]. aware [AMS20]. axiomatic [Mel91].

backoff [ACY21]. backup [BO22, HKR18]. balancing [AFS11, ABS16, GVM23, HT06, KPT11, LW16, RB06]. balls [BCP19]. balls-into-bins [BCP19]. bandwidth [CHPS20]. bare [CKV22]. bare-bones [CKV22]. barrier [AE18, EKS16]. based [ADGT12, ADL90, AT96, BMP06, BBS12, CMN02, EGZ20, FHV97, FPS05, GKM07, GHV01, Hal00, HM04, HP10, HMR00, HR99, Jos88, LDB05, MRR04, MRR06, SB12, Sin99, TK06, Vog91, YEFH03]. batched [ACY21]. Bayesian [SV19]. BDDs [EFT93]. be [FSY20, KS92a, MNV11, PST13]. Beeping [AAK13]. before [FHK17]. behavioral [CC95]. Behavioural [NRT90]. being [BT18]. belated [Dij86]. Ben [AT12]. Ben-Or [AT12]. benign [CBS09]. Berthold [Att16]. best [Att12, FM11]. between [AFR06, CM10, DG10, LLR17]. beyond [CHT20]. BG [BGR01]. BGP [FPS05]. BGP-based [FPS05]. bias [VBA22]. biased [CMQR23]. bidirectional [vLT87]. binary [AWH92]. bins [BCP19]. bipartite [GS14, HT09]. bit [BNNN90, BND91, MRZ11]. bits [AFP20, BKN19]. black [DFS06]. blocking [Gue02, Moi01, Sha19]. board [KPT11]. bones [CKV22]. boost [GKK08]. boosters [GK08]. Booting [WS07]. born [Gou86]. both [CP88]. bound [ACHL20, AK02, BUW20, CR10, CDG^{+}24, FZ15, HKK08, KA12]. Bounded [CHHM15, HK09, IL93, AMS20, BE13, BG93, FNP01, Hes99, KR90, Kes95, KKN^{+}24, NdM16, SW10]. Bounded-Contention [CHHM15]. bounded-independence [SW10]. Bounded-wait [HK09]. Bounding [BCD13, VK20]. Bounds [Haj91, AMW05, BSD88, BV03, BD97, CR10, CM22, CEL23, EK11, FK17a, GVB21, Gon95, GHS17, HM16, HP05, JYZ20, KW21, Lam06b, LW16, PST16, Tau14]. branching [BK98]. Breaking [AE18]. Breathe [FHK17]. bridge [KC99]. Broadcast [MK21, AFRT24, AFR06, BKM96, BD97, CKL06, Ciu94, GP21, GHK15, GM87, MMSA93, PPS17, PST16, Zie08, PS02]. Broadcasting [EGS09, FP10, KP05, KP07, BEF16, CGR02, EKP16, FGM01, GGL13, GKS08, HS91, JLC07]. broadcasts [BSD88, ST87]. broker [CB01]. buffered [KKS10]. buffers [KR90]. Building [JLC07]. bulletin [KPT11]. Byzantine [ACM06, ACD^{+}23, AFRT24, AMW05, AWH92, Baz00, Baz01, BDD16, BDL22, BCG22, BCG24, CDG^{+}24, CNV05, DPBRP20, DGFTT13, EMGR22, FHS22, GP21, GM21, Had87, HKK08, HT24, KT22, MR98, MMT03, MHG15, NK24, OBG08, PCR14, SAAAA95, TLV24, WBB12, ZIP95, ZB92]. Byzantine-resilient [CNV05].

C [GGT18]. Cache [Bri99, Gra99]. caching [Ger99, JPZ99, LLR99]. Calculational [Kes95]. calculi [Gor10]. calculus [LV10]. Calendar [Ano86]. can [FSY20, KS92a, MNV11]. cannot [PST13]. canonical [DRS23]. capacity [IFR20]. capture [MR03]. capture-resilient [MR03]. CaRuD [HG01]. case [AH95, AFRT24, BR13, GS91, KSW10, MDH86, PSL00]. Causal [ANH95, ABM02, BMP06, CT96, IRW13, KS98, SM94]. Causality [ABM02, Ksh98, BKT97]. causally [CBMT96]. Causing [EM09]. CB [AKT14]. CCS [EFT93]. centralized [BKS89]. chain [vGGT23]. chains [WWA04]. Change [Ano00, Att09]. Changes [Sch95]. Changing [Sch90]. channel [BKK16, FGN19, KKM19]. channels [CKR09, EM09, HWY19, SA95]. Characterization [Gra99]. Cheating [MDH86]. checkability [FRT13]. checkable [BBC^{+}23]. Checking [HU08, AKV18, BK98, EFT93, GHS10]. checkpoints [HMR00]. Chemical [TT19, CCS17, SHD21]. chip [DS86, FS12]. chirality [CDN19b]. Choice [BNBOD89]. Choices [CNS21]. Choreographies [CMS18]. churn [KSW10]. circle [FPV17, Koz12]. circuits [ACM86, Ebe91, Mar86, Udd86]. circulation [DJV00, HC93, HW97]. citizen [Caz03]. class [BC95, Caz03, MM00, MS94, PW97, Sin91]. classes [MRT08, NS22, NdM16]. classification [EGZ20]. classifying [Udd86]. clique [CHKS19, CHDL21]. cliques [BPS11, CK20, EFO22]. Clock [HS91, Cri89, DB93, FL06, HL99, Ksh04, WS07]. clocks [DL12, IRW13, Lis93, Sto00, TRA99, VK20, WS09]. Close [RT16, PPV15]. Closed [LM95]. Closing [DG10]. closure [EM09, Lis91]. cloud [AGPS24]. clustered [GT91]. clustering [EKR11]. coded [CLM17]. codes [MS16, Ver88]. Coding [CHHM15, JLC07]. coin [BGK14, CP88]. collaborative [BUW20, KG22]. collect [AD07, AFG02, AKW06]. Collecting [ML97]. collection [Jon92, Pix88]. colliding [CKP15]. collision [BYGI91, CDN08, GHK15, GHV01]. collusion [KFH12]. coloration [HT09]. Coloring [MW08, BE13, BE14, CPS17, FGP09, GHU20, GHM21, GK93]. colouring [JSX16]. Combinatorial [BE14, CR10, GMT18]. combining [HK09]. Comments [KK99]. commerce [KGM99]. commit [CG21, CW90, Gue02]. commodity [AK09]. Common [NT93, MM08, PT92]. Common2 [AGM07]. communicate [BKT15]. communicating [BR91b, Bro87, Jos88, LLR17, LM91, Mar86]. Communication [ACD^{+}23, AAS18, HMR00, HRZ20, AB93, ABH19, BSD88, BM05, BKN19, BHS18, CBMT96, CKV22, CMQR23, EM09, FHK17, GSM07, GPX07, HP01, Lam86a, Lam86b, Lin88, MS89, MDH86, PST16, Sch06, WZ17, ZYC16]. Communication-based [HMR00]. Communication-efficient [AAS18]. communication-time [ZYC16]. communications [MHN18]. Compact [BT18, GP03, KP08b, RGB13, Kor07, KKM15]. comparison [GHW12]. compatible [FRS11]. Competitive [RSZ13, KPY18, ORZ18]. compilation [LBH91]. Compiling [ACM86, Mar86]. complete [CC88, KGM99, Mel91, SAAAA95]. completely [Fin94]. complexities [BGKO23]. complexity [ACD^{+}23, ALK15, AH95, AWH92, AK02, BHMT23, CJM17, CPR20, CNV05, CEL23, DG10, EFR08, EGZ20, EJWK21, FGM01, Gel18, GRS04, GGK11, GS91, HW11, HRZ20, KA12, KPS10, LLR17, LL15, MRZ11, MOY12]. complexity-based [EGZ20]. Component [CDPP24, BBS12]. component-based [BBS12]. components [KS01]. Composable [SHD21]. Composite [And93, And94]. composition [DH07, EM09]. compositional [JPZ99, PJ91]. computability [GMT18, Hav00, MRT08, ST16]. Computable [Hav00]. Computation [AAP06, ADS22, AH95, AAE08a, ABS18, Bei07, CCS17, CFH^{+}22, CDPP24, DKZ17, GVB21, GS91, HP10, KKM15, LPSP19, MSS22, MG04, MG05, PP24, SS17, SDP15, SHD21, WZ17, ZYC16]. computational [AAR07, Tau20]. computations [ACHL20, ADL90, BKT97, CHGH19, CHCGS23a, CHCGS23b, HMR00, Ksh98, NS22, SM94]. Compute [DDN14, ACR19]. computer [Jou02]. Computing [BFM21, HK18, Lis91, AGPS24, CBS09, CMN02, Cla86, FR03, FM03, FPV17, FGP09, FRT13, FS12, GR07, HJV15, KS01, MS14, Sch95, SA95]. concatenation [ABBM02]. concerning [KSV13]. Concise [MT07]. Concurrency [San12, AGM07, ABS87, BG87, DRS23, DGK20, MdV89].

Concurrent [ABH16, JT21, LMP06, PT92, Vid90, AKT14, AG07, AADM23, Bro87, Gol91, HG01, Lin88]. Condition [MRR04, MRR06, SS01]. Condition-based [MRR04, MRR06]. conditional [FHS06, NS22]. conditions [KS98, MMSA93, MRR04, Tau20]. Confidential [GGK20]. configurable [CBP02]. congenial [Jou02]. CONGEST [CM22, CK20, LMR21, MK21]. congested [CHKS19, CHDL21]. Congestion [HIZ21, ACHL20, ABH16, ALH^{+}23, BFM21, FM11, FNP01, KKI21]. congruence [Vog91]. connected [ABH19]. connection [AR97]. connections [HJV15]. Connectivity [Had87, CHD20]. Consensus [BND91, CDN08, WBB12, WSS19, AJM^{+}23, AGT10, ACT00, AT12, AAS18, Asp03, Asp12, Asp15, AB05, BNP97, BCT17, BG93, BCG22, BPV06, BN94, CGM22, CBM21, CDG^{+}24, CNV05, DGK07, EHT07, FLM86, FN19, GK11, Gel18, GM21, GKS22b, GKS22a, HHT22, HR99, Lam06b, MM08, MNV11, MMSA93, MRR04, MRR06, PSL00, Sch97a, Sch97b, TLV24, TS11, LH00]. consistency [BOEY03, Bri99, CHK14, CT96, DGS19, Fri95, Ger99, JPZ99, KG22, Liv93, LD99]. consistent [AB19, Gra99, MMT08, NT93]. Constant [KW05, Mit97, BOEY03, EN22, KT12, LPSP06, TRA99]. constant-hopbound [EN22]. Constant-time [KW05]. Constraint [Sin99]. Constraint-based [Sin99]. constructing [EZ06, GM87, HK09, JRS02, Lin88]. construction [CHKY18, DJ09, EN18, GHSW23, HR94, MS16, Mic18, Mil87]. constructions [EFT16, FK20, Moi01]. containing [GGP07, Hua06, KP86, KT12]. content [RB06, TK06]. content-based [TK06]. Contention [BMIY08, CHHM15, FGN19, AB05, SB12, Zie08]. Contention-free [BMIY08]. continual [CFH^{+}22]. Continuous [MM08, ABS16, GGK11]. contracts [DGK20]. control [ABS87, BKS89, BG87, DRS23, Raj89]. controllability [Hie12]. controlled [ML97, MG04]. controller [EK11]. Convergence [AAG93, BBM01, KKN^{+}24]. cooks [MS18]. cooperation [MRS06]. coordination [BNBOD89, GVM23, HR05, HS08, NT93]. copies [AGE98]. CORBA [CB01]. core [CNS21]. cores [GGT18]. Correction [CHCGS23a, GKS22b]. correctness [AT12]. correlation [YG09]. cost [Baz01, BFM21, Ciu94, DB93, FPS05, HS08, INS09, MP16]. coteries [INK95]. count [FNP01]. counter [Kes95]. counterfactuals [HM04]. counters [AB19, BHMT23]. Counting [AA95, DMS02, GKM07, HSW96, LR19, MMT08]. Coupling [FMlP06]. cover [GS14]. coverage [EKS16]. covering [BBEiKS23, KY11]. crash [ACT00, AWH92, DPBRP20, VJ94]. crash-recovery [ACT00]. crashes [GRS04, KFH12]. Create [HG01]. Critique [Lam92]. Cross [vGGT23]. Cross-chain [vGGT23]. crossbar [KKS10]. Crumbling [PW97]. cryptocurrency [GKS22b, GKS22a]. cryptographic [Gol03, MR03]. Cryptography [Gol03]. CSP [LD99, Pel91]. cuts [PP24]. cycles [EFO22]. cyclic [ML97]. data [EKR11, GP03, HST12, LM91, ST16, XTB08]. database [Bro92]. databases [BG87, BR91a, KS92b]. dataflow [Jon94, Lam08]. deablock [Ruk91]. deadline [GGK11]. Deadlock [BR91b, HL89, BT87, Sin91]. Deadlock-free [HL89, Sin91]. deadlocks [TL95]. decades [FM03]. Decentralization [BKS89]. Decidability [Fin94]. Deciding [AFP20]. decision [CH01, FGP14]. decision-theoretic [CH01]. decomposition [BE10, NdM16]. decompositions [KP08b]. decoupled [SCA15]. decreasing [BFM21]. dedicated [DS08]. defective [CHCGS23a, CHCGS23b]. Defending [SV19]. defense [KPY18]. defining [PT92, Udd86]. definitions [ANH95]. deflection [Haj91]. degradation [LKJ19]. degrading [AT10]. degree [AMS20, BG93, GHU20]. degrees [FG14]. Delay [Ver88, AHS21, Bla86, CDP16, Ebe91, Mar86, Udd86]. Delay-insensitive [Ver88, Bla86, Ebe91, Mar86, Udd86]. delayed [KKM19]. delays [FM11]. Delegation [MR03]. Deletion [HG01]. delivery [CCW95, CH01]. Demand [AMS20, BC95]. Demand-aware [AMS20]. dependency [AG07]. dependent [OR96, SL87]. deployment [EKP16]. depth [DJV00, HC93, NdM16]. depth-first [DJV00, HC93]. deque [HLS06]. Derandomizing [CHPS20, CIK11]. derivation [Kes95]. derive [ST87]. Deriving [KHvB96, KK99]. Design [BKM96, Ciu94, Pel91, Bri99, Cla86, CBP02, FSS06, GND07]. Designing [KR90, Ebe91]. designs [AMS20]. despite [FHK17, RSZ13]. Detecting [CK20, SM94, Sto00, EFO22]. Detection [CG98, AGE98, ACT00, BYGI91, BT87, GLS12, GHK15, HR94, KKM15, LMM07, MS94, Mat87, MVP07, Ruk91]. detector [DGS19, EHT07, GHN09, HR99, MRT08, Sch97a, Sch97b, Zie10]. detectors [AKN12, BDT07, BR13, CDN08, Gue02, GK08, GKK08, HR05, HP05, PSW12]. Determining [MNS17]. Deterministic [CGR02, DP15, KP07, AFRT24, BE13, BT18, BBP19, Fis20, GMP23, KKS17, MP16, PST16]. devices [MR03]. DEX [PRT16]. diagnosis [Coa93]. diagnostic [CC95]. diameter [CPR20, LPSP06]. diameter-two [CPR20]. differential [GND07]. digital [KP86]. digitized [AP91]. Dijkstra [CSZ10]. dimensional [BAES95]. dining [DFP04]. directed [FO23, TLV24]. direction [BV03, BDK15, Sin97]. disagreement [DGFT11a]. DISC [Agu14, Att17, Att20, Att21a, Att22a, Att23, Att24, DS08, Fic05, Fra07, Kei11, LS13, Mos18, Pel08, Pel13, Tau10]. disconnected [MRS06]. discovery [BPS11, PST15]. discrepancies [AGE98]. discrete [ABS16, SHD21]. disjoint [EFT16, JT21]. disjoint-access [EFT16]. Disk [ACM06, CM05, GL03]. dissemination [AMZ12, CEW13, FHK17]. dissipation [Kes95]. Distance [Sli07, LPSP19, SDP15]. distinct [BK07, KK15]. Distributed [AFS11, AGPS24, ABS87, BE13, BO22, BDK15, BT87, BPS11, CHKY18, CHCGS23b, CC88, CKV22, CFH^{+}22, CPS17, Cla86, DJ09, DDS20, EKS16, EKR11, FPV17, FGP09, GM87, HKR18, HKR19, HS07, HT09, KK07, KY11, KS01, LPW13, LPSP19, LPSP06, MRS06, MSS22, OR96, Pet10, SB14, ŢH09, WWA04, ACHL20, AJM^{+}23, ADL90, ACTZ24, ALH^{+}23, ABF88, ABS18, AFK88, AFR06, AK09, BG14, BC95, BGM86, BE10, BE14, BKT97, BKT15, BBEiKS23, BBS12, BKA12, BGR01, BCD13, BR91a, BN94, BHS18, BHS22, CP88, CHFV19, CHGH19, CHPS20, CHD20, CBS09, CSC11, CG21, CP89, Cri91, EFO22, EMGR22, EZ06, EN18, FM04, FFP21, FR03, FLM86, FM03, Fis20, FO23, FGP14, FS12, GSM07, GP03, GHU20, GHM21, GGP07]. distributed [GM89, GT91, Gue02, GND07, HM16, HF89, HR05, HST12, HH87, HP15, HR94, HMR00, HJV15, HR13, HU08, HMN12, Hie12, HMN14, INS09, JSX16, JRS02, KP92, KR90, KGM99, KP08a, KKT12, KPS10, Ksh98, KW05, Lam90, LW86, Lis93, LMM07, ML97, MM00, Mat87, MRZ11, MS16, Mic18, MG04, MS14, MS97, NdM16, NYD21, PR01, PS10a, PJ91, PP24, PSS08, SDP15, SS94, Sch95, SM94, SA95, SL87, Sin91, Sto00, TT19, TRA99, WZ17, vLT87, CHCGS23a]. distribution [AADM23, Gor10, ABM17]. distribution-adaptive [AADM23]. Do [GRS04]. Do-All [GRS04]. does [FK00]. Dolev [CDG^{+}24]. dominating [JRS02, KW05, LPW13, Waw15]. Doomsday [LMM07]. drinking [WL93]. dual [PS10a]. Dynamic [BHS22, Sch06, AM05, BCF11, BV02, BKM96, CST15, DDS20, DFN18, DIM93, EKR14, GGH05, GLS10, HLS06, IRW13, JYZ20, Kor07, KP08b, MSS22, MNV11, Myo93, NW05, WSS19]. dynamic-sized [HLS06]. Dynamics [HS12, ABH16, BCT17, BHK^{+}24, CNS21, CMQR23, FM11].

each [BKT15]. Early [Sch97a, Sch97b, AFRT24]. early-stopping [AFRT24]. Easy [FLM86]. Eclecticism [FGP06]. edge [BGKO23, BE13, CHD20, GHU20, HT09, Mit97]. editing [KG22]. Editorial [Ano00, Asp07, Att09, Att20, Gue06]. effect [EKP16]. effects [CDN19a]. Efficient [AGE98, AD07, AG07, AKW06, AB05, BYGI91, BM05, BEF16, CSC11, Coa93, EZ06, GGL13, GW19, Gon95, KS92a, KKT12, MS94, SDP15, Sin97, VJ94, AAS18, ABF88, AB02, DGHP23, EN18, FHK17, GKS08, GHW12, HP01, Hua06, JRS02, MS16, MRR04, PW97]. Election [CGM12, BGJ07, BT18, BKM96, CC88, CPR20, DP14, DS18, FP11, FP15, GLS12, GMP23, IRW13, MOY12, RFS07, SAAAA95, Sin97, vLT87]. elementary [NRT90]. Embedded [CDN19b]. embedding [HIZ21]. embeddings [KKT12]. empirical [SS17]. emulation [BYGI91]. encapsulate [PSW12]. energy [DGHP23]. energy-efficient [DGHP23]. ensure [EFT16]. entropy [ABS18]. envelope [LDB05]. environment [CB01]. environments [CKR09, PST15]. equational [LV10]. equations [GND07, Raj89]. equilibria [BFM21, FOV08]. Equivalence [NS22]. equivalent [GM04]. Erlang [CCT22]. error [BR91a, SCA15]. error-prone [BR91a]. establishment [BK07]. estimation [Sli07]. Euclidean [EKP16]. evacuation [Haj91]. evaluation [AAK89, SS17]. even [CDG^{+}24, EFO22, FGP06]. events [AG07, Ano86, BKT97, GSM07]. eventual [DGS19, KG22]. eventually [AB19]. ever [GHN09]. evolution [AO86]. evolving [DL12]. examples [ST16]. exclusion [AM97, AK01, AK02, AKH03, AB02, BGM86, BKK16, CSZ10, CS94, DG10, GR19, HW11, Her00, Jou00, KA06, KA07, KA12, Sin91, Tau14, YA95, vdS87]. exclusive [DNN17, TL95]. execution [MG04, ŢH09]. Exhaustive [CP88]. existence [Bla86]. expanders [PT14, PRT16]. expansion [AW09, NdM16]. Expected [NK24]. expensive [WZ17]. explicit [ALK15]. exploration [BUW20, CIK11, DDS20]. exponential [BT18]. expressions [ACM86]. extended [AGM07, YEFH03]. Extending [IFR20, PMCP22]. extension [GP21]. extensions [KP93].

facility [HP15]. fading [FGN19]. Failure [ACT00, GK08, PSW12, AKN12, BNBOD89, BR13, DGS19, DGK07, EHT07, Gue02, GKK08, GHN09, HR05, HP05, HR99, MRT08, Sch97a, Sch97b, Zie10]. Failures [Vog91, AH95, AWH92, DGFT11b, GS91, Had87, HT24, KT22, MNV11, OBG08, WS07]. Fair [LM16, vdS87, Bla86, CIK11, Mic18]. Fairness [AFG93, Lam00, AFK88, BK98, DFP04, Hes99, PSW12]. families [LPW13]. family [AKN12, FK20, MVP07]. fan [AA95]. fan-out [AA95]. Fast [AHM13, AT96, AAE08a, BG93, BN94, CHFV19, CHD20, CHDL21, CMS13, CEW13, KKM15, Lam06a, PS10a, ABBM02, AK01, AAE08b, BKN19, HR99, KP08a, YA95]. fast-path [AK01]. Faster [Asp15, GPX07]. Fault [ADL90, ABM17, BG14, DFV20, GGP07, Jal89, KT12, ADS22, ALK15, BDT07, BKA12, Coa93, DDM16, DH07, FS12, GGK11, GR07, Hua06, KKM15, LKJ19, MS14, Par17, Pel91, SS01, ST87, ŢH09, ZYC16]. Fault-containing [GGP07, KT12, Hua06]. fault-gap [KT12]. fault-tolerance [ALK15]. Fault-tolerant [ADL90, ABM17, DFV20, BKA12, DDM16, FS12, GGK11, GR07, LKJ19, MS14, SS01, ST87, ZYC16]. faults [CDP16, CBS09, CCS17, DPBRP20, MHS14]. faulty [CKO19, KS01, KW21, SAAAA95, WBB12]. FCFS [DG10]. feasibility [CDN18]. federated [FHS22]. Feedback [JSX16]. Fence [KK15]. Fetch [CRR23]. few [AFP20]. FIFO [EM09, KKS10, LMS08]. file [AGE98]. filters [KU92]. Finding [MG04, KC99]. finite [AAP06, Bro87, Mil87]. finite-state [AAP06]. finitely [FOV08]. firing [GM04, LR19]. first [BBM01, Caz03, DJV00, HC93]. first-class [Caz03]. first-order [BBM01]. fixed [AP91, Lis91]. fixed-size [AP91]. flooding [BCF11, CMS13, HT23]. flow [GGP97, ABM17]. Flow-Updating [ABM17]. flows [AK09]. fly [Jon92]. Fooling [ACHL20]. forests [TK06]. forget [CKP12]. Forgiving [HST12]. Formal [BLT16, Ebe91, Udd86]. formation [CDN19b, DFY20a, FPV17, CDN19a]. Forming [DFY15]. foundations [TT19]. Four [GMP23]. framework [BBS12, CCT22, KG22]. free [ACR19, AM97, BNP97, BRS18, BMIY08, DMS02, FRT13, GGH05, Hav00, Hes94, Hes95, HG01, HL89, LMS08, Lam03, PPV98, PMCP22, Sin91, TV02]. freedom [DGFK15, EFT16, GKK08]. French [CPS17]. FSMs [Fei99]. Fully [BEW11, CHCGS23a, CHCGS23b, Jon94]. Fully-adaptive [BEW11]. fully-defective [CHCGS23a, CHCGS23b]. function [AAK89, HP10]. functional [Bro92]. functions [BFM21, GVB21, Lin88, ZYC16]. fused [BG14]. game [HTW07]. games [ABH16, BFM21, FM11, HHM10, HS12]. gap [DG10, KT12]. garbage [Jon92, ML97, Pix88]. Gathering [CDN18, DDN14, Koz12, BDD16, BDL22, DNN17, DPBRP20, DN17, PPV15]. General [Kor07, AH95, SB14]. Generating [CC95, EFT93, Fei99]. generation [AJM^{+}23, ABS18, HMN12]. Generic [PS02]. Genuinely [EMGR22]. Geometric [GMT18, AP91, CEW13, DFY15]. GeoQuorums [DGW05]. Getting [PPV15]. GIOP [CB01]. global [AH95, AMZ12, BBS21, CG98, HP10, KSV13, Sto00]. Good [AFRT24]. Good-case [AFRT24]. gossip [ABL13, GGK11, GGK20, MS03, MS97]. graceful [BR91a, LKJ19]. gracefully [AT10]. gracefully-degrading [AT10]. Gradient [FL06]. grail [SM94]. Graph [HST12, BGKO23, BE14, BDK15, CHFV19, CPS17, FGP09, INS09, KKI21, vdS87]. graphs [AW09, BE10, BCF11, BV03, CST15, CIK11, CKP12, DN17, GK93, LPW13, LPSP06, SW10, TLV24]. Greedy [AK09, BAES95, FG14, JSX16, SB12]. grid [BUW20]. group [Cri91, Jou00, Jou10, Sch06]. growth [CFH^{+}22]. guarantees [vGGT23]. guard [Sch90].

Handling [PS02]. handshake [AR97]. hard [ACR19]. hardness [NS22]. hardware [Cla86, IFR20]. hash [GGH05]. having [BPV06]. healer [MNV11]. healing [PT14, PRT16, SS17]. Heard [CBS09]. Heard-Of [CBS09]. heavy [BC95]. Herlihy [PSL00]. Herman [FZ15]. heterogeneous [GVM23]. Hierarchical [Ruk91, SA95]. hierarchy [MRR04, PMCP22]. high [BC95, CB01, Fri95, HK09, SA95]. high-level [Fri95]. high-performance [CB01, SA95]. high-throughput [HK09]. Highly [Gol91, ABH19, SB95, Yu11]. highly-available [Yu11]. hoarders [KFH12]. hoc [AMZ12, CGR02, GNZ17, HWY19, KP05, Pel07, DGW05]. hole [DFS06]. holy [SM94]. homonyms [DGFTT13]. hop [BYGI91, FNP01, GNZ17, RSZ13]. hopbound [EN22]. hopsets [EN22]. hot [BAES95, FK00]. hot-potato [BAES95, FK00]. Hundreds [FR03]. husbands [MDH86]. hybrid [AKS18, Fri95, WS07]. hypercube [BT95]. hypercubes [KS01]. hyperfairness [AFG93, Lam00]. HyperTree [DK08].

identical [Koz12]. identifying [CSC11]. identities [BK07, GNZ17]. identity [BPV06]. image [KW86]. images [AP91]. imitation [ABH16]. impact [MS10]. implement [LH00]. Implementation [HMN12, SA95, ANH95, BBS12, HMN14, KS98]. Implementations [CRR23, AM97, FK17a, GW19, GHW12, Gon95]. Implementing [CB01, Fri95, Lam08, ADGT08, BMP06, DGW05, HT24]. importance [BPV06]. impossibility [BHS18, FR03, FLM86, TB93]. improve [AM97, Moi01]. Improved [EGN23, Fis20, GHU20, GHM21, AK02, CM22, vLT87]. in-network [GVB21]. inaccurate [FP10]. Incentive [FRS11, LM16]. Incentive-compatible [FRS11]. incomplete [Bei07]. incorruptible [VK20]. Increment [CRR23]. incremental [KS03, PS91, Pix88]. independence [BE13, SW10]. independent [AAK13, HK18, JSX16, SW10]. independently [DL12]. indeterminate [KP86]. indulgence [DG05]. infinite [Bro87]. infinitely [CM05]. Information [CST15, AMZ12, ABL13, CC95, CHM19, DGFT11b, FHK17, FGP09, FP10, KS98]. Inherent [AKS18, DG05, FHS06]. inhibition [CT96]. input [Hie12]. insensitive [Bla86, Ebe91, Mar86, Udd86, Ver88]. inter [YG09]. inter-object [YG09]. interaction [HS12]. interactions [AFG93, MNS17]. interactive [BOEY03, Bro92]. interdomain [FRS11]. interface [Bro92]. interfaces [HMN12, LS92]. interference [DJ09, GGL13]. interleaving [ABS18, KP92]. Internet [PST13]. interprocess [Lam86a, Lam86b, TB93]. intersection [SS01]. Interval [FGM01, SS01, Vog91]. interworkings [Fei99]. Introduction [And01, Aro07, AR03, DS08, Fic05, Gue06, Had97, Kut06, Mar05, Mer93, Mer99, Mos18, Raj05, Sch92, Sha02]. invocation [Caz03]. ISPs [PST15]. Issue [Fra07, Agu14, Alv11, Att16, Att17, Att19, Att20, Att21a, Att21b, Att22a, Att22b, Att23, Att24, DS08, Fic05, Fra13, Kei11, Kut06, LS13, Mos18, Pan15, PS10b, Pel08, Pel13, Tau10, Tau15, Wat09]. issues [Cla86]. iteration [AAG93]. Iterative [TLV24, GRS04, LW05].

jamming [KPY18, RSZ13]. join [DGHP23]. joint [KS92b]. journal [Gou86].

Keeping [MS97]. key [AJM^{+}23, DFY05]. Knowledge [FHV97, FP15, MT93, CGM12, HF89, Hal00, HM04, HR05, HP10, KSV13, MM08, MDH86, NT93, PPS17, PT92]. Knowledge-based [FHV97, Hal00, HM04, HP10]. knowledge-theoretic [HR05]. Known [KP07, GPX07, GKS08].

labeling [FPSP19, Kor07, KP08b, KKP10]. labels [FPP01]. Lake [Lam92]. language [AO86]. languages [AFK88, BBM01]. Laplacian [ALH^{+}23]. large [BPS11, DKZ17, HK18, Yu11]. large-scale [Yu11]. Last [SW22]. Last-use [SW22]. latency [AFRT24, BCG24, Zie08]. latest [MS97]. lattice [AHR95]. law [FG14]. layer [KLN11]. layouts [FNP01]. Laziness [Moi01]. Lazy [LLR99, Ger99, JPZ99, Moi01]. Leader [DP14, GLS12, AAE08a, BGJ07, BT18, BKM96, CPR20, DS18, FPP01, FP11, FP15, GMP23, IRW13, MOY12, RFS07, SAAAA95, Sin97]. Leakage [BGK14]. Leakage-resilient [BGK14]. learn [CM86a]. learning [EMGR22, SV19]. legitimate [ALK15]. lemma [CPS17]. level [AR97, Fri95, FP15]. Leveraging [HWY19]. Life [CHT20]. limitations [AKS18, Bla86, CG98]. limited [BNBOD89, DFV22, FHK17, HP05, KKN^{+}24]. limited-scope [HP05]. LiMoSense [EKR14]. line [AP91, CKO19, DS03]. Linear [EN22, GHS17, AP91, DS18, FM04, FGM01, KW86, Myo93, NK24, Pet10, RT16, ZIP95]. Linear-in- [GHS17]. linear-messages [ZIP95]. Linear-size [EN22]. linear-time [FM04]. Linearizable [CRR23, HSW96, MMT08]. linearization [Hes94, Hes95]. Linial [MT22]. link [AH95, GS91, NK24, WS07]. links [SAAAA95]. Lipschitz [SS01]. list [AADM23]. lists [MT22]. live [BCG22, EKR14]. lived [AST02, BHMT23, BEW11, WSS19]. livelock [FK00]. Liveness [BCG24, AT10, AS87, SL87]. Load [KPT11, AFS11, ABS16, GVM23, LW16, RB06]. local [AM97, BGKO23, CHPS20, CHM19, CPS17, DGFT11b, KA06, KA07, KSV13, MNS17, MS16, Waw15, BBS21, GHS17]. local-spin [AM97]. Locality [FRT13, BCD13]. Localization [CKP15]. localized [GP03, KSV13, PT14]. Locally [BBC^{+}23, AFP20, CIK11, FSY20]. location [CPP02, HP15, Sli07]. Lock [DMS02, GGH05, LMS08]. Lock-free [DMS02, GGH05, LMS08]. log [CKP12]. log-space [CKP12]. logarithmic [HP15, HW11]. logging [ABM02]. logic [BK98, HJV15, LV10, PJ91]. logical [GS98, Ksh04, TRA99]. logically [CMS18, Gol91]. loneliness [GLS12]. Long [AST02, BHMT23, BEW11, GGL13]. Long-lived [BHMT23, BEW11]. long-range [GGL13]. Look [DDN14, ACR19]. look-compute-move [ACR19]. LOTOS [KHvB96, KK99]. Lovász [CPS17]. Low [CNV05, DB93, HIZ21, KKI21, Zie08, ALH^{+}23, AHS21, AB05, CDPP24, DJ09, HST12, NdM16]. Low-Congestion [HIZ21, KKI21, ALH^{+}23]. low-contention [AB05]. Low-cost [DB93]. low-delay [AHS21]. low-interference [DJ09]. Low-latency [Zie08]. low-space [CDPP24]. Lower [BV03, BD97, CEL23, FK17a, GVB21, Gon95, HKK08, KW21, Lam06b, ACHL20, AK02, BSD88, BUW20, CR10, GHS17, JYZ20, KA12, PST16]. lowest [FPS05]. lowest-cost [FPS05].

M&M [HHT22]. MAC [BMIY08, KLN11, ORZ18]. machine [BCG24, DFV22, EMGR22, KW86]. machines [BG14, LM91]. made [KS92a]. Maekawa [BC95, Sin91]. Maekawa-type [BC95, Sin91]. mailbox [AGL10]. Maintenance [LMP06]. major [AKH03]. majority [AAE08b, BER21, CMQR23, MNS17]. Making [BCG22, CHGH19]. malicious [MS03]. management [HG01, SB12]. managing [AR97]. Manhattan [CMS13]. many [CM05, FOV08, GGT18, MS18]. many-cores [GGT18]. Mapping [KW86, Myo93]. mappings [LA92]. markets [CHM19]. masking [BKA12]. Mass [ABM17]. Mass-Distribution [ABM17]. massively [CDPP24, NS22]. matching [DGHP23, Fis20, HRZ20, KPS10, KY11, PS10a]. maxima [CMN02]. maximal [AAK13, DGHP23, JSX16, SW10]. Maximum [CKR09, GGP97, HRZ20, KY11]. me [DGHP23]. means [SCA15]. meassage [KP93]. meassage-passing [KP93]. measured [PST13]. mechanical [Hes95]. Mechanism [FSS06, AK01, FPS05]. mechanisms [DMS15, LM16, Moi01]. media [AB93]. mediated [MOY12]. meet [CKP12]. Meeting [DFY20b, GGK11, CDN18]. meeting-points [CDN18]. meets [ABM17]. membrship [Cri91]. memories [BMP06, GHS10]. memory [ACM06, ANH95, AKS18, AMW05, AKH03, AHS21, ABS18, AFR06, Asp12, Att16, BHS18, BHS22, CLM17, DGW05, FP11, GLS10, Gra99, GGT18, GR07, HS07, HG01, IFR20, MT93, MNS17, SCA15, SB12, SB14, ST97, SW22]. mesh [BAES95]. message [ABM02, AWH92, AN05, BKN19, CLM17, CGM12, CCW95, CDS01, CH01, CEW13, DGFT11b, FM04, FGM01, HL89, HRZ20, KS98, OBG08, PS02]. message-complexity [FGM01]. message-logging [ABM02]. message-passing [AN05, CDS01, DGFT11b, HRZ20]. messages [BND91, GSM07, GM21, KPS10, LPSP19, MK21, ZIP95]. metamorphic [WWA04]. method [Caz03]. methodology [Lin88]. methods [CHKS19]. metric [HP15, HS07]. metric-space [HS07]. middleware [CBP02]. migration [ML97]. Minimizing [BKN19]. minimum [DGFT11b, FM04, FO23, GM87, KP08a, KK07, LPW13, MK21, SB95, Waw15]. minimum-weight [FO23, GM87]. MinMax [CBM21]. MIS [BE10, MRZ11]. Misleading [PST13]. missing [FP10, MG04, NK24]. mitigation [SCA15]. mixing [AW09]. mobile [AAP06, AMZ12, BLT16, BD97, DFY15, DPBRP20, DFV22, DFS06, DGW05, FPV17, KK15, PPV15]. mobility [Gor10]. modal [HJV15]. mode [KKS10]. Model [BK98, GHS10, WS09, ACT00, AKV18, Asp12, BBS21, CHHM15, CBS09, CM22, CW90, DDN14, DRS23, EFT93, GM89, GHS17, HM16, HRZ20, Jon94, KPT11, LMR21, MK21, NDO86, Udd86]. Modeling [NL11, KS03]. Modelling [HF89, vGHT16]. models [BSD88, BKT15, BBS12, CMQR23, EZ06, HJV15]. modify [FH04]. modular [Asp12, Myo93, Rhe98, WL93]. modules [KP86, LS92]. monitoring [DS03, EKR14, NYD21]. monitors [Gje88]. monotonicity [INS09]. mortal [WBB12]. most [CCW95]. Move [DDN14, ACR19]. MPLS [ABBM02]. MST [KKM15, LPSP06]. much [FP11]. Multi [And94, CG21, AFG93, AK09, BYGI91, DKZ17, FDS95, KKS17, MP18, PMCP22, Pix88, RSZ13]. multi-agent [KKS17]. multi-commodity [AK09]. multi-hop [BYGI91, RSZ13]. multi-mutator [Pix88]. multi-party [AFG93, DKZ17, MP18]. Multi-shot [CG21]. multi-threaded [PMCP22]. multi-token [FDS95]. Multi-writer [And94]. multicaches [Bro90, Liv93]. multicast [Gol91, MMR00]. multicomputer [HL89]. Multidimensional [MHG15]. multilevel [GS98, Tan94]. multiple [AGE98, BKK16, CKR09, HWY19, Lin88]. Multiplicity [CRR23]. multiprocess [PZ86]. multiprocessor [BT95, EGZ20]. mutator [Pix88]. mutators [Jon92]. Mutual [BGM86, AK01, AK02, AKH03, AB02, BKK16, CSZ10, CS94, DG10, GR19, HW11, Her00, Jou00, KA06, KA07, KA12, Sin91, YA95, vdS87].

Naming [ES94, CP89, PPV98]. Nash [BE10, BFM21]. Nash-Williams [BE10]. nature [JSX16]. Near [LR19, PP24, ZYC16, BPS11, EN18, PPV15]. near-cliques [BPS11]. near-gathering [PPV15]. Near-optimal [LR19, PP24, ZYC16]. Nearly [BNNPS02, HM16, FZ15]. Necessary [KS98, MMSA93]. needed [FP11]. neighborhood [BE13]. neighbors [Sli07]. nested [Ruk91]. net [GM89, NRT90]. nets [BKS89, YEFH03]. network [AFP20, AMS20, BYGI91, BFM21, CHHM15, CPP02, FPP01, GVB21, GM87, Gon95, HTW07, HS12, JLC07, MS16, PST15, Sin99]. Networks [FK00, KP07, AA95, AB19, ABH19, AAP06, AMZ12, BK07, BM05, Bei07, BEF16, BHK^{+}24, BG93, BDD16, BR91b, Bro87, BKM96, BD97, BDV07, BMIY08, CHGN14, CHCGS23a, CHCGS23b, CGM12, CDP16, CC88, CPR20, CCS17, CGR02, CKL06, CKV22, CDN08, CNS21, CEW13, CK20, DKZ17, DGHP23, DJV00, DP14, DFN18, DFS06, DGW05, EGS09, EKP16, EKR11, EKR14, FH04, FG14, FP10, GGL13, GKM07, GPX07, GKS08, GHK15, GLS10, GNZ17, GMP23, GHSW23, HKR18, HWY19, HP01, HSW96, HT06, HS07, HL89, HC93, HW97, HT09, INK95, IRW13, JYZ20, JLC07, Jon94, Jou02, KP86, KP05, MMT08, MS10, MNS17, MSS22, MP16, Mit97, MW08, NL11, NDO86, OR96, PR01, Pel07, PST16, RSZ13, RB06, SDP15, SAAAA95, SHD21, SB14, Waw15]. networks [WSS19, Yu11]. No [GS14, BYGI91]. Node [BGKO23, BKT15]. noise [CHGH19]. Noisy [FN19, ABH19, FHK17]. Non [Gue02, Sha19, DGFT11b, EM09, Moi01, SV19]. non-Bayesian [SV19]. Non-blocking [Gue02, Sha19, Moi01]. non-FIFO [EM09]. non-local [DGFT11b]. Nonatomic [KA06]. nonblocking [HLS06]. nonuniform [EHT07]. note [Hal00, Liv93]. notions [NRT90]. novel [GND07, LM95]. null [GM21]. number [ABS18, GKS22b, GKS22a]. numbers [GK11].

object [ADL90, AM97, CB01, LH00, Sli07, YG09]. object-based [ADL90]. Objects [MMT03, HK09, Tau20]. Oblivious [LDB05, Asp15, AB05, DFY15, DFY20b, DFV22, DN17, JYZ20]. obstruction [BRS18, GKK08]. obstruction-free [BRS18]. obstruction-freedom [GKK08]. obstructions [Hav00]. odd [HL99, HHH03]. offs [BER21, CKP14]. omega [ADGT08]. omission [MNV11]. on-line [DS03]. On-the-fly [Jon92]. once [BKT15, CCW95]. ONE [CPP02, BNNN90, BND91, LH00]. One-bit [BNNN90, BND91]. one-resilient [LH00]. online [CSC11]. only [BKT15, DIM93, IFR20]. onto [KW86, Myo93]. opacity [SW22]. open [GGH05]. Operational [KS92b, ŢH09]. Operations [CRR23, Fri95, GHW12, KW86, MT94, Sha19]. operator [GM89]. operators [DT01]. opinion [BHK^{+}24]. Opportunistic [AMZ12]. Optimal [AP91, BMP06, BBEiKS23, DN17, GP21, INK95, KP07, MPSL04, SAAAA95, YG09, ZIP95, ACM06, ADS22, ALH^{+}23, ABL13, BNNPS02, BGJ07, BOEY03, CDN18, CPP02, DFS06, DDM16, EN18, FM04, Fek90, FZ15, Gel18, Gon95, GHSW23, HM16, KS98, LR19, MRZ11, MVP07, Mit97, PP24, PCR14, SW10, VBA22, YEFH03, ZYC16]. optimal-message [FM04]. optimistic [ABS87, LMS08, VJ94]. Optimistically [GM21]. optimization [AK09]. Optimizing [KFH12, TK06]. Order [ABL13, BBM01, KU92]. Ordered [KKM19, CBMT96]. ordering [KS98]. orientation [KM02]. oriented [KS01, KS03, MRT08]. other [MDH86]. output [AFP20, Hie12]. Overcoming [Hie12]. overhead [DGK07]. overlay [BK07, GKM07, GHSW23, JLC07]. overview [Ver88].

P [PJ91]. Packet [KKS10]. packing [KY11]. papers [Agu14, Fra13, Gue12, Kei11, Mos18, Pan15]. Parallel [BT95, DH07, CDPP24, KS92a, KKS17, LW16, NS22]. parallelism [EFT16]. Parameterized [AKV18]. parameters [KKI21]. Parsimonious [BCF11]. Partial [ADGT12, CDG^{+}24]. partially [BM05, CGM12, NYD21, VK20, WS07]. particles [DFY20a]. Partition [CHK14]. partitioned [BGM86]. parts [FHS22]. party [AFG93, DKZ17, MP18]. passing [AN05, CLM17, CGM12, CDS01, DGFT11b, HRZ20, KP93, OBG08]. Passive [MHN18]. passively [AAP06]. path [ABBM02, ACM86, AK01, Hua06, OR96]. paths [ABBM02, CHDL21, CM22]. Patricia [Sha19]. patrolling [KK15]. Pattern [CDN19a, CDN19b]. patterns [BKN19, DFY15, Mic18]. Paxos [ACM06, CM05, GL03, Lam06a]. payment [vGGT23]. pays [Moi01]. Peer [GKM07, BEF16, DK08, JLC07, KSW10, RB06]. peer-to-peer [BEF16, DK08, JLC07, KSW10, RB06]. perfectly [BNNPS02]. performance [ACY21, BC95, BR91a, CCT22, CSZ10, CB01, LLR17, SA95]. PerformERL [CCT22]. Performing [CDS01, CKL06]. periodic [BNNPS02]. periphery [CNS21]. permanent [MHS14]. Permissionless [KPPT23]. perspectives [Lyn03]. Petri [YEFH03]. Phase [CNS21, CMQR23]. philosophers [DFP04, Jou02, WL93]. PIF [BDV07]. placement [BO22, HKR18]. planar [GK93, Waw15]. planarity [LMR21]. plane [BBP19, DP15, HP01]. platform [CBP02]. Plausible [TRA99]. plurality [BCT17, FN19]. PODC [Alv11, Att17, Att19, Att20, Att21a, Att21b, Att22b, Att24, Fra13, Gue12, Lyn03, Pan15, PS10b, Tau15, Wat09]. PODC’2004 [Kut06]. points [CDN18]. policy [CPP02, FSS06, RGB13]. polygon [DFY20b]. polylogarithmic [BHMT23]. polynomial [BBP19, BDL22, DP15, FM11]. polynomially [BFM21]. population [AAR07, AAE08a, AAE08b, BER21, CFH^{+}22, CEL23, DFV20, DS18, EJWK21, MOY12]. posets [TK06]. potato [BAES95, FK00]. power [AAR07, CM10, DGFT11a, FG14, HS91, Kes95, KKN^{+}24, Ksh04, LH00, MRT08]. power-law [FG14]. powerful [LL15]. powerline [PST16]. Practical [Lis93, RB06, AKT14, PW97]. practically} [VK20]. practice [SB12]. Precision [NYD21]. predicates [CG98, Sto00]. Preface [Hoa86]. preorders [CC95]. prerelease [SW22]. presence [AH95, GS91, WBB12, Zie08]. Preserving [VK20]. prevention [HMR00]. price [DG05]. pricing [HTW07]. primal [PS10a]. primal-dual [PS10a]. primitives [FHS06, GHW12]. priority [KU92]. private [Bei07]. Probabilistic [AM05, Cri89, MS89, Asp12, BK98, Ciu94, GS98, KKT12, PZ86]. probabilistic-write [Asp12]. problem [AGT10, AGL10, BER21, BD97, CS94, EK11, FK17b, Fin94, GGP97, GM04, GHV01, Hua06, Jou02, Waw15]. problems [BBS21, BBC^{+}23, BGKO23, FLM86, Hie12, PS10a]. process [BKS89, Gor10, GHV01, TV02, WS07]. processes [ABS16, AMW05, BR91b, CM10, CM86a, CM05, ES94, Gel18, Gje88, GT91, Jal89, Jos88, MMT03, Mar86, NDO86, WBB12, vdS87]. processing [KW86]. processor [Cri91]. processor-group [Cri91]. processors [CDS01, vLT87]. profit [AMZ12]. program [EM09, San12]. programmable [DFY20a]. programming [ANH95, AFK88, HM04, Myo93]. programs [BKA12, CM86b, FHV97, Hal00, HR94, KP92, LW86, NYD21, PJ91]. Progress [Hes99, AT10, Tau20]. prone [BR91a]. Proof [KKP10, AT12, BG87, Dij86, FPSP19, HH87, Hes94, Hes95, NDO86, PJ91]. proof-labeling [FPSP19]. proofs [FFP21, FLM86]. propagation [BMP06]. propagation-based [BMP06]. properties [AFP20, CHFV19, LA92, MS94, SS94, SL87]. Property [LMR21, AT10, SW22]. Protocol [YEFH03, AAE08b, CMN02, HR99, KHvB96, KK99, LMM07, LM91, MOY12, PS91, SB95, vGHT16]. protocols [ABM02, AAR07, AAE08a, Asp03, BMP06, BLT16, BER21, BKN19, BKM96, BMIY08, CEL23, DFV20, DFS06, DS18, EJWK21, Fin94, GP21, GGP07, Gol03, Gon95, GND07, JMP10, Lin88, LM95, MMR00, MS89, Mil87, MMSA93, MRR04, OR96, Pac91, PS02, PZ86, Sin99, vGGT23]. prove [LA92]. Proving [BBM01, JPR99, SL87]. pseudo [BGM93]. pseudo-stabilization [BGM93]. public [DFY05]. public-key [DFY05]. purely [CHKY18].

QoS [AFS11, KKS10]. queries [PSS08, Yu11]. Queues [CRR23, KU92, LMS08, ST16]. queuing [KKS10]. quorum [Baz00, Baz01, CMN02, GV10, Jou10, MR98, NW05, PW97, Yu06]. quorum-based [CMN02]. quorums [AM05].

Radio [KP07, BYGI91, BD97, CHGN14, CGR02, CDN08, DGHP23, EGS09, EKP16, FP10, GGL13, GPX07, GKS08, GHK15, KP05, MW08, NL11, Pel07]. Rambo [GLS10]. random [ABS18, BEF16, CIK11, CEW13, GKM07, KKS17]. randomised [BEF16, JSX16]. Randomization [MNV11, AKW06, MS10]. Randomized [AAK89, Asp03, BGJ07, BKK16, DFP04, FGP14, FPSP19, GHK15, HW11, LW05, PPV98, RFS07, TV02, AT12, AAS18, Asp15, DDM16, GW19, HHT22, LW16, MRZ11, PSL00, VBA22]. randomness [Her92]. range [GGL13]. Rank [KU92]. rational [DMS15]. re [MG04, YEFH03]. re-execution [MG04]. re-synthesis [YEFH03]. Reaching [AJM^{+}23, Cri91]. reaction [CCS17, SHD21]. Read [CRR23, FH04, HG01, BRS18, DIM93, GHW12, CPP02]. Read-modify-write [FH04]. READ-ONE-WRITE-ALL [CPP02]. Read/Write [CRR23, BRS18, DIM93]. reading [Vid90]. real [MPSL04, MS14, SL87]. real-time [MPSL04, MS14, SL87]. realistic [CW90]. really [BPV06]. recall [NYD21]. Recognizing [AS87]. reconcile [SS01]. reconciliation [MP18]. Reconciling [FS12, MS14]. reconfigurable [CBP02, GLS10]. reconfiguration [KT22, MSS22, WWA04]. Recoverable [GR19]. recovery [ABBM02, ACT00, BG87, DGK07, HH87, VJ94]. recurrence [Raj89]. recursion [MdV89]. RedBlue [FK20]. reduce [Her92]. reduced [ABS87]. Redundancy [FFP21]. reference [DMS02]. Refined [GV10]. refinement [JPR99, Vog91]. register [MT94]. registers [And93, And94, BNP97, BRS18, HHT22, HT24, LW05]. regret [KPT11]. regular [BBM01, BEF16, BV03]. Reischuk [CDG^{+}24]. related [CSZ10, INK95]. relations [HMN12, HMN14]. Relationships [AFR06, SM94]. relaxed [ST16]. relevant [AG07]. Reliability [Ciu94, ADGT08, ŢH09]. Reliable [ABH19, BSD88, PPS17, AFRT24, AFR06, BM05, CH01, EM09, MMR00]. relocatable [EKS16]. Remote [Caz03]. removing [KSV13]. Renaming [AKN12, OBG08, BEW11, CR10]. Rendezvous [CDP16, AKV18, CKP12, CKP14, MP16]. rendezvousing [AHM13]. repeated [BCP19]. replace [Sha19]. replicas [CPP02]. replicated [BG87]. replicating [YG09]. replication [BCG24, FG01]. report [GM04]. representative [KP92]. request [CB01]. requests [RB06, TL95]. required [AR97]. requirements [Had87]. requires [DS18]. research [AKH03, FM03]. resilience [ACM06, ADS22, PCR14]. Resilient [BOEY03, BGK14, CNV05, LH00, LM16, MR03]. Resilient-optimal [BOEY03]. resistant [KSW10]. resolution [CDN18, FGN19]. resource [KPY18, Rhe98]. resource-competitive [KPY18]. resources [Jou10, TL95, YEFH03]. respect [PPS17]. response [FM11, KR90, Kes95]. restartable [CDS01]. Restoration [ABBM02]. restricted [Had87, LPW13]. restrictions [CHPS20]. results [BHS18, FR03, TB93]. resume [IFR20]. Reviewer [Ano21, Att12]. revisited [ACD^{+}23, Vid90]. Revisiting [ADS22]. revoking [DFY05]. rewriting [BBM01]. rigorous [CDN19a]. ring [AAK89, BT18, GS91, KKS17]. Rings [LMP06, BGJ07, DDN14, DNN17, DDS20, DN17, FDS95, HL99, HHH03, INK95, Sli07, vLT87]. RMR [GHW12, HW11]. RMR-complexity [HW11]. RMR-efficient [GHW12]. RNC [EN22]. robot [ACR19, BLT16, WWA04]. robots [CDN18, CDN19b, CKP15, CKO19, DFY15, DPBRP20, DFY20b, DFV22, DN17, FPV17, KKN^{+}24, KW21, PPV15]. robust [AAE08b, CHGH19, GLS10, HK09, JLC07, KS92a]. robustness [MRT08]. rollback [ABS87, IFR20]. rollback-only [IFR20]. rooted [BBC^{+}23, DJV00, Mit97, WSS19]. rotor [KKS17]. rotor-router [KKS17]. round [HK18, NK24]. rounding [Fis20]. router [KKS17]. routers [TK06]. routing [BAES95, BCD13, DS86, EN18, FK00, FPS05, FSS06, FRS11, FG97, FGM01, FG14, Haj91, HP01, HL89, LPSP19, RGB13, RT16, vGHT16]. rumor [AE18, DDM16, FN19].

Sade [ORZ18]. safe [EM09]. Safety [DS03, AS87, SL87, SW22]. same [CP88]. sampled [FSY20]. sampling [GKM07, Yu11]. satisfiability [TS11]. satisfy [LS92]. Scalable [AB19, DMS15, DFY05, NW05, AHM13, HKK08, JLC07, YA95, ZIM15]. scale [Yu11]. scan [AP91]. scan-line [AP91]. scanner [FK17a]. scheduler [AB05, BGJ07]. schedulers [LM95]. schedules [BNNPS02, MPSL04]. scheduling [BHS22, HM16, MRS06, PS10a]. scheme [GS14]. schemes [EN18, FG97, FGM01, FPSP19, Kor07, KP08b, KKP10, RT16]. scope [HP05]. scrip [KFH12]. Search [CKO19, AKT14, SM94]. Searching [DFS06, LLR17, BDK15, DNN17, INS09, KW21]. secret [DMS15]. Secretary [CHM19]. section [Gue12]. Secure [DKZ17, MMR00, Yu11, GNZ17]. security [GS98]. selected [Agu14, Fra13, Gue12, Kei11, Pan15]. selection [JSX16, LLR17]. Self [AB93, BCP19, DJV00, DPBRP20, DIM93, DT01, FDS95, HT06, Her92, HC93, HW97, HL99, KP93, AKT14, BBM01, BGJ07, BT18, BKN19, BV02, CSZ10, CMN02, Dij86, DK08, FZ15, FMlP06, GK93, GGP97, GGP07, Hua06, KC99, KT12, KKM15, LR19, Mil87, MOY12, PT14, PRT16, SS17, VBA22, ZB92]. self-adjusting [AKT14, ZB92]. self-healing [PT14, PRT16, SS17]. self-stabilising [LR19]. Self-stabilization [AB93, DIM93, DT01, Her92, Dij86, FZ15, FMlP06, KT12, VBA22]. Self-stabilizing [BCP19, DJV00, DPBRP20, FDS95, HT06, HC93, HW97, HL99, KP93, BBM01, BGJ07, BT18, BKN19, CSZ10, CMN02, DK08, GK93, GGP97, GGP07, Hua06, KC99, KKM15, MOY12]. self-synchronizing [Mil87]. selfish [HTW07]. Semantics [Bro87, Gje88, KP86, Mel91, MdV89, PS02, ŢH09, Vog91]. semi [BUW20]. semi-synchronous [BUW20]. semiwords [Vog91]. sense [BV03, BDK15, Sin97]. sensitivity [FGP09, NYD21]. sensor [BMIY08, EKR11, EKR14, Yu11]. sensors [AAP06, EKS16]. separator [KP08b]. sequences [DFY15, HU08, KP92]. Sequential [Ger99, JPZ99, LD99]. Sequentially [MMT08, Gra99]. Serializable [KS92b, Bro92]. servers [MR03]. service [GLS10, HKR19, KHvB96, KK99]. Set [CRR23, AAK13, ADGT12, BRS18, CHT20, GK11, GW19, HP05, JT21, JSX16, JMP10, KW05, LPW13, MP18, MRT08, SW10, TV02, Waw15, Yu11, Zie10]. Set-Linearizable [CRR23]. Sets [CRR23, HK18, JRS02]. setting [EKP16]. settings [ABS16]. several [Jon92]. shades [GMP23]. Shape [DFY20a]. shapes [Mic18]. Shared [AKH03, ACM06, AMW05, AFR06, Asp12, BKT15, CLM17, ES94, GR07, HK09, KKM19, LH00, MMT03, MT93, PPV98, Tau20]. Shared-memory [AKH03, Asp12, GR07]. sharing [DMS15, HHM10]. short [WSS19]. short-lived [WSS19]. shortcut [KKI21]. shortcuts [ALH^{+}23, HIZ21]. shortest [CHDL21, CM22, Hua06, OR96]. shortest-path [OR96]. shot [CG21, GKS08]. shrinks [FGP06]. shuffle [DL12]. sides [CP88]. signals [Raj89]. signature [LDB05]. signature-based [LDB05]. Signed [Yu06]. Simple [BCT17, DDM16, MS16, MP18, ABS16, AAE08b, BN94, HR99, JSX16, PR01, ST87]. Simulating [ST87]. simulation [BGR01, BT95, CP88, DFV20, GT91]. simultaneous [AGT10, NT93]. since [AKH03]. Single [CM22, BYGI91, FK17a, GNZ17, HK18]. single-hop [BYGI91]. single-scanner [FK17a]. Single-source [CM22]. SINR [HM16, ORZ18]. size [AP91, BKN19, EN22, HL99, HHH03, Lis91, Pet10, TRA99]. sized [HLS06]. skeletons [Pet10]. sketches [SDP15]. Sketching [XTB08]. skip [AADM23, AW09]. skip-list [AADM23]. slicing [MG05]. sliding [PS91, XTB08]. sliding-window [PS91]. small [EFO22, EN22, FGP06, FG14, JRS02, KPS10, LPSP19, MNS17, PP24]. small-world [FG14]. smart [DGK20]. Smoothed [DFN18]. smoothing [HT06, MPSL04, MS10]. SMR [NK24]. Snap [BDV07]. Snap-stabilization [BDV07]. snapshot [FK17a]. snapshots [AHR95]. social [BHK^{+}24]. Soft [SCA15]. Soft-error [SCA15]. Software [ST97, GGT18]. Solo [HS08]. Solo-valency [HS08]. solution [Mic18]. solutions [CS94]. solvability [MRR04]. solve [EHT07]. solvers [ALH^{+}23]. solving [DGFT11b, TS11]. Some [JYZ20, Lyn03, PR01, TB93, KW21]. sound [KGM99]. soup [MS18]. source [CM22]. space [BGJ07, BER21, CDPP24, CKP12, CKP14, EFR08, Gel18, HS07, Her92, MOY12, RT16, Tau14, VK20]. spanners [CHKY18, DJ09, EZ06, EGN23, Par17, Pet10]. spanning [FM04, FO23, GM87, KP08a, KK07, MK21, SB95]. Sparse [HP01, BE10, PR01]. spatial [LV10]. spawning [Mel91]. speaking [FHK17]. Special [Agu14, Alv11, Att16, Att17, Att19, Att20, Att21a, Att21b, Att22a, Att22b, Att23, Att24, Fra07, Fra13, Gue12, Kei11, LS13, Pan15, PS10b, Pel08, Pel13, Tau10, Tau15, Wat09, DS08, Fic05, Kut06, Mos18]. Specification [LM91, Bro92, HH87, KS92b, PS91]. Specifications [LW86, Hal00, KHvB96, KK99]. specified [Fin94]. Specifying [LS92]. spectrum [CT96, HHM10]. speculative [ŢH09]. Speed [CCS17]. speeds [KK15]. spin [AM97]. spinning [KA06, KA07]. spite [CDP16]. splay [AADM23]. splay-list [AADM23]. splitter [AST02]. splitting [GHU20]. spoil [MS18]. spreading [ABL13, AE18, CST15, DDM16, FN19]. squad [GM04]. squads [LR19]. SR [AO86]. stabilising [LR19]. stability [CDPP24, WSS19]. Stabilization [BGM93, AB93, AN05, BV02, BDV07, Dij86, DIM93, DH07, DT01, FZ15, FMlP06, Her92, KT12, VK20, VBA22]. stabilizing [BBM01, BGJ07, BCP19, BT18, BKN19, CBM21, CSZ10, CMN02, DJV00, DPBRP20, DK08, FDS95, GK93, GGP97, GGP07, HT06, HC93, HW97, HL99, Hua06, KC99, KP93, KKM15, MOY12]. Stable [DS18, ZIM15, KPS10, MS94, MS16, SS94]. Stacks [CRR23, AGM07, ST16]. stamps [IL93]. stars [PST13]. state [AAP06, BG14, BCG24, CEL23, Jos88, LS92, Mil87, VK20]. state-based [Jos88]. state-machine [BCG24]. states [ALK15, CSC11]. stealing [HLS06]. Step [MdV89, BHMT23]. stigmergy [Koz12]. Stochastic [GVM23, BKN19]. stopping [AFRT24]. stopwatches [DL12]. stories [MDH86]. strategies [CIK11]. streaming [EZ06]. streams [MPSL04, XTB08]. stretch [HST12]. Strong [KG22, SS94, SW22]. structure [HST12, Tau20]. structures [GP03, ST16]. Structuring [CHGN14, Sin99]. study [DL12, MDH86, PSL00]. Sub [HP15, HW11]. Sub-logarithmic [HP15, HW11]. Sublinear [EFO22]. Sublinear-time [EFO22]. Sublogarithmic [BE10, GS14]. sublogarithmic-time [GS14]. success [vGGT23]. sufficient [KS98, MMSA93]. sum [FHS22]. Superstabilizing [Her00]. support [SW22]. survivor [JMP10]. suspend [IFR20]. suspend/resume [IFR20]. swarms [ZIM15]. switches [KKS10]. SWMR [HT24]. SWSR [HT24]. sybils [KFH12]. Symbolic [BKA12, EFT93]. symmetric [ES94]. symmetry [FP15]. synchronization [CM10, Cri89, DB93, EGZ20, FL06, Fri95, GM04, HS91, Lam03, MG04, Moi01, NK24, TB93, WS07]. synchronized [Lis93, VK20]. synchronizing [Mil87]. Synchronous [Baz00, CBMT96, MRR06, AFRT24, BV02, BUW20, CDS01, Cri91, GSM07, GRS04, Gol91, GM21, NYD21, OBG08, WS07]. synchrony [ADGT08, ADGT12, AMZ12, CDG^{+}24, WS09]. synthesis [BKA12, San12, YEFH03]. Synthesizing [Raj89, VBA22]. system [ADL90, BKT15, CKP15, GSM07, Hie12, LS92, MS97, NDO86, PJ91, Sch97a, Sch97b]. systems [AM05, AG07, ADGT08, AMW05, AKV18, AAG93, AN05, AFR06, BG14, BC95, BGM86, Baz00, Baz01, BBM01, BEF16, BR13, CP88, CGM12, CBS09, CSC11, CP89, Cri91, DGFT11b, DIM93, DK08, FHS22, FS12, Gje88, GM89, GVM23, GS98, Gue02, GV10, HHT22, HF89, HMN12, HT24, Jou10, JMP10, KFH12, KP93, KT12, Koz12, KSW10, KS03, Lis93, LM91, MR98, MM00, MHG15, MT93, NW05, NRT90, OBG08, PT92, PW97, Pel91, PMCP22, Pix88, SS94, SL87, Sin91, Sto00, TRA99, Udd86, WS07, Yu06]. systems-on-chip [FS12]. Systolic [CM86b, KR90, KW86, Lis91, Myo93, Raj89]. systolizing [LBH91].

tables [GGH05]. talkative [BT18]. talking [Jou02]. TAP [CHD20]. tasks [CDS01, DGFT11b]. taxonomy [Gor10]. technique [ACHL20, BG87, LM95]. Techniques [CG98, MG05]. temporal [NDO86]. Terminating [Mic18]. Termination [HT23, ACR19, AN05, Fin94, HR94, LMM07, Mat87, MVP07]. test [GW19, HU08, HMN12, HMN14, TV02]. test-and-set [GW19, TV02]. testing [CCT22, CHFV19, Hie12, LMR21, MHN18, San12]. their [CG98, KS98]. them [KKM19]. theorem [Lam90]. theoretic [CH01, HR05]. theoretical [SS17]. theory [ACY21, Bla86, FM03, FG01, San12, SB12]. Theta [WS09]. Theta-Model [WS09]. third [CSZ10]. threaded [PMCP22]. threads [Lam08, SCA15]. threat [PST15]. Three [Lam92]. Threshold [JMP10]. throughput [CKR09, HK09, RSZ13]. Tight [AMW05, HP05, LW16, Tau14, BUW20, CDG^{+}24]. Time [Ano00, BER21, BHS18, CKP14, GKS08, GHSW23, MP16, SL87, AK02, AW09, BK98, BKT97, BOEY03, BDL22, DH07, DS18, EFO22, FM04, FZ15, FP15, GM04, GS14, Haj91, IL93, KR90, Kes95, KA12, KW05, MPSL04, Mit97, MS14, OR96, ZIP95, ZYC16]. Time-communication [BHS18]. Time-dependent [SL87, OR96]. time-equivalent [GM04]. Time-optimal [GHSW23]. Time-space [BER21]. time-stamps [IL93]. time-to-fault [DH07]. Timed [HMN14, San12]. timeliness [ADGT12]. timeouts [CCW95]. Timestamping [GSM07]. timestamps [EFR08, MHN18]. timing [AT96, CW90, LA92]. timing-based [AT96]. TLA [LLR99]. TM [GGT18]. toggle [Pac91]. token [DJV00, FDS95, HC93, HW97]. tolerance [ALK15, BG14, ŢH09]. tolerant [ADS22, ADL90, ABM17, BKA12, DFV20, DDM16, FS12, GGK11, GR07, Jal89, LKJ19, MS14, Par17, Pel91, SS01, ST87, ZYC16]. Tolerating [MHS14, MS03]. top [PSS08]. top- [PSS08]. Topology [KP07, ACR19, CR10, EGS09, GPX07, GKS08, HR13, PPS17, PST15]. Torus [KM02, DS86]. tossing [BGK14]. touching [PPV15]. Trace [Bla86, Jon94]. tracing [DFY05]. track [MS97]. tracking [AG07, ABM02]. Tractable [AHS21]. trade [BER21, CKP14]. trade-offs [BER21, CKP14]. tradeoff [ZYC16]. tradeoffs [BSD88, LLR17, MP16]. traffic [HTW07]. Transaction [CW90, CG21]. transactional [AKS18, BHS18, BHS22, GGT18, GHS10, HS07, IFR20, SCA15, SB12, SB14, ST97, SW22]. transactions [ABF88, IFR20, KGM99, Ruk91]. transduction [JPR99]. transfer [KPPT23, LM91]. Transient [BDT07, MHS14]. transition [CNS21, CMQR23, Hie12, LS92]. transitive [Lis91]. Tree [LM16, AKT14, BDV07, FO23, GM87, KKT12, MK21, Mit97, NdM16, SB95]. tree-depth [NdM16]. trees [BBC^{+}23, CKP14, DN17, FM04, KP08a, KK07, Kor07, KP08b]. trends [AKH03]. tries [Sha19]. true [MdV89]. trust [ACTZ24]. tuning [GM21]. Turing [DFV22]. TuringMobile [DFV22]. two [BAES95, CM10, CPR20, FM03, TV02]. two-dimensional [BAES95]. two-process [TV02]. type [BC95, GK08, Sin91]. types [Had87, LH00].

UDG [EGS09, FP10]. ultrasparse [Pet10]. Unbeatable [CGM22]. unbounded [AGM07, EFR08]. uncontended [CJM17]. Understanding [CP89]. undirected [CIK11, KP05]. unidirectional [HL99]. unified [DNN17]. uniform [FPV17, HR05, HW97, HHH03]. Unifying [AN05]. union [JT21]. Universal [BV02, EFT16, FG97, BPV06, FK20, ZIM15]. universally [ALH^{+}23]. unknown [EGS09, FPP01]. unreliable [AB93, AB19, CHGN14]. unstructured [MW08]. until [HG01]. Updating [ABM17]. Upper [PST16, FK17a, FZ15]. upperbound [vLT87]. use [BNP97, SW22]. useless [CSC11, HMR00]. uses [Lis93]. Using [AM97, CCW95, HM04, LD99, LA92, AHR95, AKW06, ABL13, BG14, BE10, BE13, BBM01, Coa93, CPP02, CIK11, EKS16, ES94, GSM07, GLS12, GHW12, JLC07, JPR99, KP92, Koz12, LM91, MG04, PPV98, PT14, Sin97, TS11, VK20, Moi01].

valency [HS08]. value [AB05, MHS14]. value-oblivious [AB05]. variables [ES94, PPV98]. Vector [BKT97]. vectors [MT07]. Verification [KP92, PZ86, PSL00, TS11, BLT16, Gje88, Gra99, KK07, KKM15, MM00, MS89, Pel91]. verify [LD99]. verifying [AFP20, EJWK21, vGHT16]. version [MT07]. versus [Ciu94, CKP14, MMT08, MP16]. Vertex [Par17, GS14, PP24]. very [MNS17]. via [ALH^{+}23, DS03, Fis20, IFR20, KKT12, MM08, Sli07, Yu11]. views [ACHL20, GMT18]. virtualization [PST15]. visibility [DFV22, KKN^{+}24]. VLSI [ACM86, Mar86]. Vöcking [Att16]. voting [Tan94].

Wait [DGFK15, Hes94, Hes95, HG01, ACR19, AM97, BNP97, EFT16, FRT13, Hav00, HK09, PPV98, PMCP22, TV02]. Wait-free [Hes94, Hes95, HG01, ACR19, AM97, BNP97, FRT13, Hav00, PPV98, PMCP22, TV02]. Wait-freedom [DGFK15, EFT16]. Wake [DGHP23, GM04]. walks [CIK11, GKM07, KKS17]. walls [PW97]. WAN [MMR00]. Wardrop [FOV08]. Weak [HJV15, ADGT08, DNN17, HR99, Sch97a, Sch97b]. weakest [AKN12, DGS19, EHT07, GKK08, GHN09, Zie10]. weakness [FHS06]. weight [FO23, GM87]. weighted [EGN23, FM11, KY11]. which [FK00]. while [VK20, Vid90]. whiteboard [BKT15]. Who [GNZ17]. WiFi [ACY21]. Williams [BE10]. win [GM21]. Window [SB12, PS91]. Window-based [SB12]. Windowed [ACY21]. windows [XTB08]. WinFS [MT07]. wireless [BMIY08, CKV22, HM16, RSZ13]. without [ALK15, CDN19b, DFP04, HIZ21, KPT11, LLR17, PPV15, WS09]. WOOT [KG22]. work [CKL06, HLS06, KKM19]. world [FG14]. worlds [FGP06]. worst [KSW10]. worst-case [KSW10]. WRITE [CPP02, CRR23, Asp12, BRS18, DIM93, FH04, GHW12, GHV01]. Write-All} [GHV01]. writer [And94]. writing [Vid90]. written [KHvB96, KK99].

X [FG01]. X-Ability [FG01]. Xheal [PT14].

References

Aharonson:1995:CNA

[AA95] Eran Aharonson and Hagit Attiya. Counting networks with arbitrary fan-out. Distributed Computing, 8(4):163–169, June 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

            tronic). URL https://link.springer.com/article/10.1007/
            BF02242734.

Aksenov:2023:SLD

[AADM23] Vitaly Aksenov, Dan Alistarh, Alexandra Drozdova, and Amirkeivan Mohtashami. The splay-list: a distribution-adaptive concurrent skip-list. Distributed Computing, 36(3):395–418, September 2023. CODEN DICOEB. ISSN 0178-2770 (print),

            1432-0452 (electronic). URL https://link.springer.com/
            article/10.1007/s00446-022-00441-x.

Angluin:2008:FCP

[AAE08a] Dana Angluin, James Aspnes, and David Eisenstat. Fast computation by population protocols with a leader. Distributed Computing, 21(3):183–199, September 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-008-0067-z.

Angluin:2008:SPP

[AAE08b] Dana Angluin, James Aspnes, and David Eisenstat. A simple population protocol for fast robust approximate majority. Distributed Computing, 21(2):87–102, July 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-008-0059-z.

Arora:1993:CIS

[AAG93] Anish Arora, Paul Attie, and Mohamed Gouda. Convergence of iteration systems. Distributed Computing, 7(1):43–53, November 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02278855.

Abrahamson:1989:RFE

[AAK89] Karl Abrahamson, Andrew Adler, and David Kirkpatrick. Randomized function evaluation on a ring. Distributed Computing, 3 (3):107–117, September 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01784021.

Afek:2013:BMI

[AAK13] Yehuda Afek, Noga Alon, and Fabian Kuhn. Beeping a maximal independent set. Distributed Computing, 26(4):195–208, August 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-012-0175-7.

Angluin:2006:CNP

[AAP06] Dana Angluin, James Aspnes, and René Peralta. Computation in networks of passively mobile finite-state sensors. Distributed Computing, 18(4):235–253, March 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-005-0138-3.

Angluin:2007:CPP

[AAR07] Dana Angluin, James Aspnes, and Eric Ruppert. The computational power of population protocols. Distributed Computing, 20 (4):279–304, November 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-007-0040-2.

Alistarh:2018:CER

[AAS18] Dan Alistarh, James Aspnes, and Jared Saia. Communicationefficient randomized consensus. Distributed Computing, 31(6): 489–501, November 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0315-1.

Afek:1993:SSU

[AB93] Yehuda Afek and Geoffrey M. Brown. Self-stabilization over unreliable communication media. Distributed Computing, 7(1): 27–34, November 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02278853.

Attiya:2002:AEM

[AB02] Hagit Attiya and Vita Bortnikov. Adaptive and efficient mutual exclusion. Distributed Computing, 15(3):177–189, July 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460100068.

Aumann:2005:ELC

[AB05] Yonatan Aumann and Michael A. Bender. Efficient lowcontention asynchronous consensus with the value-oblivious adversary scheduler. Distributed Computing, 17(3):191–207, March 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-004-0113-4.

Almeida:2019:SEC

[AB19] Paulo Sérgio Almeida and Carlos Baquero. Scalable eventually consistent counters over unreliable networks. Distributed Computing, 32(1):69–89, February 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-017-0322-2.

Afek:2002:RPC

[ABBM02] Yehuda Afek, Anat Bremler-Barr, and Michael Merritt. Restoration by path concatenation: fast recovery of MPLS paths. Distributed Computing, 15(4):273–283, December 2002. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
002-0080-6.

Ancilotti:1988:AED

[ABF88] Paolo Ancilotti, Antonia Bertolino, and Mario Fusani. An approach to efficient distributed transactions. Distributed Computing, 2(4):201–212, December 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01872846.

Ackermann:2016:CID

[ABH16] Heiner Ackermann, Petra Berenbrink, and Martin Hoefer. Concurrent imitation dynamics in congestion games. Distributed Computing, 29(2):105–125, April 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0223-6.

Alon:2019:RCH

[ABH19] Noga Alon, Mark Braverman, and Bernhard Haeupler. Reliable communication over highly connected noisy networks. Distributed Computing, 32(6):505–515, December 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-017-0303-5.

Avin:2013:OOI

[ABL13] Chen Avin, Michael Borokhovich, and Zvi Lotker. Order optimal information spreading using algebraic gossip. Distributed Computing, 26(2):99–117, April 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-013-0187-y.

Alvisi:2002:CTC

[ABM02] Lorenzo Alvisi, Karan Bhatia, and Keith Marzullo. Causality tracking in causal message-logging protocols. Distributed Computing, 15(1):1–15, January 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s446-002-8026-7.

Almeida:2017:FTA

[ABM17] Paulo Sérgio Almeida, Carlos Baquero, and Miguel A. Mosteiro. Fault-tolerant aggregation: Flow-Updating meets Mass-Distribution. Distributed Computing, 30(4):281–291, August 2017. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/s00446-016-
0288-5.

Agrawal:1987:DOC

[ABS87] Divyakant Agrawal, Arthur J. Bernstein, and Soumitra Sengupta. Distributed optimistic concurrency control with reduced rollback. Distributed Computing, 2(1):45–59, March 1987. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01786254.

Akbari:2016:SAA

[ABS16] Hoda Akbari, Petra Berenbrink, and Thomas Sauerwald. A simple approach for adapting continuous load balancing processes to discrete settings. Distributed Computing, 29(2):143–161, April 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-016-0266-y.

Antoniadis:2018:EDC

[ABS18] Karolos Antoniadis, Peva Blanchard, and Julien Stainer. The entropy of a distributed computation random number generation from memory interleaving. Distributed Computing, 31(5):389–417, October 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-017-0311-5.

Abraham:2023:CCB

[ACD^{+}23] Ittai Abraham, T.-H. Hubert Chan, Danny Dolev, Kartik Nayak, Rafael Pass, Ling Ren, and Elaine Shi. Communication complexity of Byzantine Agreement, revisited. Distributed Computing, 36(1):3–28, March 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-022-00428-8.

Abboud:2020:FVN

[ACHL20] Amir Abboud, Keren Censor-Hillel, and Christoph Lenzen. Fooling views: a new lower bound technique for distributed computations under congestion. Distributed Computing, 33(6):545–559, December 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-020-00373-4.

Anantharaman:1986:CPE

[ACM86] T. S. Anantharaman, E. M. Clarke, and B. Mishra. Compiling path expressions into VLSI circuits. Distributed Computing, 1(3): 150–166, September 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01661169.

Abraham:2006:BDP

[ACM06] Ittai Abraham, Gregory Chockler, and Dahlia Malkhi. Byzantine Disk Paxos: optimal resilience with Byzantine shared memory. Distributed Computing, 18(5):387–408, April 2006. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0151-6.

Alcantara:2019:TLC

[ACR19] Manuel Alcántara, Armando Castañeda, and Sergio Rajsbaum. The topology of look-compute-move robot wait-free algorithms with hard termination. Distributed Computing, 32(3):235–255, June 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-018-0345-3.

Aguilera:2000:FDC

[ACT00] Marcos Kawazoe Aguilera, Wei Chen, and Sam Toueg. Failure detection and consensus in the crash-recovery model. Distributed Computing, 13(2):99–125, April 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050070.

Alpos:2024:ADT

[ACTZ24] Orestis Alpos, Christian Cachin, Björn Tackmann, and Luca Zanolini. Asymmetric distributed trust. Distributed Computing, 37(3):247–277, September 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-024-00469-1.

Anderton:2021:WBA

[ACY21] William C. Anderton, Trisha Chakraborty, and Maxwell Young. Windowed backoff algorithms for WiFi: theory and performance under batched arrivals. Distributed Computing, 34(5):367–393, October 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-021-00403-9.

Afek:2007:EAC

[AD07] Yehuda Afek and Yaron De Levie. Efficient adaptive collect algorithms. Distributed Computing, 20(3):221–238, October 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-007-0041-1.

Aguilera:2008:IOS

[ADGT08] Marcos K. Aguilera, Carole Delporte-Gallet, and Sam Toueg. On implementing omega in systems with weak reliability and synchrony assumptions. Distributed Computing, 21(4):285–314, October 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-008-0068-y.

Aguilera:2012:PSB

[ADGT12] Marcos K. Aguilera, Carole Delporte-Gallet, and Sam Toueg. Partial synchrony based on set timeliness. Distributed Computing, 25(3):249–260, June 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0158-8.

Ahamad:1990:FTA

[ADL90] Mustaque Ahamad, Partha Dasgupta, and Richard J. LeBlanc, Jr. Fault-tolerant atomic computations in an object-based distributed system. Distributed Computing, 4(2):69–80, June 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01786632.

Abraham:2022:RAF

[ADS22] Ittai Abraham, Danny Dolev, and Gilad Stern. Revisiting asynchronous fault tolerant computation with optimal resilience. Distributed Computing, 35(4):333–355, August 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
021-00416-4.

Avin:2018:BBR

[AE18] Chen Avin and Robert Elsässer. Breaking the log n barrier on rumor spreading. Distributed Computing, 31(6):503–513, November 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-017-0312-4.

Attie:1993:FHM

[AFG93] Paul C. Attie, Nissim Francez, and Orna Grumberg. Fairness and hyperfairness in multi-party interactions. Distributed Computing, 6(4):245–254, July 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02242712.

Attiya:2002:ACA

[AFG02] Hagit Attiya, Arie Fouren, and Eli Gafni. An adaptive collect algorithm with applications. Distributed Computing, 15(2):87–96, April 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460100067.

Apt:1988:AFL

[AFK88] Krzysztof R. Apt, Nissim Francez, and Shmuel Katz. Appraising fairness in languages for distributed programming. Distributed Computing, 2(4):226–241, December 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01872848.

Arfaoui:2020:DVN

[AFP20] Heger Arfaoui, Pierre Fraigniaud, and Andrzej Pelc. Deciding and verifying network properties locally with few output bits. Distributed Computing, 33(2):169–187, April 2020. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
019-00355-1.

Aspnes:2006:RBB

[AFR06] James Aspnes, Faith Ellen Fich, and Eric Ruppert. Relationships between broadcast and shared memory in reliable anonymous distributed systems. Distributed Computing, 18(3):209–219, February 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-005-0145-4.

Albouy:2024:GCE

[AFRT24] Timothé Albouy, Davide Frey, Michel Raynal, and François Taïani. Good-case early-stopping latency of synchronous Byzantine reliable broadcast: the deterministic case. Distributed Computing, 37(2):121–143, June 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-024-00464-6.

Ackermann:2011:DAQ

[AFS11] Heiner Ackermann, Simon Fischer, and Marcel Schöngens. Distributed algorithms for QoS load balancing. Distributed Computing, 23(5–6):321–330, April 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0125-1.

Agarwal:2007:EDT

[AG07] Anurag Agarwal and Vijay K. Garg. Efficient dependency tracking for relevant events in concurrent systems. Distributed Computing, 19(3):163–183, January 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-006-0004-y.

Abdel-Ghaffar:1998:EDD

[AGE98] Khaled A. S. Abdel-Ghaffar and Amr El Abbadi. Efficient detection of discrepancies in multiple file copies. Distributed Computing, 11(2):59–72, April 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050042.

Aguilera:2010:MP

[AGL10] Marcos K. Aguilera, Eli Gafni, and Leslie Lamport. The mailbox problem. Distributed Computing, 23(2):113–134, October 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-010-0102-8.

Afek:2007:CES

[AGM07] Yehuda Afek, Eli Gafni, and Adam Morrison. Common2 extended to stacks and unbounded concurrency. Distributed Computing, 20 (4):239–252, November 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-007-0023-3.

Afek:2024:DCC

[AGPS24] Yehuda Afek, Gal Giladi, and Boaz Patt-Shamir. Distributed computing with the cloud. Distributed Computing, 37(1):1–18, March 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-024-00460-w.

Afek:2010:SCP

[AGT10] Yehuda Afek, Eli Gafni, and Corentin Travers. The ksimultaneous consensus problem. Distributed Computing, 22(3): 185–195, March 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-009-0090-8.

Aguilera:2014:SIS

[Agu14] Marcos K. Aguilera. Special issue with selected papers from DISC 2012. Distributed Computing, 27(6):391, December 2014. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0235-2.

Afek:1995:CGC

[AH95] Y. Afek and D. Hendler. On the complexity of global computation in the presence of link failures: the general case. Distributed Computing, 8(3):115–120, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF02242713.

Afek:2013:FSR

[AHM13] Yehuda Afek, Michael Hakimi, and Adam Morrison. Fast and scalable rendezvousing. Distributed Computing, 26(4):243–269, August 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-013-0185-0.

Attiya:1995:ASU

[AHR95] Hagit Attiya, Maurice Herlihy, and Ophir Rachman. Atomic snapshots using lattice agreement. Distributed Computing, 8 (3):121–132, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02242714.

Anta:2021:TLD

[AHS21] Antonio Fernández Anta, Theophanis Hadjistasi, and Alexander A. Schwarzmann. Tractable low-delay atomic memory. Distributed Computing, 34(1):33–58, February 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
020-00379-y.

Abraham:2023:RCA

[AJM^{+}23] Ittai Abraham, Philipp Jovanovic, Mary Maller, Sarah Meiklejohn, Gilad Stern, and Alin Tomescu. Reaching consensus for asynchronous distributed key generation. Distributed Computing, 36(3):219–252, September 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-022-00436-8.

Anderson:2001:NFP

[AK01] James H. Anderson and Yong-Jik Kim. A new fast-path mechanism for mutual exclusion. Distributed Computing, 14(1):17–29, January 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/PL00008923.

Anderson:2002:ILB

[AK02] James H. Anderson and Yong-Jik Kim. An improved lower bound for the time complexity of mutual exclusion. Distributed Computing, 15(4):221–253, December 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-002-0084-2.

Awerbuch:2009:GDO

[AK09] Baruch Awerbuch and Rohit Khandekar. Greedy distributed optimization of multi-commodity flows. Distributed Computing, 21 (5):317–329, February 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0074-0.

Anderson:2003:SMM

[AKH03] James H. Anderson, Yong-Jik Kim, and Ted Herman. Sharedmemory mutual exclusion: major research trends since 1986. Distributed Computing, 16(2–3):75–110, September 2003. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
003-0088-6.

Afek:2012:RWF

[AKN12] Yehuda Afek, Petr Kuznetsov, and Israel Nir. Renaming and the weakest family of failure detectors. Distributed Computing, 25(6): 411–425, December 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0177-5.

Alistarh:2018:ILH

[AKS18] Dan Alistarh, Justin Kopinsky, and Nir Shavit. Inherent limitations of hybrid transactional memory. Distributed Computing, 31(3):167–185, June 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0305-3.

Afek:2014:CTP

[AKT14] Yehuda Afek, Haim Kaplan, and Robert E. Tarjan. The CB tree: a practical concurrent self-adjusting search tree. Distributed Computing, 27(6):393–417, December 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0229-0.

Aminof:2018:PMC

[AKV18] Benjamin Aminof, Tomer Kotek, and Helmut Veith. Parameterized model checking of rendezvous systems. Distributed Computing, 31(3):187–222, June 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-017-0302-6.

Attiya:2006:EAC

[AKW06] Hagit Attiya, Fabian Kuhn, and Roger Wattenhofer. Efficient adaptive collect using randomization. Distributed Computing, 18 (3):179–188, February 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0143-6.

Anagnostides:2023:AUO

[ALH^{+}23] Ioannis Anagnostides, Christoph Lenzen, Bernhard Haeupler, Goran Zuzic, and Themis Gouleakis. Almost universally optimal distributed Laplacian solvers via low-congestion shortcuts.

Distributed Computing, 36(4):475–499, December 2023. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
023-00454-0.

Abujarad:2015:CAA

[ALK15] Fuad Abujarad, Yiyan Lin, and Sandeep S. Kulkarni. The complexity of automated addition of fault-tolerance without explicit legitimate states. Distributed Computing, 28(3):201–219, June 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-014-0227-2.

Alvisi:2011:SIP

[Alv11] Lorenzo Alvisi. Special issue on PODC 2009. Distributed Computing, 24(1):??, September 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0136-6.

Anderson:1997:ULS

[AM97] James H. Anderson and Mark Moir. Using local-spin kexclusion algorithms to improve wait-free object implementations. Distributed Computing, 11(1):1–20, December 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050039.

Abraham:2005:PQD

[AM05] Ittai Abraham and Dahlia Malkhi. Probabilistic quorums for dynamic systems. Distributed Computing, 18(2):113–124, November 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-005-0139-2.

Avin:2020:DAN

[AMS20] Chen Avin, Kaushik Mondal, and Stefan Schmid. Demand-aware network designs of bounded degree. Distributed Computing, 33 (3–4):311–325, June 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-019-00351-5.

Alon:2005:TBS

[AMW05] Noga Alon, Michael Merritt, and Rebecca N. Wright. Tight bounds for shared memory systems accessed by Byzantine processes. Distributed Computing, 18(2):99–109, November 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-005-0125-8.

Anta:2012:OID

[AMZ12] Antonio Fernández Anta, Alessia Milani, and Shmuel Zaks. Opportunistic information dissemination in mobile ad-hoc networks: the profit of global synchrony. Distributed Computing, 25(4): 279–296, August 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0165-9.

Arora:2005:UST

[AN05] Anish Arora and Mikhail Nesterenko. Unifying stabilization and termination in message-passing systems. Distributed Computing, 17(3):279–290, March 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-004-0111-6.

Anderson:1993:CR

[And93] James H. Anderson. Composite registers. Distributed Computing, 6(3):141–154, April 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02242703.

Anderson:1994:MWC

[And94] James H. Anderson. Multi-writer composite registers. Distributed Computing, 7(4):175–195, May 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF02280833.

Anderson:2001:I

[And01] James H. Anderson. Introduction. Distributed Computing, 14 (4):191, December 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460100066.

Ahamad:1995:CMD

[ANH95] Mustaque Ahamad, Gil Neiger, and Phillip W. Hutto. Causal memory: definitions, implementation, and programming. Distributed Computing, 9(1):37–49, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01784241.

Anonymous:1986:CE

[Ano86] Anonymous. Calendar of events. Distributed Computing, 1(1):73– 74, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01843571.

Anonymous:2000:ETC

[Ano00] Anonymous. Editorial: Time for change. Distributed Computing, 13(4):187, November 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460000043.

Anonymous:2021:RA

[Ano21] Anonymous. Reviewer acknowledgement (2020). Distributed Computing, 34(3):179, June 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00395-6.

Andrews:1986:ESL

[AO86] Gregory R. Andrews and Ronald A. Olsson. The evolution of the SR language. Distributed Computing, 1(3):133–149, September 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01661168.

Alnuweiri:1991:OGA

[AP91] Hussein M. Alnuweiri and Viktor K. Prasanna. Optimal geometric algorithms for digitized images on fixed-size linear arrays and scan-line arrays. Distributed Computing, 5(2):55–65, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02259747.

Attiya:1997:LHR

[AR97] Hagit Attiya and Rinat Rappoport. The level of handshake required for managing a connection. Distributed Computing, 11 (1):41–57, December 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050041.

Attiya:2003:I

[AR03] Hagit Attiya and Sergio Rajsbaum. Introduction. Distributed Computing, 16(2–3):69–70, September 2003. CODEN DICOEB.

ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-003-0100-1.

Arora:2007:I

[Aro07] Anish Arora. Introduction. Distributed Computing, 20(1):1, July 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-007-0036-y.

Alpern:1987:RSL

[AS87] Bowen Alpern and Fred B. Schneider. Recognizing safety and liveness. Distributed Computing, 2(3):117–126, September 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01782772.

Aspnes:2003:RPA

[Asp03] James Aspnes. Randomized protocols for asynchronous consensus. Distributed Computing, 16(2–3):165–175, September 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-002-0081-5.

Aspnes:2007:E

[Asp07] James Aspnes. Editorial. Distributed Computing, 19(4):253, March 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-006-0016-7.

Aspnes:2012:MAS

[Asp12] James Aspnes. A modular approach to shared-memory consensus, with applications to the probabilistic-write model. Distributed Computing, 25(2):179–188, May 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0134-8.

Aspnes:2015:FRC

[Asp15] James Aspnes. Faster randomized consensus with an oblivious adversary. Distributed Computing, 28(1):21–29, February 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0195-y.

Afek:2002:LLA

[AST02] Yehuda Afek, Gideon Stupp, and Dan Touitou. Long lived adaptive splitter and applications. Distributed Computing, 15(2):67– 86, April 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460100060.

Alur:1996:FTB

[AT96] Rajeev Alur and Gadi Taubenfeld. Fast timing-based algorithms. Distributed Computing, 10(1):1–10, July 1996. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s004460050020.

Aguilera:2010:APG

[AT10] Marcos K. Aguilera and Sam Toueg. Adaptive progress: a gracefully-degrading liveness property. Distributed Computing, 22 (5–6):303–334, August 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-010-0106-4.

Aguilera:2012:CPB

[AT12] Marcos K. Aguilera and Sam Toueg. The correctness proof of Ben-Or’s randomized consensus algorithm. Distributed Computing, 25 (5):371–381, October 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0162-z.

Attiya:2009:EAA

[Att09] Hagit Attiya. Editorial: It’s all about change. Distributed Computing, 22(1):1, April 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-009-0085-5.

Attiya:2012:ABR

[Att12] Hagit Attiya. Announcement: best reviewer award 2011. Distributed Computing, 25(1):1, March 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-012-0161-0.

Attiya:2016:SIM

[Att16] Hagit Attiya. Special issue in memory of Berthold Vöcking. Distributed Computing, 29(2):75, April 2016. CODEN DICOEB.

ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-016-0268-9.

Attiya:2017:SID

[Att17] Hagit Attiya. Special issue on DISC 2013, 2014 and PODC 2014. Distributed Computing, 30(5):307, October 2017. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
017-0314-2.

Attiya:2019:SIP

[Att19] Hagit Attiya. Special issue on PODC 2015 and PODC 2016. Distributed Computing, 32(6):459, December 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
019-00366-y.

Attiya:2020:ESI

[Att20] Hagit Attiya. Editorial: Special issue of PODC 2017 and DISC 2017. Distributed Computing, 33(3–4):207, June 2020. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
020-00374-3.

Attiya:2021:SIPa

[Att21a] Hagit Attiya. Special issue on PODC 2018 and DISC 2018. Distributed Computing, 34(4):227, August 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-021-00400-y.

Attiya:2021:SIPb

[Att21b] Hagit Attiya. Special issue on PODC 2019. Distributed Computing, 34(6):411, December 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00411-9.

Attiya:2022:SID

[Att22a] Hagit Attiya. Special issue on DISC 2020. Distributed Computing, 35(6):475, December 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-022-00438-6.

Attiya:2022:SIP

[Att22b] Hagit Attiya. Special issue on PODC 2020. Distributed Computing, 35(4):303, August 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-022-00434-w.

Attiya:2023:SID

[Att23] Hagit Attiya. Special issue on DISC 2019. Distributed Computing, 36(1):1, March 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-023-00445-1.

Attiya:2024:SIP

[Att24] Hagit Attiya. Special issue on PODC 2021 and DISC 2021. Distributed Computing, 37(1):65, March 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-023-00456-y.

Aspnes:2009:EMT

[AW09] James Aspnes and Udi Wieder. The expansion and mixing time of skip graphs with applications. Distributed Computing, 21 (6):385–393, March 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0071-3.

Amdur:1992:MCB

[AWH92] Eugene S. Amdur, Samuel M. Weber, and Vassos Hadzilacos. On the message complexity of binary Byzantine agreement under crash failures. Distributed Computing, 5(4):175–186, April 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02277665.

Ben-Aroya:1995:GHP

[BAES95] Ishai Ben-Aroya, Tamar Eilam, and Assaf Schuster. Greedy hotpotato routing on the two-dimensional mesh. Distributed Computing, 9(1):3–19, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01784239.

Bazzi:2000:SBQ

[Baz00] Rida A. Bazzi. Synchronous Byzantine quorum systems. Distributed Computing, 13(1):45–52, January 2000. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050004.

Bazzi:2001:ACA

[Baz01] Rida A. Bazzi. Access cost for asynchronous Byzantine quorum systems. Distributed Computing, 14(1):41–48, January 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008925.

Balliu:2023:LCP

[BBC^{+}23] Alkida Balliu, Sebastian Brandt, Yi-Jun Chang, Dennis Olivetti, Jan Studený, Jukka Suomela, and Aleksandr Tereshchenko. Locally checkable problems in rooted trees. Distributed Computing, 36(3):277–311, September 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-022-00435-9.

Ben-Basat:2023:ODC

[BBEiKS23] Ran Ben-Basat, Guy Even, Ken ichi Kawarabayashi, and Gregory Schwartzman. Optimal distributed covering algorithms. Distributed Computing, 36(1):45–55, March 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-021-00391-w.

Beauquier:2001:PCS

[BBM01] J. Beauquier, B. Bérard, and F. Magniette. Proving convergence of self-stabilizing systems using first-order rewriting and regular languages. Distributed Computing, 14(2):83–95, April 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008931.

Bouchard:2019:AAP

[BBP19] Sébastien Bouchard, Marjorie Bournat, and Franck Petit. Asynchronous approach in the plane: a deterministic polynomial algorithm. Distributed Computing, 32(4):317–337, August 2019. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
018-0338-2.

Bonakdarpour:2012:FAD

[BBS12] Borzoo Bonakdarpour, Marius Bozga, and Joseph Sifakis. A framework for automated distributed implementation of component-based models. Distributed Computing, 25(5):383–409, October 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-012-0168-6.

Balliu:2021:AGP

[BBS21] Alkida Balliu, Sebastian Brandt, and Jukka Suomela. Almost global problems in the LOCAL model. Distributed Computing, 34 (4):259–281, August 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-020-00375-2.

Baldoni:1995:CHP

[BC95] Roberto Baldoni and Bruno Ciciani. A class of high performance Maekawa-type algorithms for distributed systems under heavy demand. Distributed Computing, 8(4):171–180, June 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02242735.

Bose:2013:BLD

[BCD13] Prosenjit Bose, Paz Carmi, and Stephane Durocher. Bounding the locality of distributed routing algorithms. Distributed Computing, 26(1):39–58, February 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0179-3.

Baumann:2011:PFD

[BCF11] Hervé Baumann, Pierluigi Crescenzi, and Pierre Fraigniaud. Parsimonious flooding in dynamic graphs. Distributed Computing, 24 (1):31–44, September 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-011-0133-9.

Bravo:2022:MBC

[BCG22] Manuel Bravo, Gregory Chockler, and Alexey Gotsman. Making Byzantine consensus live. Distributed Computing, 35(6):503–532, December 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-022-00432-y.

Bravo:2024:LLB

[BCG24] Manuel Bravo, Gregory Chockler, and Alexey Gotsman. Liveness and latency of Byzantine state-machine replication. Distributed Computing, 37(2):177–205, June 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-024-00466-4.

Becchetti:2019:SSR

[BCP19] L. Becchetti, A. Clementi, and G. Posta. Self-stabilizing repeated balls-into-bins. Distributed Computing, 32(1):59–68, February 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-017-0320-4.

Becchetti:2017:SDP

[BCT17] Luca Becchetti, Andrea Clementi, and Luca Trevisan. Simple dynamics for plurality consensus. Distributed Computing, 30 (4):293–306, August 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-016-0289-4.

Bruschi:1997:LBB

[BD97] Danilo Bruschi and Massimiliano Del Pinto. Lower bounds for the broadcast problem in mobile radio networks. Distributed Computing, 10(3):129–135, March 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050030.

Bouchard:2016:BGN

[BDD16] Sébastien Bouchard, Yoann Dieudonné, and Bertrand Ducourthial. Byzantine gathering in networks. Distributed Computing, 29(6): 435–457, November 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-016-0276-9.

Borowiecki:2015:DGS

[BDK15] Piotr Borowiecki, Dariusz Dereniowski, and Łukasz Kuszner. Distributed graph searching with a sense of direction. Distributed Computing, 28(3):155–170, June 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0236-1.

Bouchard:2022:BGP

[BDL22] Sébastien Bouchard, Yoann Dieudonné, and Anissa Lamani. Byzantine gathering in polynomial time. Distributed Computing, 35(3):235–263, June 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-022-00419-9.

Beauquier:2007:TFD

[BDT07] Joffroy Beauquier, Sylvie Delaët, and Sébastien Tixeuil. Transient fault detectors. Distributed Computing, 20(1):39–51, July 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-007-0029-x.

Bui:2007:SSP

[BDV07] Alain Bui, Ajoy K. Datta, and Vincent Villain. Snap-stabilization and PIF in tree networks. Distributed Computing, 20(1):3–19, July 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-007-0030-4.

Barenboim:2010:SDM

[BE10] Leonid Barenboim and Michael Elkin. Sublogarithmic distributed MIS algorithm for sparse graphs using Nash-Williams decomposition. Distributed Computing, 22(5–6):363–379, August 2010. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
009-0088-2.

Barenboim:2013:DDE

[BE13] Leonid Barenboim and Michael Elkin. Distributed deterministic edge coloring using bounded neighborhood independence. Distributed Computing, 26(5–6):273–287, October 2013. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
012-0167-7.

Barenboim:2014:CAD

[BE14] Leonid Barenboim and Michael Elkin. Combinatorial algorithms for distributed graph coloring. Distributed Computing, 27(2):79– 93, April 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-013-0203-2.

Berenbrink:2016:ERB

[BEF16] Petra Berenbrink, Robert Elsässer, and Tom Friedetzky. Efficient randomised broadcasting in random regular networks with applications in peer-to-peer systems. Distributed Computing, 29 (5):317–339, October 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-016-0264-0.

Beimel:2007:PCI

[Bei07] Amos Beimel. On private computation in incomplete networks. Distributed Computing, 19(3):237–252, January 2007. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
006-0010-0.

Berenbrink:2021:TST

[BER21] Petra Berenbrink, Robert Elsässer, and Tomasz Radzik. Timespace trade-offs in population protocols for the majority problem. Distributed Computing, 34(2):91–111, April 2021. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
020-00385-0.

Brodsky:2011:FAA

[BEW11] Alex Brodsky, Faith Ellen, and Philipp Woelfel. Fully-adaptive algorithms for long-lived renaming. Distributed Computing, 24 (2):??, October 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-011-0137-5.

Bilo:2021:CAN

[BFM21] Vittorio Bilò, Michele Flammini, and Luca Moscardelli. Computing approximate Nash equilibria in network congestion games with polynomially decreasing cost functions. Distributed Computing, 34(1):1–14, February 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-020-00381-4.

Bernstein:1987:PTC

[BG87] Philip A. Bernstein and Nathan Goodman. A proof technique for concurrency control and recovery algorithms for replicated databases. Distributed Computing, 2(1):32–44, March 1987.

CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01786253.

Berman:1993:FCN

[BG93] Piotr Berman and Juan A. Garay. Fast consensus in networks of bounded degree. Distributed Computing, 7(2):67–73, December 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02280836.

Balasubramanian:2014:FTD

[BG14] Bharath Balasubramanian and Vijay K. Garg. Fault tolerance in distributed systems using fused state machines. Distributed Computing, 27(4):287–311, August 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0209-4.

Beauquier:2007:RSS

[BGJ07] Joffroy Beauquier, Maria Gradinariu, and Colette Johnen. Randomized self-stabilizing and space optimal leader election under arbitrary scheduler on rings. Distributed Computing, 20(1):75–93, July 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-007-0034-0.

Boyle:2014:LRC

[BGK14] Elette Boyle, Shafi Goldwasser, and Yael Tauman Kalai. Leakageresilient coin tossing. Distributed Computing, 27(3):147–164, June 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-013-0206-z.

Balliu:2023:NEA

[BGKO23] Alkida Balliu, Mohsen Ghaffari, Fabian Kuhn, and Dennis Olivetti. Node and edge averaged complexities of local graph problems. Distributed Computing, 36(4):451–473, December 2023. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
023-00453-1.

Barbara:1986:MEP

[BGM86] Daniel Barbara and Hector Garcia-Molina. Mutual exclusion in partitioned distributed systems. Distributed Computing, 1(2):119– 132, June 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01786230.

Burns:1993:SPS

[BGM93] James E. Burns, Mohamed G. Gouda, and Raymond E. Miller. Stabilization and pseudo-stabilization. Distributed Computing, 7 (1):35–42, November 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02278854.

Borowsky:2001:BDS

[BGR01] E. Borowsky, E. Gafni, and S. Rajsbaum. The BG distributed simulation algorithm. Distributed Computing, 14(3):127–146, July 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008933.

Berenbrink:2024:AOD

[BHK^{+}24] Petra Berenbrink, Martin Hoefer, Dominik Kaaser, Pascal Lenzner, Malin Rau, and Daniel Schmand. Asynchronous opinion dynamics in social networks. Distributed Computing, 37(3): 207–224, September 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-024-00467-3.

Baig:2023:LLC

[BHMT23] Mirza Ahad Baig, Danny Hendler, Alessia Milani, and Corentin Travers. Long-lived counters with polylogarithmic amortized step complexity. Distributed Computing, 36(1):29–43, March 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-022-00439-5.

Busch:2018:TCI

[BHS18] Costas Busch, Maurice Herlihy, and Gokarna Sharma. Timecommunication impossibility results for distributed transactional memory. Distributed Computing, 31(6):471–487, November 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-017-0318-y.

Busch:2022:DSD

[BHS22] Costas Busch, Maurice Herlihy, and Gokarna Sharma. Dynamic scheduling in distributed transactional memory. Distributed Computing, 35(1):19–36, February 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00410-w.

Baier:1998:MCP

[BK98] Christel Baier and Marta Kwiatkowska. Model checking for a probabilistic branching time logic with fairness. Distributed Computing, 11(3):125–155, August 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050046.

Bazzi:2007:EDI

[BK07] Rida A. Bazzi and Goran Konjevod. On the establishment of distinct identities in overlay networks. Distributed Computing, 19 (4):267–287, March 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0012-y.

Bonakdarpour:2012:SSM

[BKA12] Borzoo Bonakdarpour, Sandeep S. Kulkarni, and Fuad Abujarad. Symbolic synthesis of masking fault-tolerant distributed programs. Distributed Computing, 25(1):83–108, March 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-011-0139-3.

Bienkowski:2016:RME

[BKK16] Marcin Bienkowski, Marek Klonowski, and Dariusz R. Kowalski. Randomized mutual exclusion on a multiple access channel. Distributed Computing, 29(5):341–359, October 2016. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
016-0265-z.

Brunekreef:1996:DAD

[BKM96] Jacob Brunekreef, Joost-Pieter Katoen, and Sjouke Mauw. Design and analysis of dynamic leader election protocols in broadcast networks. Distributed Computing, 9(4):??, February 1996.

CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050017.

Boczkowski:2019:MMS

[BKN19] Lucas Boczkowski, Amos Korman, and Emanuele Natale. Minimizing message size in stochastic communication patterns: fast self-stabilizing protocols with 3 bits. Distributed Computing, 32 (3):173–191, June 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-018-0330-x.

Back:1989:DPN

[BKS89] R.-J. R. Back and R. Kurki-Suonio. Decentralization of process nets with centralized control. Distributed Computing, 3(2):73–87, June 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01558665.

Basten:1997:VTC

[BKT97] Twan Basten, Thomas Kunz, and David J. Taylor. Vector time and causality among abstract events in distributed computations. Distributed Computing, 11(1):21–39, December 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050040.

Becker:2015:AEN

[BKT15] Florent Becker, Adrian Kosowski, and Ioan Todinca. Allowing each node to communicate only once in a distributed system: shared whiteboard models. Distributed Computing, 28(3):189– 200, June 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-014-0221-8.

Black:1986:EDI

[Bla86] David L. Black. On the existence of delay-insensitive fair arbiters: Trace theory and its limitations. Distributed Computing, 1(4): 205–225, December 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01660033.

Berard:2016:FVM

[BLT16] Béatrice Bérard, Pascal Lafourcade, and Sébastien Tixeuil. Formal verification of mobile robot protocols. Distributed Computing, 29(6):459–487, November 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-016-0271-1.

Beimel:2005:ERC

[BM05] Amos Beimel and Lior Malka. Efficient reliable communication over partially authenticated networks. Distributed Computing, 18(1):1–19, July 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-004-0119-y.

Busch:2008:CFM

[BMIY08] Costas Busch, Malik Magdon-Ismail, and Bülent Yener. Contentionfree MAC protocols for asynchronous wireless sensor networks. Distributed Computing, 21(1):23–42, June 2008. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/s00446-007-
0053-x.

Baldoni:2006:OPB

[BMP06] Roberto Baldoni, Alessia Milani, and Sara Tucci Piergiovanni. Optimal propagation-based protocols implementing causal memories. Distributed Computing, 18(6):461–474, June 2006. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0128-5.

Burns:1994:FSD

[BN94] James E. Burns and Gil Neiger. Fast and simple distributed consensus. Distributed Computing, 8(2):59–64, October 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02280828.

Bar-Noy:1989:CCL

[BNBOD89] Amotz Bar-Noy, Michael Ben-Or, and Danny Dolev. Choice coordination with limited failure. Distributed Computing, 3(2):61–72, June 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01558664.

Bar-Noy:1991:CAO

[BND91] Amotz Bar-Noy and Danny Dolev. Consensus algorithms with one-bit messages. Distributed Computing, 4(3):105–110, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452

(electronic). URL https://link.springer.com/article/10.
1007/BF01798957.

Bar-Noy:1990:OBA

[BNNN90] Amotz Bar-Noy, Joseph Naor, and Moni Naor. One-bit algorithms. Distributed Computing, 4(1):3–8, March 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01783661.

Bar-Noy:2002:NOP

[BNNPS02] Amotz Bar-Noy, Aviv Nisgav, and Boaz Patt-Shamir. Nearly optimal perfectly periodic schedules. Distributed Computing, 15 (4):207–220, December 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-002-0085-1.

Bazzi:1997:URA

[BNP97] Rida A. Bazzi, Gil Neiger, and Gary L. Peterson. On the use of registers in achieving wait-free consensus. Distributed Computing, 10(3):117–127, March 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050029.

Barenboim:2022:DBP

[BO22] Leonid Barenboim and Gal Oren. Distributed backup placement. Distributed Computing, 35(5):455–473, October 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
022-00423-z.

Ben-Or:2003:ROI

[BOEY03] Michael Ben-Or and Ran El-Yaniv. Resilient-optimal interactive consistency in constant time. Distributed Computing, 16(4): 249–262, December 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-002-0083-3.

Brakerski:2011:DDL

[BPS11] Zvika Brakerski and Boaz Patt-Shamir. Distributed discovery of large near-cliques. Distributed Computing, 24(2):79–89, October 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-011-0132-x.

Buhrman:2006:IHI

[BPV06] Harry Buhrman, Alessandro Panconesi, and Paul Vitanyi. On the importance of having an identity or, is consensus really universal? Distributed Computing, 18(3):167–176, February 2006. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0121-z.

Brahmadathan:1991:AGP

[BR91a] K. Brahmadathan and K. V. S. Ramarao. Achieving graceful performance in distributed error-prone databases. Distributed Computing, 4(4):163–174, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01784718.

Brookes:1991:DAN

[BR91b] S. D. Brookes and A. W. Roscoe. Deadlock analysis in networks of communicating processes. Distributed Computing, 4(4):209–230, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01784721.

Bonnet:2013:AAS

[BR13] François Bonnet and Michel Raynal. Anonymous asynchronous systems: the case of failure detectors. Distributed Computing, 26(3):141–158, June 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0169-5.

Brinksma:1999:CCD

[Bri99] Ed Brinksma. Cache consistency by design. Distributed Computing, 12(2–3):61–74, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050058.

Broy:1987:SFI

[Bro87] Manfred Broy. Semantics of finite and infinite networks of concurrent communicating agents. Distributed Computing, 2(1):13–31, March 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01786252.

Brown:1990:AM

[Bro90] Geoffrey M. Brown. Asynchronous multicaches. Distributed Computing, 4(1):31–36, March 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01783663.

Broy:1992:AFS

[Bro92] Manfred Broy. Algebraic and functional specification of an interactive serializable database interface. Distributed Computing, 6(1):5–18, July 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02276638.

Bouzid:2018:AOF

[BRS18] Zohir Bouzid, Michel Raynal, and Pierre Sutra. Anonymous obstruction-free (n, k)-set agreement with n − k + 1 atomic read/ write registers. Distributed Computing, 31(2):99–117, April 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-017-0301-7.

Babaoglu:1988:RBC

[BSD88] Özalp Babaoglu, Pat Stephenson, and Rogério Drummond. Reliable broadcasts and communication models: tradeoffs and lower bounds. Distributed Computing, 2(4):177–189, December 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01872844.

Bracha:1987:DDD

[BT87] Gabriel Bracha and Sam Toueg. Distributed deadlock detection. Distributed Computing, 2(3):127–138, September 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01782773.

Boukerche:1995:PSH

[BT95] Azzedine Boukerche and Carl Tropper. Parallel simulation on the hypercube multiprocessor. Distributed Computing, 8(4):181–190, June 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02242736.

Blin:2018:CDS

[BT18] Lélia Blin and Sébastien Tixeuil. Compact deterministic selfstabilizing leader election on a ring: the exponential advantage of being talkative. Distributed Computing, 31(2):139–166, April 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-017-0294-2.

Brandt:2020:TLB

[BUW20] Sebastian Brandt, Jara Uitto, and Roger Wattenhofer. A tight lower bound for semi-synchronous collaborative grid exploration. Distributed Computing, 33(6):471–484, December 2020. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
020-00369-0.

Boldi:2002:UDS

[BV02] Paolo Boldi and Sebastiano Vigna. Universal dynamic synchronous self–stabilization. Distributed Computing, 15(3):137– 153, July 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460100062.

Boldi:2003:LBS

[BV03] Paolo Boldi and Sebastiano Vigna. Lower bounds for sense of direction in regular graphs. Distributed Computing, 16(4):279–286, December 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-003-0092-x.

Bar-Yehuda:1991:EES

[BYGI91] Reuven Bar-Yehuda, Oded Goldreich, and Alon Itai. Efficient emulation of single-hop radio network with collision detection on multi-hop radio network with no collision detection. Distributed Computing, 5(2):67–71, September 1991. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02259748.

Cazzola:2003:RMI

[Caz03] Walter Cazzola. Remote method invocation as a first-class citizen. Distributed Computing, 16(4):287–306, December 2003. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
003-0094-8.

Coulson:2001:ICG

[CB01] Geoff Coulson and Shakuntala Baichoo. Implementing the CORBA GIOP in a high-performance object request broker environment. Distributed Computing, 14(2):113–126, April 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008928.

Charron-Bost:2021:MAS

[CBM21] Bernadette Charron-Bost and Shlomo Moran. MinMax algorithms for stabilizing consensus. Distributed Computing, 34 (3):195–206, June 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-021-00392-9.

Charron-Bost:1996:SAC

[CBMT96] Bernadette Charron-Bost, Friedemann Mattern, and Gerard Tel. Synchronous, asynchronous, and causally ordered communication. Distributed Computing, 9(4):173–191, February 1996. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050018.

Coulson:2002:DCR

[CBP02] Geoff Coulson, Gordon S. Blair, and Nikos Parlavantzas. The design of a configurable and reconfigurable middleware platform. Distributed Computing, 15(2):109–126, April 2002. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460100064.

Charron-Bost:2009:HMC

[CBS09] Bernadette Charron-Bost and André Schiper. The Heard-Of model: computing in distributed systems with benign faults. Distributed Computing, 22(1):49–71, April 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-009-0084-6.

Chan:1988:DEC

[CC88] M. Y. Chan and F. Y. L. Chin. Distributed election in complete networks. Distributed Computing, 3(1):19–22, March 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01788564.

Celikkan:1995:GDI

[CC95] Ufuk Celikkan and Rance Cleaveland. Generating diagnostic information for behavioral preorders. Distributed Computing, 9 (2):61–75, October 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050010.

Chen:2017:SFC

[CCS17] Ho-Lin Chen, Rachel Cummings, and David Soloveichik. Speed faults in computation by chemical reaction networks. Distributed Computing, 30(5):373–390, October 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-015-0255-6.

Cazzola:2022:PPT

[CCT22] Walter Cazzola, Francesco Cesarini, and Luca Tansini. PerformERL: a performance testing framework for Erlang. Distributed Computing, 35(5):439–454, October 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
022-00429-7.

Chaudhuri:1995:UAT

[CCW95] Soma Chaudhuri, Brian A. Coan, and Jennifer L. Welch. Using adaptive timeouts to achieve at-most-once message delivery. Distributed Computing, 9(3):109–117, December 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050013.

Civit:2024:BCT

[CDG^{+}24] Pierre Civit, Muhammad Ayaz Dzulfikar, Seth Gilbert, Vincent Gramoli, Rachid Guerraoui, Jovan Komatovic, and Manuel Vidigueira. Byzantine consensus is Θ(n^{2}): the Dolev–Reischuk bound is tight even in partial synchrony! Distributed Computing, 37(2):89–119, June 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-023-00458-w.

Chockler:2008:CCD

[CDN08] Gregory Chockler, Murat Demirbas, and Tina Nolte. Consensus and collision detectors in radio networks. Distributed Computing, 21(1):55–84, June 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0056-2.

Cicerone:2018:GRM

[CDN18] Serafino Cicerone, Gabriele Di Stefano, and Alfredo Navarra. Gathering of robots on meeting-points: feasibility and optimal resolution algorithms. Distributed Computing, 31(1):1–50, February 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452

(electronic). URL https://link.springer.com/article/10.
1007/s00446-017-0293-3.

Cicerone:2019:AAP

[CDN19a] Serafino Cicerone, Gabriele Di Stefano, and Alfredo Navarra. Asynchronous Arbitrary Pattern Formation: the effects of a rigorous approach. Distributed Computing, 32(2):91–132, April 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-018-0325-7.

Cicerone:2019:EPF

[CDN19b] Serafino Cicerone, Gabriele Di Stefano, and Alfredo Navarra. Embedded pattern formation by asynchronous robots without chirality. Distributed Computing, 32(4):291–315, August 2019. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
018-0333-7.

Chalopin:2016:RNS

[CDP16] Jérémie Chalopin, Yoann Dieudonné, and Andrzej Pelc. Rendezvous in networks in spite of delay faults. Distributed Computing, 29(3):187–205, June 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-015-0259-2.

Czumaj:2024:CSL

[CDPP24] Artur Czumaj, Peter Davies-Peck, and Merav Parter. Component stability in low-space massively parallel computation. Distributed Computing, 37(1):35–64, March 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-024-00461-9.

Chlebus:2001:PTS

[CDS01] Bogdan S. Chlebus, Roberto De Prisco, and Alex A. Shvartsman. Performing tasks on synchronous restartable messagepassing processors. Distributed Computing, 14(1):49–64, January 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008926.

Czerner:2023:LBS

[CEL23] Philipp Czerner, Javier Esparza, and Jérôme Leroux. Lower bounds on the state complexity of population protocols. Distributed Computing, 36(3):209–218, September 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
023-00450-4.

Czumaj:2013:FMD

[CEW13] Artur Czumaj, Robert Elsässer, and Xin Wang. Fast message dissemination in random geometric networks. Distributed Computing, 26(1):1–24, February 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0154-4.

Cho:2022:DCC

[CFH^{+}22] Da-Jung Cho, Matthias Függer, Corbin Hopper, Manish Kushwaha, Thomas Nowak, and Quentin Soubeyran. Distributed computation with continual population growth. Distributed Computing, 35(6):547–569, December 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00404-8.

Chase:1998:DGP

[CG98] Craig M. Chase and Vijay K. Garg. Detection of global predicates: Techniques and their limitations. Distributed Computing, 11(4):191–201, October 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050049.

Chockler:2021:MSD

[CG21] Gregory Chockler and Alexey Gotsman. Multi-shot distributed transaction commit. Distributed Computing, 34(4):301–318, August 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-021-00389-4.

Chalopin:2012:EPA

[CGM12] Jérémie Chalopin, Emmanuel Godard, and Yves Métivier. Election in partially anonymous networks with arbitrary knowledge in message passing systems. Distributed Computing, 25(4):297–311, August 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-012-0163-y.

Castaneda:2022:UC

[CGM22] Armando Castañeda, Yannai A. Gonczarowski, and Yoram Moses. Unbeatable consensus. Distributed Computing, 35(2):123–143, April 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-021-00417-3.

Chlebus:2002:DBA

[CGR02] Bogdan S. Chlebus, Leszek Gasieniec, and Wojciech Rytter. Deterministic broadcasting in ad hoc radio networks. Distributed Computing, 15(1):27–38, January 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s446-002-8028-1.

Chu:2001:DTA

[CH01] Francis C. Chu and Joseph Y. Halpern. A decision-theoretic approach to reliable message delivery. Distributed Computing, 14 (1):1–16, January 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/PL00008922.

Censor-Hillel:2023:CDC

[CHCGS23a] Keren Censor-Hillel, Shir Cohen, Ran Gelles, and Gal Sela. Correction to: Distributed computations in fully-defective networks. Distributed Computing, 36(4):529, December 2023. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-

023-00457-x. See [CHCGS23b].

Censor-Hillel:2023:DCF

[CHCGS23b] Keren Censor-Hillel, Shir Cohen, Ran Gelles, and Gal Sela. Distributed computations in fully-defective networks. Distributed Computing, 36(4):501–528, December 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-

023-00452-2. See correction [CHCGS23a].

Censor-Hillel:2020:FDA

[CHD20] Keren Censor-Hillel and Michal Dory. Fast distributed approximation for TAP and 2-edge-connectivity. Distributed Computing, 33(2):145–168, April 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-019-00353-3.

Censor-Hillel:2021:FAS

[CHDL21] Keren Censor-Hillel, Michal Dory, and Dean Leitersdorf. Fast approximate shortest paths in the congested clique. Distributed Computing, 34(6):463–487, December 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-020-00380-5.

Censor-Hillel:2019:FDA

[CHFV19] Keren Censor-Hillel, Eldar Fischer, and Yadu Vasudev. Fast distributed algorithms for testing graph properties. Distributed Computing, 32(1):41–57, February 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-018-0324-8.

Censor-Hillel:2019:MAD

[CHGH19] Keren Censor-Hillel, Ran Gelles, and Bernhard Haeupler. Making asynchronous distributed computations robust to noise. Distributed Computing, 32(5):405–421, October 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
018-0343-5.

Censor-Hillel:2014:SUR

[CHGN14] Keren Censor-Hillel, Seth Gilbert, and Calvin Newport. Structuring unreliable radio networks. Distributed Computing, 27(1):1–19, February 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-013-0198-8.

Censor-Hillel:2015:BCC

[CHHM15] Keren Censor-Hillel, Bernhard Haeupler, and Muriel Médard. Bounded-contention coding for the additive network model. Distributed Computing, 28(5):297–308, October 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
015-0244-9.

Cheng:2014:PC

[CHK14] Steven Cheng, Lisa Higham, and Jalal Kawash. Partition consistency. Distributed Computing, 27(5):363–389, October 2014. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0205-0.

Censor-Hillel:2019:AMC

[CHKS19] Keren Censor-Hillel, Petteri Kaski, and Jukka Suomela. Algebraic methods in the congested clique. Distributed Computing, 32(6): 461–478, December 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-016-0270-2.

Censor-Hillel:2018:DCP

[CHKY18] Keren Censor-Hillel, Telikepalli Kavitha, and Amir Yehudayoff. Distributed construction of purely additive spanners. Distributed Computing, 31(3):223–240, June 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-017-0306-2.

Chen:2019:SML

[CHM19] Ning Chen, Martin Hoefer, and Peihan Miao. Secretary markets with local information. Distributed Computing, 32(5):361–378, October 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-018-0327-5.

Censor-Hillel:2020:DLD

[CHPS20] Keren Censor-Hillel, Merav Parter, and Gregory Schwartzman. Derandomizing local distributed algorithms under bandwidth restrictions. Distributed Computing, 33(3–4):349–366, June 2020.

CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-020-00376-1.

Chan:2020:LBS

[CHT20] David Yu Cheng Chan, Vassos Hadzilacos, and Sam Toueg. Life beyond set agreement. Distributed Computing, 33(3–4):255–277, June 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-020-00372-5.

Cooper:2011:DRW

[CIK11] Colin Cooper, David Ilcinkas, and Adrian Kosowski. Derandomizing random walks in undirected graphs using locally fair exploration strategies. Distributed Computing, 24(2):??, October 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-011-0138-4.

Ciuffoletti:1994:RVC

[Ciu94] Augusto Ciuffoletti. Reliability versus cost: Design of a probabilistic broadcast algorithm. Distributed Computing, 7(3): 115–127, March 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02277856.

Capdevielle:2017:UCA

[CJM17] Claire Capdevielle, Colette Johnen, and Alessia Milani. On the uncontended complexity of anonymous agreement. Distributed Computing, 30(6):459–468, December 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-017-0297-z.

Czumaj:2020:DCC

[CK20] Artur Czumaj and Christian Konrad. Detecting cliques in CON-GEST networks. Distributed Computing, 33(6):533–543, December 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452

(electronic). URL https://link.springer.com/article/10.
1007/s00446-019-00368-w.

Chlebus:2006:PWB

[CKL06] Bogdan S. Chlebus, Dariusz R. Kowalski, and Andrzej Lingas. Performing work in broadcast networks. Distributed Computing, 18(6):435–451, June 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0153-4.

Czyzowicz:2019:SLF

[CKO19] Jurek Czyzowicz, Evangelos Kranakis, and Jaroslav Opatrny. Search on a line with faulty robots. Distributed Computing, 32 (6):493–504, December 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0296-0.

Czyzowicz:2012:HMW

[CKP12] Jurek Czyzowicz, Adrian Kosowski, and Andrzej Pelc. How to meet when you forget: log-space rendezvous in arbitrary graphs. Distributed Computing, 25(2):165–178, May 2012. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
011-0141-9.

Czyzowicz:2014:TVS

[CKP14] Jurek Czyzowicz, Adrian Kosowski, and Andrzej Pelc. Time versus space trade-offs for rendezvous in trees. Distributed Computing, 27(2):95–109, April 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-013-0201-4.

Czyzowicz:2015:LSC

[CKP15] Jurek Czyzowicz, Evangelos Kranakis, and Eduardo Pacheco. Localization for a system of colliding robots. Distributed Computing, 28(4):245–252, August 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0234-3.

Chlebus:2009:MTM

[CKR09] Bogdan S. Chlebus, Dariusz R. Kowalski, and Mariusz A. Rokicki. Maximum throughput of multiple access channels in adversarial environments. Distributed Computing, 22(2):93–116, October 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-009-0086-4.

Chlebus:2022:DBB

[CKV22] Bogdan S. Chlebus, Dariusz R. Kowalski, and Shailesh Vaya. Distributed bare-bones communication in wireless networks. Distributed Computing, 35(1):59–80, February 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
021-00413-7.

Clarke:1986:DCI

[Cla86] Edmund M. Clarke. Distributed computing issues in hardware design. Distributed Computing, 1(4):185–186, December 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01660030.

Cadambe:2017:CSA

[CLM17] Viveck R. Cadambe, Nancy Lynch, and Peter Musial. A coded shared atomic memory algorithm for message passing architectures. Distributed Computing, 30(1):49–73, February 2017. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
016-0275-x.

Chandy:1986:HPL

[CM86a] K. M. Chandy and Jayadev Misra. How processes learn. Distributed Computing, 1(1):40–52, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01843569.

Chandy:1986:SAP

[CM86b] K. Mani Chandy and J. Misra. Systolic algorithms as programs. Distributed Computing, 1(3):177–183, September 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01661171.

Chockler:2005:ADP

[CM05] Gregory Chockler and Dahlia Malkhi. Active Disk Paxos with infinitely many processes. Distributed Computing, 18(1):73–84, July 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-005-0123-x.

Chalopin:2010:PSB

[CM10] Jérémie Chalopin and Yves Métivier. On the power of synchronization between two adjacent processes. Distributed Computing, 23(3):177–196, November 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0115-3.

Chechik:2022:SSS

[CM22] Shiri Chechik and Doron Mukhtar. Single-source shortest paths in the CONGEST model with improved bounds. Distributed Computing, 35(4):357–374, August 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00412-8.

Chiao:2002:SSQ

[CMN02] Yiwei Chiao, Masaaki Mizuno, and Mitchell L. Neilsen. A self-stabilizing quorum-based protocol for maxima computing. Distributed Computing, 15(1):49–55, January 2002. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s446-002-8030-8.

Cruciani:2023:PTE

[CMQR23] Emilio Cruciani, Hlafo Alfie Mimun, Matteo Quattropani, and Sara Rizzo. Phase transition of the k-majority dynamics in biased communication models. Distributed Computing, 36(2):107–135, June 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-023-00444-2.

Clementi:2013:FFM

[CMS13] Andrea Clementi, Angelo Monti, and Riccardo Silvestri. Fast flooding over Manhattan. Distributed Computing, 26(1):25–38, February 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-012-0182-8.

Carbone:2018:CL

[CMS18] Marco Carbone, Fabrizio Montesi, and Carsten Schürmann. Choreographies, logically. Distributed Computing, 31(1):51–67, February 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-017-0295-1.

Cruciani:2021:PTC

[CNS21] Emilio Cruciani, Emanuele Natale, and Giacomo Scornavacca. Phase transition of the 2-choices dynamics on core–periphery networks. Distributed Computing, 34(3):207–225, June 2021. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
021-00396-5.

Correia:2005:LCB

[CNV05] Miguel Correia, Nuno Ferreira Neves, and Paulo Veríssimo. Low complexity Byzantine-resilient consensus. Distributed Computing, 17(3):237–249, March 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-004-0110-7.

Coan:1993:EAU

[Coa93] Brian A. Coan. Efficient agreement using fault diagnosis. Distributed Computing, 7(2):87–98, December 1993. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02280838.

Cavalli:1988:EAS

[CP88] Ana R. Cavalli and Etienne Paul. Exhaustive analysis and simulation for distributed systems, both sides of the same coin. Distributed Computing, 2(4):213–225, December 1988. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01872847.

Comer:1989:UND

[CP89] Douglas E. Comer and Larry L. Peterson. Understanding naming in distributed systems. Distributed Computing, 3(2):51–60, June 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01558663.

Cook:2002:OLR

[CPP02] Stephen A. Cook, Jan Pachl, and Irwin S. Pressman. The optimal location of replicas in a network using a READ-ONE-WRITE-ALL policy. Distributed Computing, 15(1):57–66, January 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s446-002-8031-5.

Chatterjee:2020:CLE

[CPR20] Soumyottam Chatterjee, Gopal Pandurangan, and Peter Robinson. The complexity of leader election in diameter-two networks. Distributed Computing, 33(2):189–205, April 2020. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
019-00354-2.

Chung:2017:DAL

[CPS17] Kai-Min Chung, Seth Pettie, and Hsin-Hao Su. Distributed algorithms for the Lovász local lemma and graph coloring. (French) []. Distributed Computing, 30(4):261–280, August 2017. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
016-0287-6.

Castaneda:2010:NCT

[CR10] Armando Castañeda and Sergio Rajsbaum. New combinatorial topology bounds for renaming: the lower bound. Distributed Computing, 22(5–6):287–301, August 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0108-2.

Cristian:1989:PCS

[Cri89] Flaviu Cristian. Probabilistic clock synchronization. Distributed Computing, 3(3):146–158, September 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01784024.

Cristian:1991:RAP

[Cri91] Flaviu Cristian. Reaching agreement on processor-group membrship in synchronous distributed systems. Distributed Computing, 4(4):175–187, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01784719.

Castaneda:2023:SLI

[CRR23] Armando Castañeda, Sergio Rajsbaum, and Michel Raynal. Setlinearizable implementations from read/write operations: Sets, fetch & increment, stacks and queues with multiplicity. Distributed Computing, 36(2):89–106, June 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-022-00440-y.

Choy:1994:ASM

[CS94] Manhoi Choy and Ambuj K. Singh. Adaptive solutions to the mutual exclusion problem. Distributed Computing, 8(1):1–17, August 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02283567.

Chen:2011:EOA

[CSC11] Lung-Pin Chen, Der-Johng Sun, and William Chu. Efficient online algorithm for identifying useless states in distributed systems. Distributed Computing, 23(5–6):359–372, April 2011. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
010-0124-2.

Clementi:2015:ISD

[CST15] Andrea Clementi, Riccardo Silvestri, and Luca Trevisan. Information spreading in dynamic graphs. Distributed Computing, 28 (1):55–73, February 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0219-2.

Chernoy:2010:PDT

[CSZ10] Viacheslav Chernoy, Mordechai Shalom, and Shmuel Zaks. On the performance of Dijkstra’s third self-stabilizing algorithm for mutual exclusion and related algorithms. Distributed Computing, 23(1):43–60, September 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-010-0104-6.

Critchlow:1996:ISA

[CT96] Carol Critchlow and Kim Taylor. The inhibition spectrum and the achievement of causal consistency. Distributed Computing, 10(1):11–27, July 1996. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050021.

Coan:1990:TCR

[CW90] Brian A. Coan and Jennifer Lundelius Welch. Transaction commit in a realistic timing model. Distributed Computing, 4(2):87–103, June 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01786634.

Drummond:1993:LCC

[DB93] Rogério Drummond and Özalp Babaoĝlu. Low-cost clock synchronization. Distributed Computing, 6(4):193–203, July 1993.

CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02242707.

Doerr:2016:SOR

[DDM16] Benjamin Doerr, Carola Doerr, and Shlomo Moran. Simple and optimal randomized fault-tolerant rumor spreading. Distributed Computing, 29(2):89–104, April 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0238-z.

DAngelo:2014:GRU

[DDN14] Gianlorenzo D’Angelo, Gabriele Di Stefano, and Alfredo Navarra. Gathering on rings under the look–compute–move model. Distributed Computing, 27(4):255–285, August 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0212-9.

DiLuna:2020:DED

[DDS20] G. Di Luna, S. Dobrev, and N. Santoro. Distributed exploration of dynamic rings. Distributed Computing, 33(1):41–67, February 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-018-0339-1.

Dinitz:2018:SAD

[DFN18] Michael Dinitz, Jeremy T. Fineman, and Calvin Newport. Smoothed analysis of dynamic networks. Distributed Computing, 31(4):273–287, August 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0300-8.

Duflot:2004:RDP

[DFP04] Marie Duflot, Laurent Fribourg, and Claudine Picaronny. Randomized dining philosophers without fairness assumption. Distributed Computing, 17(1):65–76, February 2004. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
003-0102-z.

Dobrev:2006:SBH

[DFS06] Stefan Dobrev, Paola Flocchini, and Nicola Santoro. Searching for a black hole in arbitrary networks: optimal mobile agents protocols. Distributed Computing, 19(1):1–99999, September 2006. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
006-0154-y.

DiLuna:2020:FTS

[DFV20] Giuseppe A. Di Luna, Paola Flocchini, and Giovanni Viglietta. Fault-tolerant simulation of population protocols. Distributed Computing, 33(6):561–578, December 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-020-00377-0.

DiLuna:2022:TTM

[DFV22] Giuseppe A. Di Luna, Paola Flocchini, and Giovanni Viglietta. TuringMobile: a Turing machine of oblivious mobile robots with limited visibility and its applications. Distributed Computing, 35 (2):105–122, April 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-021-00406-6.

Dodis:2005:SPK

[DFY05] Yevgeniy Dodis, Nelly Fazio, and Moti Yung. Scalable publickey tracing and revoking. Distributed Computing, 17(4):323–347, May 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-004-0120-5.

Das:2015:FSG

[DFY15] Shantanu Das, Paola Flocchini, and Masafumi Yamashita. Forming sequences of geometric patterns with oblivious mobile robots. Distributed Computing, 28(2):131–145, April 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0220-9.

DiLuna:2020:SFP

[DFY20a] Giuseppe A. Di Luna, Paola Flocchini, and Yukiko Yamauchi. Shape formation by programmable particles. Distributed Computing, 33(1):69–101, February 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-019-00350-6.

DiLuna:2020:MPA

[DFY20b] Giuseppe Antonio Di Luna, Paola Flocchini, and Masafumi Yamashita. Meeting in a polygon by anonymous oblivious robots. Distributed Computing, 33(5):445–469, October 2020. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
019-00362-2.

Dutta:2005:IPI

[DG05] Partha Dutta and Rachid Guerraoui. The inherent price of indulgence. Distributed Computing, 18(1):85–98, July 2005. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0124-9.

Danek:2010:CCG

[DG10] Robert Danek and Wojciech Golab. Closing the complexity gap between FCFS mutual exclusion and mutual exclusion. Distributed Computing, 23(2):87–111, October 2010. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
010-0096-2.

Delporte-Gallet:2015:WFA

[DGFK15] Carole Delporte-Gallet, Hugues Fauconnier, and Petr Kuznetsov. Wait-freedom with advice. Distributed Computing, 28(1):3–19, February 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-014-0231-6.

Delporte-Gallet:2011:DPA

[DGFT11a] Carole Delporte-Gallet, Hugues Fauconnier, and Andreas Tielmann. The disagreement power of an adversary. Distributed Computing, 24(3–4):137–147, November 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-010-0122-4.

Delporte-Gallet:2011:MIA

[DGFT11b] Carole Delporte-Gallet, Hugues Fauconnier, and Sam Toueg. The minimum information about failures for solving non-local tasks in message-passing systems. Distributed Computing, 24(5):255–269, December 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-011-0146-4.

Delporte-Gallet:2013:BAH

[DGFTT13] Carole Delporte-Gallet, Hugues Fauconnier, and Hung Tran-The. Byzantine agreement with homonyms. Distributed Computing, 26 (5–6):321–340, October 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-013-0190-3.

Dani:2023:WJM

[DGHP23] Varsha Dani, Aayush Gupta, Thomas P. Hayes, and Seth Pettie. Wake up and join me! An energy-efficient algorithm for maximal matching in radio networks. Distributed Computing, 36(3): 373–384, September 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-022-00426-w.

Dutta:2007:OCF

[DGK07] Partha Dutta, Rachid Guerraoui, and Idit Keidar. The overhead of consensus failure recovery. Distributed Computing, 19 (5–6):373–386, April 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0017-6.

Dickerson:2020:ACS

[DGK20] Thomas Dickerson, Paul Gazzillo, and Eric Koskinen. Adding concurrency to smart contracts. Distributed Computing, 33(3– 4):209–225, June 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-019-00357-z.

Dubois:2019:WFD

[DGS19] Swan Dubois, Rachid Guerraoui, and Pierre Sens. The weakest failure detector for eventual consistency. Distributed Computing, 32(6):479–492, December 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-016-0292-9.

Dolev:2005:GIA

[DGW05] Shlomi Dolev, Seth Gilbert, and Jennifer L. Welch. GeoQuorums: implementing atomic memory in mobile ad hoc networks.

Distributed Computing, 18(2):125–155, November 2005. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0140-9.

Dolev:2007:PCT

[DH07] Shlomi Dolev and Ted Herman. Parallel composition for time-to-fault adaptive stabilization. Distributed Computing, 20(1):29–38, July 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-007-0028-y.

Dijkstra:1986:BPS

[Dij86] Edsger W. Dijkstra. A belated proof of self-stabilization. Distributed Computing, 1(1):5–6, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01843566.

Dolev:1993:SSD

[DIM93] Shlomi Dolev, Amos Israeli, and Shlomo Moran. Self-stabilization of dynamic systems assuming only read/write atomicity. Distributed Computing, 7(1):3–16, November 1993. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02278851.

Damian:2009:DCL

[DJ09] Mirela Damian and Nagesh Javali. Distributed construction of low-interference spanners. Distributed Computing, 22(1):15–28, April 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-009-0083-7.

Datta:2000:SSD

[DJV00] Ajoy K. Datta, Colette Johnen, and Vincent Villain. Selfstabilizing depth-first token circulation in arbitrary rooted networks. Distributed Computing, 13(4):207–218, November 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008919.

Dolev:2008:HSS

[DK08] Shlomi Dolev and Ronen I. Kat. HyperTree for self-stabilizing peer-to-peer systems. Distributed Computing, 20(5):375–388, February 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-007-0038-9.

Dani:2017:SMP

[DKZ17] Varsha Dani, Valerie King, and Mahdi Zamani. Secure multiparty computation in large networks. Distributed Computing, 30(3):193–229, June 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-016-0284-9.

Dima:2012:SSS

[DL12] Cătălin Dima and Ruggero Lanotte. A study on shuffle, stopwatches and independently evolving clocks. Distributed Computing, 25(1):5–33, March 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-011-0148-2.

Detlefs:2002:LFR

[DMS02] David L. Detlefs, Paul A. Martin, and Guy L. Steele Jr. Lockfree reference counting. Distributed Computing, 15(4):255–271, December 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-002-0079-z.

Dani:2015:SMR

[DMS15] Varsha Dani, Mahnush Movahedi, and Jared Saia. Scalable mechanisms for rational secret sharing. Distributed Computing, 28 (3):171–187, June 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0233-4.

DiStefano:2017:OGO

[DN17] Gabriele Di Stefano and Alfredo Navarra. Optimal gathering of oblivious robots in anonymous graphs and its application on trees and rings. Distributed Computing, 30(2):75–86, April 2017. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
016-0278-7.

DAngelo:2017:UAG

[DNN17] Gianlorenzo D’Angelo, Alfredo Navarra, and Nicolas Nisse. A unified approach for gathering and exclusive searching on rings under weak assumptions. Distributed Computing, 30(1):17–48, February 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-016-0274-y.

Dereniowski:2014:LEA

[DP14] Dariusz Dereniowski and Andrzej Pelc. Leader election for anonymous asynchronous agents in arbitrary networks. Distributed Computing, 27(1):21–38, February 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-013-0196-x.

Dieudonne:2015:DPA

[DP15] Yoann Dieudonné and Andrzej Pelc. Deterministic polynomial approach in the plane. Distributed Computing, 28(2):111–129, April 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-014-0216-5.

Defago:2020:SSG

[DPBRP20] Xavier Défago, Maria Potop-Butucaru, and Philippe Raipin-Parvédy. Self-stabilizing gathering of mobile robots under crash or Byzantine faults. Distributed Computing, 33(5):393–421, October 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-019-00359-x.

Daymude:2023:CAM

[DRS23] Joshua J. Daymude, Andréa W. Richa, and Christian Scheideler. The canonical amoebot model: algorithms and concurrency control. Distributed Computing, 36(2):159–192, June 2023. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
023-00443-3.

Dally:1986:TRC

[DS86] William J. Dally and Charles L. Seitz. The torus routing chip. Distributed Computing, 1(4):187–196, December 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01660031.

Dolev:2003:SAL

[DS03] Shlomi Dolev and Frank Stomp. Safety assurance via on-line monitoring. Distributed Computing, 16(4):269–277, December 2003.

CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-003-0089-5.

Dolev:2008:ISI

[DS08] Shlomi Dolev and Alex Shvartsman. Introduction to special issue dedicated to the DISC 20th anniversary. Distributed Computing, 20(6):389–390, April 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0054-4.

Doty:2018:SLE

[DS18] David Doty and David Soloveichik. Stable leader election in population protocols requires linear time. Distributed Computing, 31 (4):257–271, August 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-016-0281-z.

Ducourthial:2001:SSO

[DT01] Bertrand Ducourthial and Sébastien Tixeuil. Self-stabilization with r-operators. Distributed Computing, 14(3):147–162, July 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008934.

Ebergen:1991:FAD

[Ebe91] Jo C. Ebergen. A formal approach to designing delay-insensitive circuits. Distributed Computing, 5(3):107–119, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02252954.

Eden:2022:STD

[EFO22] Talya Eden, Nimrod Fiat, and Rotem Oshman. Sublinear-time distributed algorithms for detecting small cliques and even cycles. Distributed Computing, 35(3):207–234, June 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
021-00409-3.

Ellen:2008:SCU

[EFR08] Faith Ellen, Panagiota Fatourou, and Eric Ruppert. The space complexity of unbounded timestamps. Distributed Computing, 21(2):103–115, July 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0060-6.

Enders:1993:GBS

[EFT93] Reinhard Enders, Thomas Filkorn, and Dirk Taubner. Generating BDDs for symbolic model checking in CCS. Distributed Computing, 6(3):155–164, April 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02242704.

Ellen:2016:UCE

[EFT16] Faith Ellen, Panagiota Fatourou, and Corentin Travers. Universal constructions that ensure disjoint-access parallelism and waitfreedom. Distributed Computing, 29(4):251–277, August 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-015-0261-8.

Elkin:2023:IWA

[EGN23] Michael Elkin, Yuval Gitlitz, and Ofer Neiman. Improved weighted additive spanners. Distributed Computing, 36(3):385– 394, September 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-022-00433-x.

Emek:2009:BUR

[EGS09] Yuval Emek, Leszek Gaşieniec, and Chang Su. Broadcasting in UDG radio networks with unknown topology. Distributed Computing, 21(5):331–351, February 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-008-0075-z.

Ellen:2020:CBC

[EGZ20] Faith Ellen, Rati Gelashvili, and Leqi Zhu. A complexity-based classification for multiprocessor synchronization. Distributed Computing, 33(2):125–144, April 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-019-00361-3.

Eisler:2007:WFD

[EHT07] Jonathan Eisler, Vassos Hadzilacos, and Sam Toueg. The weakest failure detector to solve nonuniform consensus. Distributed Computing, 19(4):335–359, March 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-006-0019-4.

Esparza:2021:CVP

[EJWK21] Javier Esparza, Stefan Jaax, and Chana Weil-Kennedy. The complexity of verifying population protocols. Distributed Computing, 34(2):133–177, April 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-021-00390-x.

Emek:2011:NBC

[EK11] Yuval Emek and Amos Korman. New bounds for the controller problem. Distributed Computing, 24(3–4):177–186, November 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-010-0119-z.

Emek:2016:EDS

[EKP16] Yuval Emek, Erez Kantor, and David Peleg. On the effect of the deployment setting on broadcasting in Euclidean radio networks. Distributed Computing, 29(6):409–434, November 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-016-0273-z.

Eyal:2011:DDC

[EKR11] Ittay Eyal, Idit Keidar, and Raphael Rom. Distributed data clustering in sensor networks. Distributed Computing, 24(5):207–222, December 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-011-0143-7.

Eyal:2014:LLM

[EKR14] Ittay Eyal, Idit Keidar, and Raphael Rom. LiMoSense: live monitoring in dynamic sensor networks. Distributed Computing, 27 (5):313–328, October 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0213-8.

Eftekhari:2016:DAB

[EKS16] Mohsen Eftekhari, Evangelos Kranakis, and Sunil Shende. Distributed algorithms for barrier coverage using relocatable sensors. Distributed Computing, 29(5):361–376, October 2016. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
016-0267-x.

Engelhardt:2009:CCC

[EM09] Kai Engelhardt and Yoram Moses. Causing communication closure: safe program composition with reliable non-FIFO channels. Distributed Computing, 22(2):73–91, October 2009. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
009-0081-9.

El-Mhamdi:2022:GDB

[EMGR22] El-Mahdi El-Mhamdi, Rachid Guerraoui, and Sébastien Rouault. Genuinely distributed Byzantine machine learning. Distributed Computing, 35(4):305–331, August 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-022-00427-9.

Elkin:2018:EDC

[EN18] Michael Elkin and Ofer Neiman. On efficient distributed construction of near optimal routing schemes. Distributed Computing, 31(2):119–137, April 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0304-4.

Elkin:2022:LSH

[EN22] Michael Elkin and Ofer Neiman. Linear-size hopsets with small hopbound, and constant-hopbound hopsets in RNC. Distributed Computing, 35(5):419–437, October 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-022-00431-z.

Egecioglu:1994:NSP

[ES94] Ömer Eğecioğlu and Ambuj K. Singh. Naming symmetric processes using shared variables. Distributed Computing, 8(1):19–38, August 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02283568.

Elkin:2006:EAC

[EZ06] Michael Elkin and Jian Zhang. Efficient algorithms for constructing (1+ ∈, β)-spanners in the distributed and streaming models. Distributed Computing, 18(5):375–385, April 2006. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0147-2.

Flatebo:1995:SSM

[FDS95] Mitchell Flatebo, Ajoy Kumar Datta, and Anneke A. Schoone. Self-stabilizing multi-token rings. Distributed Computing, 8(3): 133–142, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02242715.

Feijs:1999:GFI

[Fei99] L. M. G. Feijs. Generating FSMs from interworkings. Distributed Computing, 12(1):31–40, March 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050053.

Fekete:1990:AOA

[Fek90] A. D. Fekete. Asymptotically optimal algorithms for approximate agreement. Distributed Computing, 4(1):9–29, March 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01783662.

Feuilloley:2021:RDP

[FFP21] Laurent Feuilloley, Pierre Fraigniaud, and Mor Perry. Redundancy in distributed proofs. Distributed Computing, 34(2):113– 132, April 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-020-00386-z.

Fraigniaud:1997:URS

[FG97] Pierre Fraigniaud and Cyril Gavoille. Universal routing schemes. Distributed Computing, 10(2):65–78, February 1997. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050025.

Frolund:2001:XAT

[FG01] Svend Frølund and Rachid Guerraoui. X-Ability: a theory of replication. Distributed Computing, 14(4):231–249, December 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460100065.

Fraigniaud:2014:GRS

[FG14] Pierre Fraigniaud and George Giakkoupis. Greedy routing in small-world networks with power-law degrees. Distributed Computing, 27(4):231–253, August 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0210-y.

Fraigniaud:2001:IRS

[FGM01] Pierre Fraigniaud, Cyril Gavoille, and Bernard Mans. Interval routing schemes allow broadcasting with linear messagecomplexity. Distributed Computing, 14(4):217–229, December 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460100058.

Fineman:2019:CRF

[FGN19] Jeremy T. Fineman, Seth Gilbert, and Calvin Newport. Contention resolution on a fading channel. Distributed Computing, 32 (6):517–533, December 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-018-0323-9.

Fraigniaud:2006:ESE

[FGP06] Pierre Fraigniaud, Cyril Gavoille, and Christophe Paul. Eclecticism shrinks even small worlds. Distributed Computing, 18 (4):279–291, March 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0137-4.

Fraigniaud:2009:DCA

[FGP09] Pierre Fraigniaud, Cyril Gavoille, and Andrzej Pelc. Distributed computing with advice: information sensitivity of graph coloring. Distributed Computing, 21(6):395–403, March 2009. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
008-0076-y.

Fraigniaud:2014:RDD

[FGP14] Pierre Fraigniaud, Mika Göös, and David Peleg. Randomized distributed decision. Distributed Computing, 27(6):419–434, December 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-014-0211-x.

Fatourou:2004:RMW

[FH04] Panagiota Fatourou and Maurice Herlihy. Read-modify-write networks. Distributed Computing, 17(1):33–46, February 2004. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
003-0097-5.

Feinerman:2017:BBS

[FHK17] Ofer Feinerman, Bernhard Haeupler, and Amos Korman. Breathe before speaking: efficient information dissemination despite noisy, limited and anonymous communication. Distributed Computing, 30(5):339–355, October 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-015-0249-4.

Fich:2006:IWC

[FHS06] Faith Ellen Fich, Danny Hendler, and Nir Shavit. On the inherent weakness of conditional primitives. Distributed Computing, 18 (4):267–277, March 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0136-5.

Florian:2022:SPA

[FHS22] Martin Florian, Sebastian Henningsen, and Björn Scheuermann. The sum of its parts: Analysis of federated Byzantine agreement systems. Distributed Computing, 35(5):399–417, October 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-022-00430-0.

Fagin:1997:KBP

[FHV97] Ronald Fagin, Joseph Y. Halpern, and Moshe Y. Vardi. Knowledge-based programs. Distributed Computing, 10(4):199– 225, July 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050038.

Fich:2005:ISI

[Fic05] Faith Ellen Fich. Introduction to the special issue DISC 2003. Distributed Computing, 18(2):111, November 2005. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0132-9.

Finkel:1994:DTP

[Fin94] Alain Finkel. Decidability of the termination problem for completely specified protocols. Distributed Computing, 7(3):129–135, March 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02277857.

Fischer:2020:IDD

[Fis20] Manuela Fischer. Improved deterministic distributed matching via rounding. Distributed Computing, 33(3–4):279–291, June 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-018-0344-4.

Feige:2000:NWH

[FK00] Uriel Feige and Robert Krauthgamer. Networks on which hotpotato routing does not livelock. Distributed Computing, 13 (1):53–58, January 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050005.

Fatourou:2017:LUB

[FK17a] Panagiota Fatourou and Nikolaos D. Kallimanis. Lower and upper bounds for single-scanner snapshot implementations. Distributed Computing, 30(4):231–260, August 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-016-0286-7.

Feinerman:2017:AP

[FK17b] Ofer Feinerman and Amos Korman. The ANTS problem. Distributed Computing, 30(3):149–168, June 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
016-0285-8.

Fatourou:2020:RFU

[FK20] Panagiota Fatourou and Nikolaos D. Kallimanis. The RedBlue family of universal constructions. Distributed Computing, 33(6): 485–513, December 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-020-00370-7.

Fan:2006:GCS

[FL06] Rui Fan and Nancy Lynch. Gradient clock synchronization. Distributed Computing, 18(4):255–266, March 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0135-6.

Fischer:1986:EIP

[FLM86] Michael J. Fischer, Nancy A. Lynch, and Michael Merritt. Easy impossibility proofs for distributed consensus problems. Distributed Computing, 1(1):26–39, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01843568.

Fischer:2003:ATD

[FM03] Michael J. Fischer and Michael Merritt. Appraising two decades of distributed computing theory research. Distributed Computing, 16(2–3):239–247, September 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-003-0096-6.

Faloutsos:2004:LTO

[FM04] Michalis Faloutsos and Mart Molle. A linear-time optimalmessage distributed algorithm for minimum spanning trees. Distributed Computing, 17(2):151–170, August 2004. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
004-0107-2.

Fanelli:2011:BRD

[FM11] Angelo Fanelli and Luca Moscardelli. On best response dynamics in weighted congestion games with polynomial delays. Distributed Computing, 24(5):245–254, December 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-011-0145-5.

Fribourg:2006:CSS

[FMlP06] Laurent Fribourg, Stéphane Messika, and laudine Picaronny. Coupling and self-stabilization. Distributed Computing, 18(3): 221–232, February 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0142-7.

Fraigniaud:2019:NRS

[FN19] Pierre Fraigniaud and Emanuele Natale. Noisy rumor spreading and plurality consensus. Distributed Computing, 32(4):257–276, August 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-018-0335-5.

Flammini:2001:ALB

[FNP01] Michele Flammini, Enrico Nardelli, and Guido Proietti. ATM layouts with bounded hop count and congestion. Distributed Computing, 14(2):65–73, April 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/PL00008929.

Fischer:2023:DAD

[FO23] Orr Fischer and Rotem Oshman. A distributed algorithm for directed minimum-weight spanning tree. Distributed Computing, 36(1):57–87, March 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-021-00398-3.

Fischer:2008:AWE

[FOV08] Simon Fischer, Lars Olbrich, and Berthold Vöcking. Approximating Wardrop equilibria with finitely many agents. Distributed Computing, 21(2):129–139, July 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-008-0057-1.

Fusco:2010:BUR

[FP10] Emanuele G. Fusco and Andrzej Pelc. Broadcasting in UDG radio networks with missing and inaccurate information. Distributed Computing, 22(3):167–183, March 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0093-5.

Fusco:2011:HMM

[FP11] Emanuele G. Fusco and Andrzej Pelc. How much memory is needed for leader election. Distributed Computing, 24(2):??, October 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-011-0131-y.

Fusco:2015:KLS

[FP15] Emanuele G. Fusco and Andrzej Pelc. Knowledge, level of symmetry, and time of leader election. Distributed Computing, 28 (4):221–232, August 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0237-0.

Fraigniaud:2001:ALU

[FPP01] Pierre Fraigniaud, Andrzej Pelc, and Stéphane Pérennes. Assigning labels in an unknown anonymous network with a leader. Distributed Computing, 14(3):163–183, July 2001. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/PL00008935.

Feigenbaum:2005:BBM

[FPS05] Joan Feigenbaum, Christos Papadimitriou, and Scott Shenker. A BGP-based mechanism for lowest-cost routing. Distributed Computing, 18(1):61–72, July 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-005-0122-y.

Fraigniaud:2019:RPL

[FPSP19] Pierre Fraigniaud, Boaz Patt-Shamir, and Mor Perry. Randomized proof-labeling schemes. Distributed Computing, 32(3):217– 234, June 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-018-0340-8.

Flocchini:2017:DCM

[FPV17] Paola Flocchini, Giuseppe Prencipe, and Giovanni Viglietta. Distributed computing by mobile robots: uniform circle formation. Distributed Computing, 30(6):413–457, December 2017. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
016-0291-x.

Fich:2003:HIR

[FR03] Faith Fich and Eric Ruppert. Hundreds of impossibility results for distributed computing. Distributed Computing, 16(2–3): 121–163, September 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-003-0091-y.

Fraigniaud:2007:SID

[Fra07] Pierre Fraigniaud. Special issue DISC 2005. Distributed Computing, 20(3):163, October 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-007-0039-8.

Fraigniaud:2013:SIS

[Fra13] Pierre Fraigniaud. Special issue with selected papers from PODC 2011. Distributed Computing, 26(5–6):271, October 2013. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0197-9.

Friedman:1995:IHC

[Fri95] Roy Friedman. Implementing hybrid consistency with high-level synchronization operations. Distributed Computing, 9(3):119–129, December 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050014.

Feigenbaum:2011:ICI

[FRS11] Joan Feigenbaum, Vijay Ramachandran, and Michael Schapira. Incentive-compatible interdomain routing. Distributed Computing, 23(5–6):301–319, April 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0126-8.

Fraigniaud:2013:LCW

[FRT13] Pierre Fraigniaud, Sergio Rajsbaum, and Corentin Travers. Locality and checkability in wait-free computing. Distributed Computing, 26(4):223–242, August 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-013-0188-x.

Fugger:2012:RFT

[FS12] Matthias Függer and Ulrich Schmid. Reconciling fault-tolerant distributed computing and systems-on-chip. Distributed Computing, 24(6):323–355, January 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0151-7.

Feigenbaum:2006:MDP

[FSS06] Joan Feigenbaum, Rahul Sami, and Scott Shenker. Mechanism design for policy routing. Distributed Computing, 18(4):293–305, March 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-005-0134-7.

Feng:2020:WCS

[FSY20] Weiming Feng, Yuxin Sun, and Yitong Yin. What can be sampled locally? Distributed Computing, 33(3–4):227–253, June 2020. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
018-0332-8.

Feng:2015:NOU

[FZ15] Yuan Feng and Lijun Zhang. A nearly optimal upper bound for the self-stabilization time in Herman’s algorithm. Distributed Computing, 28(4):233–244, August 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-015-0241-z.

Gelashvili:2018:OSC

[Gel18] Rati Gelashvili. On the optimal space complexity of consensus for anonymous processes. Distributed Computing, 31(4):317–326, August 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-018-0331-9.

Gerth:1999:SCL

[Ger99] Rob Gerth. Sequential consistency and the lazy caching algorithm. Distributed Computing, 12(2–3):57–59, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050057.

Gao:2005:LFD

[GGH05] H. Gao, J. F. Groote, and W. H. Hesselink. Lock-free dynamic hash tables with open addressing. Distributed Computing, 18(1):21–42, July 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-004-0115-2.

Georgiou:2011:MDC

[GGK11] Chryssis Georgiou, Seth Gilbert, and Dariusz R. Kowalski. Meeting the deadline: on the complexity of fault-tolerant continuous gossip. Distributed Computing, 24(5):223–244, December 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-011-0144-6.

Georgiou:2020:CG

[GGK20] Chryssis Georgiou, Seth Gilbert, and Dariusz R. Kowalski. Confidential gossip. Distributed Computing, 33(5):367–392, October 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-019-00367-x.

Galcik:2013:EBR

[GGL13] František Galčík, Leszek Ga̧sieniec, and Andrzej Lingas. Efficient broadcasting in radio networks with long-range interference. Distributed Computing, 26(1):59–74, February 2013. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
012-0176-6.

Ghosh:1997:SSA

[GGP97] Sukumar Ghosh, Arobinda Gupta, and Sriram V. Pemmaraju. A self-stabilizing algorithm for the maximum flow problem. Distributed Computing, 10(4):167–180, July 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s004460050034.

Ghosh:2007:FCS

[GGP07] Sukumar Ghosh, Arobinda Gupta, and Sriram V. Pemmaraju. Fault-containing self-stabilizing distributed protocols. Distributed Computing, 20(1):53–73, July 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-007-0032-2.

Gramoli:2018:TST

[GGT18] Vincent Gramoli, Rachid Guerraoui, and Vasileios Trigonakis. TM^{2}C: a software transactional memory for many-cores. Distributed Computing, 31(5):367–388, October 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
017-0310-6.

Gouda:2007:A

[GH07] Mohamed G. Gouda and F. Furman Haddix. The alternator. Distributed Computing, 20(1):21–28, July 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-007-0033-1.

Ghaffari:2015:RBR

[GHK15] Mohsen Ghaffari, Bernhard Haeupler, and Majid Khabbazian. Randomized broadcast in radio networks with collision detection. Distributed Computing, 28(6):407–422, December 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0230-7.

Ghaffari:2021:IDC

[GHM21] Mohsen Ghaffari, Juho Hirvonen, and Yannic Maus. Improved distributed Δ-coloring. Distributed Computing, 34(4):239–258, August 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-021-00397-4.

Guerraoui:2009:WFD

[GHN09] Rachid Guerraoui, Maurice Herlihy, and Calvin Newport. On the weakest failure detector ever. Distributed Computing, 21(5): 353–366, February 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-009-0079-3.

Guerraoui:2010:MCT

[GHS10] Rachid Guerraoui, Thomas A. Henzinger, and Vasu Singh. Model checking transactional memories. Distributed Computing, 22(3): 129–145, March 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-009-0092-6.

Goos:2017:LLB

[GHS17] Mika Göös, Juho Hirvonen, and Jukka Suomela. Linear-in-Δ lower bounds in the LOCAL model. Distributed Computing, 30 (5):325–338, October 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-015-0245-8.

Gotte:2023:TOC

[GHSW23] Thorsten Götte, Kristian Hinnenthal, Christian Scheideler, and Julian Werthmann. Time-optimal construction of overlay networks. Distributed Computing, 36(3):313–347, September 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-023-00442-4.

Ghaffari:2020:IDD

[GHU20] Mohsen Ghaffari, Juho Hirvonen, and Jara Uitto. Improved distributed degree splitting and edge coloring. Distributed Computing, 33(3–4):293–310, June 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-018-00346-8.

Groote:2001:AAW

[GHV01] Jan Friso Groote, Wim H. Hesselink, and Rogier Vermeulen. An algorithm for the asynchronous Write-All problem based on process collision. Distributed Computing, 14(2):75–81, April 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008930.

Golab:2012:REI

[GHW12] Wojciech Golab, Vassos Hadzilacos, and Philipp Woelfel. RMRefficient implementations of comparison primitives using read and write operations. Distributed Computing, 25(2):109–162, May 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-011-0150-8.

Gjessing:1988:SVM

[Gje88] Stein Gjessing. Semantics and verification of monitors and systems of monitors and processes. Distributed Computing, 2(4): 190–200, December 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01872845.

Ghosh:1993:SSA

[GK93] Sukumar Ghosh and Mehmet Hakan Karaata. A self-stabilizing algorithm for coloring planar graphs. Distributed Computing, 7 (1):55–59, November 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02278856.

Guerraoui:2008:FDT

[GK08] Rachid Guerraoui and Petr Kouznetsov. Failure detectors as type boosters. Distributed Computing, 20(5):343–358, February 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-007-0043-z.

Gafni:2011:SCN

[GK11] Eli Gafni and Petr Kuznetsov. On set consensus numbers. Distributed Computing, 24(3–4):??, November 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
011-0142-8.

Guerraoui:2008:WFD

[GKK08] Rachid Guerraoui, Michał Kapałka, and Petr Kouznetsov. The weakest failure detectors to boost obstruction-freedom. Distributed Computing, 20(6):415–433, April 2008. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/s00446-007-
0046-9.

Ganesh:2007:PCS

[GKM07] A. J. Ganesh, A.-M. Kermarrec, and L. Massoulié. Peer counting and sampling in overlay networks based on random walks. Distributed Computing, 20(4):267–278, November 2007. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
007-0027-z.

Gasieniec:2008:TES

[GKS08] Leszek Ga̧sieniec, Erez Kantor, and Chang Su. Time efficient kshot broadcasting in known topology radio networks. Distributed Computing, 21(2):117–127, July 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-008-0058-0.

Guerraoui:2022:CNC

[GKS22a] Rachid Guerraoui, Petr Kuznetsov, and Dragos-Adrian Seredinschi. The consensus number of a cryptocurrency. Distributed Computing, 35(1):1–15, February 2022. CODEN DICOEB.

ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-021-00399-2.

See correction [GKS22b].

Guerraoui:2022:CCN

[GKS22b] Rachid Guerraoui, Petr Kuznetsov, and Dragos-Adrian Seredinschi. Correction to: The consensus number of a cryptocurrency. Distributed Computing, 35(1):17, February 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-

022-00422-0. See [GKS22a].

Gafni:2003:DP

[GL03] Eli Gafni and Leslie Lamport. Disk Paxos. Distributed Computing, 16(1):1–20, February 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-002-0070-8.

Gilbert:2010:RRR

[GLS10] Seth Gilbert, Nancy A. Lynch, and Alexander A. Shvartsman. Rambo: a robust, reconfigurable atomic memory service for dynamic networks. Distributed Computing, 23(4):225–272, December 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452

(electronic). URL https://link.springer.com/article/10.
1007/s00446-010-0117-1.

Ghaffari:2012:LEU

[GLS12] Mohsen Ghaffari, Nancy Lynch, and Srikanth Sastry. Leader election using loneliness detection. Distributed Computing, 25(6): 427–450, December 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0172-x.

Gold:1987:DAC

[GM87] Yaron I. Gold and Shlomo Moran. Distributed algorithms for constructing a minimum-weight spanning tree in a broadcast network. Distributed Computing, 2(3):139–148, September 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01782774.

Glasgow:1989:ONM

[GM89] Janice I. Glasgow and Glenn H. MacEwen. An operator net model for distributed systems. Distributed Computing, 3(4):196–209, December 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01784888.

Goldstein:2004:WRP

[GM04] Darin Goldstein and Nick Meyer. The wake up and report problem is time-equivalent to the firing squad synchronization problem. Distributed Computing, 17(1):21–31, February 2004. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
003-0095-7.

Goren:2021:OTS

[GM21] Guy Goren and Yoram Moses. Optimistically tuning synchronous Byzantine consensus: another win for null messages. Distributed Computing, 34(5):395–410, October 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-021-00393-8.

Gorain:2023:FSD

[GMP23] Barun Gorain, Avery Miller, and Andrzej Pelc. Four shades of deterministic leader election in anonymous networks. Distributed Computing, 36(4):419–449, December 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-023-00451-3.

Goubault:2018:GCV

[GMT18] Éric Goubault, Samuel Mimram, and Christine Tasson. Geometric and combinatorial views on asynchronous computability. Distributed Computing, 31(4):289–316, August 2018. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
018-0328-4.

Gupta:2007:DND

[GND07] Indranil Gupta, Mahvesh Nagda, and Christo Frank Devaraj. The design of novel distributed protocols from differential equations. Distributed Computing, 20(2):95–114, August 2007. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
007-0024-2.

Gilbert:2017:WYS

[GNZ17] Seth Gilbert, Calvin Newport, and Chaodong Zheng. Who are you? Secure identities in single hop ad hoc networks. Distributed Computing, 30(2):103–125, April 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-016-0280-0.

Goldman:1991:HCL

[Gol91] Kenneth J. Goldman. Highly concurrent logically synchronous multicast. Distributed Computing, 4(4):189–207, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01784720.

Goldreich:2003:CCP

[Gol03] Oded Goldreich. Cryptography and cryptographic protocols. Distributed Computing, 16(2–3):177–199, September 2003. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
002-0077-1.

Gong:1995:ENA

[Gon95] Li Gong. Efficient network authentication protocols: Lower bounds and optimal implementations. Distributed Computing, 9 (3):131–145, December 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050015.

Gorla:2010:TPC

[Gor10] Daniele Gorla. A taxonomy of process calculi for distribution and mobility. Distributed Computing, 23(4):273–299, December 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-010-0120-6.

Gouda:1986:NJB

[Gou86] Mohamed G. Gouda. A new journal is born. Distributed Computing, 1(1):3–4, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01843565.

Gavoille:2003:CLD

[GP03] Cyril Gavoille and David Peleg. Compact and localized distributed data structures. Distributed Computing, 16(2–3):111– 120, September 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-002-0073-5.

Ganesh:2021:OEP

[GP21] Chaya Ganesh and Arpita Patra. Optimal extension protocols for Byzantine broadcast and agreement. Distributed Computing, 34 (1):59–77, February 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-020-00384-1.

Gasieniec:2007:FCK

[GPX07] Leszek Ga̧sieniec, David Peleg, and Qin Xin. Faster communication in known topology radio networks. Distributed Computing, 19(4):289–300, March 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0011-z.

Guerraoui:2007:AFT

[GR07] Rachid Guerraoui and Eric Ruppert. Anonymous and faulttolerant shared-memory computing. Distributed Computing, 20 (3):165–177, October 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-007-0042-0.

Golab:2019:RME

[GR19] Wojciech Golab and Aditya Ramaraju. Recoverable mutual exclusion. Distributed Computing, 32(6):535–564, December 2019. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
019-00364-0.

Graf:1999:CSC

[Gra99] Susanne Graf. Characterization of a sequentially consistent memory and verification of a cache memory by abstraction. Distributed Computing, 12(2–3):75–90, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050059.

Georgiou:2004:CSI

[GRS04] Chryssis Georgiou, Alexander Russell, and Alex A. Shvartsman. The complexity of synchronous iterative Do-All with crashes. Distributed Computing, 17(1):47–63, February 2004. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
003-0099-3.

Goldreich:1991:CCP

[GS91] Oded Goldreich and Liuba Shrira. On the complexity of computation in the presence of link failures: the case of a ring. Distributed Computing, 5(3):121–131, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF02252955.

Gray:1998:LAM

[GS98] James W. Gray III and Paul F. Syverson. A logical approach to multilevel security of probabilistic systems. Distributed Computing, 11(2):73–90, April 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050043.

Goos:2014:NST

[GS14] Mika Göös and Jukka Suomela. No sublogarithmic-time approximation scheme for bipartite vertex cover. Distributed Computing, 27(6):435–443, December 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-013-0194-z.

Garg:2007:TME

[GSM07] Vijay K. Garg, Chakarat Skawratananond, and Neeraj Mittal. Timestamping messages and events in a distributed system using synchronous communication. Distributed Computing, 19(5– 6):387–402, April 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0018-5.

Groselj:1991:DSC

[GT91] Bojan Grošelj and Carl Tropper. The distributed simulation of clustered processes. Distributed Computing, 4(3):111– 121, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01798958.

Guerraoui:2002:NBA

[Gue02] Rachid Guerraoui. Non-blocking atomic commit in asynchronous distributed systems with failure detectors. Distributed Computing, 15(1):17–25, January 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s446-002-8027-4.

Guerraoui:2006:EI

[Gue06] Rachid Guerraoui. Editorial: Introduction. Distributed Computing, 18(3):177, February 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-005-0146-3.

Guerraoui:2012:SSS

[Gue12] Rachid Guerraoui. Special section with selected papers from PODC 2010. Distributed Computing, 25(2):163, May 2012. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
012-0166-8.

Guerraoui:2010:RQS

[GV10] Rachid Guerraoui and Marko Vukolić. Refined quorum systems. Distributed Computing, 23(1):1–42, September 2010. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
010-0103-7.

Gillani:2021:LBN

[GVB21] Iqra Altaf Gillani, Pooja Vyavahare, and Amitabha Bagchi. Lower bounds for in-network computation of arbitrary functions. Distributed Computing, 34(3):181–193, June 2021. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
021-00394-7.

Goren:2023:SCH

[GVM23] Guy Goren, Shay Vargaftik, and Yoram Moses. Stochastic coordination in heterogeneous load balancing systems. Distributed Computing, 36(3):253–275, September 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-022-00437-7.

Giakkoupis:2019:ERT

[GW19] George Giakkoupis and Philipp Woelfel. Efficient randomized test-and-set implementations. Distributed Computing, 32(6): 565–586, December 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-019-00349-z.

Hadzilacos:1987:CRB

[Had87] Vassos Hadzilacos. Connectivity requirements for Byzantine agreement under restricted types of failures. Distributed Computing, 2(2):95–103, June 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01667081.

Hadzilacos:1997:I

[Had97] Vassos Hadzilacos. Introduction. Distributed Computing, 10 (2):63, February 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050024.

Hajek:1991:BET

[Haj91] B. Hajek. Bounds on evacuation time for deflection routing. Distributed Computing, 5(1):1–6, June 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/BF02311228.

Halpern:2000:NKB

[Hal00] Joseph Y. Halpern. A note on knowledge-based programs and specifications. Distributed Computing, 13(3):145–153, July 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008916.

Havlicek:2000:COW

[Hav00] John Havlicek. Computable obstructions to wait-free computability. Distributed Computing, 13(2):59–83, April 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050068.

Huang:1993:SSD

[HC93] Shing-Tsaan Huang and Nian-Shing Chen. Self-stabilizing depthfirst token circulation on networks. Distributed Computing, 7 (1):61–66, November 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02278857.

Herman:1992:SSR

[Her92] Ted Herman. Self-stabilization: randomness to reduce space. Distributed Computing, 6(2):95–98, September 1992. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02252680.

Herman:2000:SME

[Her00] Ted Herman. Superstabilizing mutual exclusion. Distributed Computing, 13(1):1–17, January 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050001.

Hesselink:1994:WFL

[Hes94] Wim H. Hesselink. Wait-free linearization with an assertional proof. Distributed Computing, 8(2):65–80, October 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02280829.

Hesselink:1995:WFL

[Hes95] Wim H. Hesselink. Wait-free linearization with a mechanical proof. Distributed Computing, 9(1):21–36, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01784240.

Hesselink:1999:PUB

[Hes99] Wim H. Hesselink. Progress under bounded fairness. Distributed Computing, 12(4):197–207, September 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s004460050066.

Halpern:1989:MKA

[HF89] Joseph Y. Halpern and Ronald Fagin. Modelling knowledge and action in distributed systems. Distributed Computing, 3(4): 159–177, December 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01784885.

Hesselink:2001:WFC

[HG01] Wim H. Hesselink and Jan Friso Groote. Wait-free concurrent memory management by Create and Read until Deletion (CaRuD). Distributed Computing, 14(1):31–39, January 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008924.

He:1987:ASP

[HH87] Jifeng He and C. A. R. Hoare. Algebraic specification and proof of a distributed recovery algorithm. Distributed Computing, 2(1):1– 12, March 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01786251.

Huang:2003:AUR

[HHH03] Shing-Tsaan Huang, Ying-Sung Huang, and Su-Shen Hung. Alternators on uniform rings of odd size. Distributed Computing, 16 (4):263–268, December 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-003-0087-7.

Halldorsson:2010:SSG

[HHM10] Magnús M. Halldórsson, Joseph Y. Halpern, and Vahab S. Mirrokni. On spectrum sharing games. Distributed Computing, 22 (4):235–248, May 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-010-0098-0.

Hadzilacos:2022:ARR

[HHT22] Vassos Hadzilacos, Xing Hu, and Sam Toueg. On atomic registers and randomized consensus in M&M systems. Distributed Computing, 35(1):81–103, February 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00405-7.

Hierons:2012:OCP

[Hie12] Robert M. Hierons. Overcoming controllability problems in distributed testing from an input output transition system. Distributed Computing, 25(1):63–81, March 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-011-0153-5.

Haeupler:2021:LCS

[HIZ21] Bernhard Haeupler, Taisuke Izumi, and Goran Zuzic. Lowcongestion shortcuts without embedding. Distributed Computing, 34(1):79–90, February 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-020-00383-2.

Hella:2015:WMD

[HJV15] Lauri Hella, Matti Järvisalo, and Jonni Virtema. Weak models of distributed computing, with connections to modal logic. Distributed Computing, 28(1):31–53, February 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0202-3.

Hendler:2009:BWC

[HK09] Danny Hendler and Shay Kutten. Bounded-wait combining: constructing robust and high-throughput shared objects. Distributed Computing, 21(6):405–431, March 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-009-0078-4.

Halldorsson:2018:CLI

[HK18] Magnús M. Halldórsson and Christian Konrad. Computing large independent sets in a single round. Distributed Computing, 31 (1):69–82, February 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0298-y.

Holtby:2008:LBS

[HKK08] Dan Holtby, Bruce M. Kapron, and Valerie King. Lower bound for scalable Byzantine agreement. Distributed Computing, 21(4): 239–248, October 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0069-x.

Halldorsson:2018:DBP

[HKR18] Magnús M. Halldórsson, Sven Köhler, and Dror Rawitz. Distributed backup placement in networks. Distributed Computing, 31(2):83–98, April 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0299-x.

Halldorsson:2019:DAS

[HKR19] Magnús M. Halldórsson, Sven Köhler, and Dror Rawitz. Distributed approximation of k-service assignment. Distributed Computing, 32(1):27–40, February 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-017-0321-3.

Hilbers:1989:DFM

[HL89] Peter A. J. Hilbers and Johan J. Lukkien. Deadlock-free message routing in multicomputer networks. Distributed Computing, 3(4): 178–186, December 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01784886.

Huang:1999:SSC

[HL99] Shing-Tsaan Huang and Tzong-Jye Liu. Self-stabilizing 2^{m}-clock for unidirectional rings of odd size. Distributed Computing, 12 (1):41–46, March 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050054.

Hendler:2006:DSN

[HLS06] Danny Hendler, Yossi Lev, and Nir Shavit. A dynamic-sized nonblocking work stealing deque. Distributed Computing, 18(3): 189–207, February 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0144-5.

Halpern:2004:UCK

[HM04] Joseph Y. Halpern and Yoram Moses. Using counterfactuals in knowledge-based programming. Distributed Computing, 17 (2):91–106, August 2004. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-004-0108-1.

Halldorsson:2016:NOB

[HM16] Magnús M. Halldórsson and Pradipta Mitra. Nearly optimal bounds for distributed wireless scheduling in the SINR model. Distributed Computing, 29(2):77–88, April 2016. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0222-7.

Hierons:2012:IRT

[HMN12] Robert M. Hierons, Mercedes G. Merayo, and Manuel Núñez. Implementation relations and test generation for systems with distributed interfaces. Distributed Computing, 25(1):35–62, March 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-011-0149-1.

Hierons:2014:TIR

[HMN14] Robert M. Hierons, Mercedes G. Merayo, and Manuel Núñez. Timed implementation relations for the distributed test architecture. Distributed Computing, 27(3):181–201, June 2014. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0208-5.

Helary:2000:CBP

[HMR00] J.-M. Hélary, A. Mostefaoui, and M. Raynal. Communicationbased prevention of useless checkpoints in distributed computations. Distributed Computing, 13(1):29–43, January 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050003.

Hoare:1986:P

[Hoa86] C. A. R. Hoare. Preface. Distributed Computing, 1(1):1, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01843564.

Hassin:2001:SCN

[HP01] Yehuda Hassin and David Peleg. Sparse communication networks and efficient routing in the plane. Distributed Computing, 14(4): 205–215, December 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460100054.

Herlihy:2005:TBS

[HP05] M. Herlihy and L. D. Penso. Tight bounds for k-set agreement with limited-scope failure detectors. Distributed Computing, 18 (2):157–166, November 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0141-8.

Halpern:2010:KBA

[HP10] Joseph Y. Halpern and Sabina Petride. A knowledge-based analysis of global function computation. Distributed Computing, 23 (3):197–224, November 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-010-0111-7.

Hegeman:2015:SLD

[HP15] James W. Hegeman and Sriram V. Pemmaraju. Sub-logarithmic distributed algorithms for metric facility location. Distributed Computing, 28(5):351–374, October 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-015-0243-x.

Helary:1994:TCD

[HR94] Jean-Michel Hélary and Michel Raynal. Towards the construction of distributed detection programs, with an application to distributed termination. Distributed Computing, 7(3):137–147, March 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02277858.

Hurfin:1999:SFA

[HR99] Michel Hurfin and Michel Raynal. A simple and fast asynchronous consensus protocol based on a weak failure detector. Distributed Computing, 12(4):209–223, September 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s004460050067.

Halpern:2005:KTA

[HR05] Joseph Y. Halpern and Aleta Ricciardi. A knowledge-theoretic analysis of uniform distributed coordination and failure detectors. Distributed Computing, 17(3):223–236, March 2005. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
004-0109-0.

Herlihy:2013:TDA

[HR13] Maurice Herlihy and Sergio Rajsbaum. The topology of distributed adversaries. Distributed Computing, 26(3):173–192, June 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-013-0189-9.

Huang:2020:CCA

[HRZ20] Zengfeng Huang, Bozidar Radunovic, and Qin Zhang. Communication complexity of approximate maximum matching in the message-passing model. Distributed Computing, 33(6):515–531, December 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-020-00371-6.

Halpern:1991:CSP

[HS91] Joseph Y. Halpern and Ichiro Suzuki. Clock synchronization and the power of broadcasting. Distributed Computing, 5(2): 73–82, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02259749.

Herlihy:2007:DTM

[HS07] Maurice Herlihy and Ye Sun. Distributed transactional memory for metric-space networks. Distributed Computing, 20(3):195–208, October 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-007-0037-x.

Hendler:2008:SVC

[HS08] Danny Hendler and Nir Shavit. Solo-valency and the cost of coordination. Distributed Computing, 21(1):43–54, June 2008. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
007-0051-z.

Hoefer:2012:DNI

[HS12] Martin Hoefer and Siddharth Suri. Dynamics in network interaction games. Distributed Computing, 25(5):359–370, October 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-012-0171-y.

Hayes:2012:FGD

[HST12] Thomas P. Hayes, Jared Saia, and Amitabh Trehan. The forgiving graph: a distributed data structure for low stretch under adversarial attack. Distributed Computing, 25(4):261–278, August 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-012-0160-1.

Herlihy:1996:LCN

[HSW96] Maurice Herlihy, Nir Shavit, and Orli Waarts. Linearizable counting networks. Distributed Computing, 9(4):??, February 1996.

CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050019.

Herlihy:2006:SSS

[HT06] Maurice Herlihy and Srikanta Tirthapura. Self-stabilizing smoothing and balancing networks. Distributed Computing, 18(5):345– 357, April 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-005-0130-y.

Huang:2009:DEC

[HT09] Shing-Tsaan Huang and Chi-Hung Tzeng. Distributed edge coloration for bipartite networks. Distributed Computing, 22(1):3–14, April 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-009-0082-8.

Hussak:2023:TAF

[HT23] Walter Hussak and Amitabh Trehan. Termination of amnesiac flooding. Distributed Computing, 36(2):193–207, June 2023. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
023-00448-y.

Hu:2024:ISR

[HT24] Xing Hu and Sam Toueg. On implementing SWMR registers from SWSR registers in systems with Byzantine failures. Distributed Computing, 37(2):145–175, June 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-024-00465-5.

Hayrapetyan:2007:NPG

[HTW07] Ara Hayrapetyan, Éva Tardos, and Tom Wexler. A network pricing game for selfish traffic. Distributed Computing, 19(4):255–266, March 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-006-0020-y.

Hierons:2008:CSD

[HU08] R. M. Hierons and H. Ural. Checking sequences for distributed test architectures. Distributed Computing, 21(3):223– 238, September 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0062-4.

Huang:2006:EFC

[Hua06] Tetz C. Huang. An efficient fault-containing self-stabilizing algorithm for the shortest path problem. Distributed Computing, 19 (2):149–161, October 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0001-1.

Huang:1997:SST

[HW97] Shing-Tsaan Huang and Lih-Chyau Wuu. Self-stabilizing token circulation in uniform networks. Distributed Computing, 10 (4):181–187, July 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050035.

Hendler:2011:RME

[HW11] Danny Hendler and Philipp Woelfel. Randomized mutual exclusion with sub-logarithmic RMR-complexity. Distributed Computing, 24(1):3–19, September 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0128-6.

Halldorsson:2019:LMC

[HWY19] Magnús M. Halldórsson, Yuexuan Wang, and Dongxiao Yu. Leveraging multiple channels in ad hoc networks. Distributed Computing, 32(2):159–172, April 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-018-0329-3.

Issa:2020:EHT

[IFR20] Shady Issa, Pascal Felber, and Paolo Romano. Extending hardware transactional memory capacity via rollback-only transactions and suspend/resume. Distributed Computing, 33(3–4):327– 348, June 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-019-00363-1.

Israeli:1993:BTS

[IL93] Amos Israeli and Ming Li. Bounded time-stamps. Distributed Computing, 6(4):205–209, July 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF02242708.

Ibaraki:1995:OCR

[INK95] Toshihide Ibaraki, Hiroshi Nagamochi, and Tsunehiko Kameda. Optimal coteries for rings and related networks. Distributed Computing, 8(4):191–201, June 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF02242737.

Ilcinkas:2009:CMD

[INS09] David Ilcinkas, Nicolas Nisse, and David Soguet. The cost of monotonicity in distributed graph searching. Distributed Computing, 22(2):117–127, October 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-009-0089-1.

Ingram:2013:LEA

[IRW13] Rebecca Ingram, Tsvetomira Radeva, and Jennifer L. Welch. A leader election algorithm for dynamic networks with causal clocks. Distributed Computing, 26(2):75–97, April 2013. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0184-1.

Jalote:1989:FTP

[Jal89] Pankaj Jalote. Fault tolerant processes. Distributed Computing, 3 (4):187–195, December 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01784887.

Jain:2007:BSR

[JLC07] Kamal Jain, László Lovász, and Philip A. Chou. Building scalable and robust peer-to-peer overlay networks for broadcasting using network coding. Distributed Computing, 19(4):301–311, March 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-006-0014-9.

Junqueira:2010:TPS

[JMP10] Flavio P. Junqueira, Keith Marzullo, and Lucia Draque Penso. Threshold protocols in survivor set systems. Distributed Computing, 23(2):135–149, October 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0107-3.

Jonker:1992:FGC

[Jon92] Jan E. Jonker. On-the-fly garbage collection for several mutators. Distributed Computing, 5(4):187–199, April 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02277666.

Jonsson:1994:FAT

[Jon94] Bengt Jonsson. A fully abstract trace model for dataflow and asynchronous networks. Distributed Computing, 7(4):197–212, May 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02280834.

Josephs:1988:SBA

[Jos88] Mark B. Josephs. A state-based approach to communicating processes. Distributed Computing, 3(1):9–18, March 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01788563.

Joung:2000:AGM

[Jou00] Yuh-Jzer Joung. Asynchronous group mutual exclusion. Distributed Computing, 13(4):189–206, November 2000. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/PL00008918.

Joung:2002:CTP

[Jou02] Yuh-Jzer Joung. The congenial talking philosophers problem in computer networks. Distributed Computing, 15(3):155–175, July 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460100069.

Joung:2010:QSG

[Jou10] Yuh-Jzer Joung. On quorum systems for group resources allocation. Distributed Computing, 22(3):197–214, March 2010. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
010-0094-4.

Jonsson:1999:PRU

[JPR99] Bengt Jonsson, Amir Pnueli, and Camilla Rump. Proving refinement using transduction. Distributed Computing, 12(2–3):129– 149, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050062.

Janssen:1999:CAS

[JPZ99] Wil Janssen, Mannes Poel, and Job Zwiers. The compositional approach to sequential consistency and lazy caching. Distributed Computing, 12(2–3):105–127, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460050061.

Jia:2002:EDA

[JRS02] Lujun Jia, Rajmohan Rajaraman, and Torsten Suel. An efficient distributed algorithm for constructing small dominating sets. Distributed Computing, 15(4):193–205, December 2002. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
002-0078-0.

Jeavons:2016:FNS

[JSX16] Peter Jeavons, Alex Scott, and Lei Xu. Feedback from nature: simple randomised distributed algorithms for maximal independent set selection and greedy colouring. Distributed Computing, 29(5):377–393, October 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-016-0269-8.

Jayanti:2021:CDS

[JT21] Siddhartha V. Jayanti and Robert E. Tarjan. Concurrent disjoint set union. Distributed Computing, 34(6):413–436, December 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-020-00388-x.

Jahja:2020:SLB

[JYZ20] Irvan Jahja, Haifeng Yu, and Yuda Zhao. Some lower bounds in dynamic networks with oblivious adversaries. Distributed Computing, 33(1):1–40, February 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-019-00360-4.

Kim:2006:NME

[KA06] Yong-Jik Kim and J. H. Aderson. Nonatomic mutual exclusion with local spinning. Distributed Computing, 19(1):19–61, September 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452

(electronic). URL https://link.springer.com/article/10.
1007/s00446-006-0003-z.

Kim:2007:AME

[KA07] Yong-Jik Kim and James H. Anderson. Adaptive mutual exclusion with local spinning. Distributed Computing, 19(3):197–236, January 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-006-0009-6.

Kim:2012:TCL

[KA12] Yong-Jik Kim and James H. Anderson. A time complexity lower bound for adaptive mutual exclusion. Distributed Computing, 24 (6):271–297, January 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-011-0152-6.

Karaata:1999:SSA

[KC99] Mehmet Hakan Karaata and Pranay Chaudhuri. A self-stabilizing algorithm for bridge finding. Distributed Computing, 12(1):47–53, March 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050055.

Keidar:2011:SIS

[Kei11] Idit Keidar. Special issue with selected papers from DISC 2009. Distributed Computing, 24(3–4):??, November 2011. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
011-0140-x.

Kessels:1995:CDC

[Kes95] Joep L. W. Kessels. Calculational derivation of a counter with bounded response time and bounded power dissipation. Distributed Computing, 8(3):143–149, March 1995. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02242716.

Kash:2012:OSS

[KFH12] Ian A. Kash, Eric J. Friedman, and Joseph Y. Halpern. Optimizing scrip systems: crashes, altruists, hoarders, sybils and collusion. Distributed Computing, 25(5):335–357, October 2012. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
012-0170-z.

Karayel:2022:SEC

[KG22] Emin Karayel and Edgar Gonzàlez. Strong eventual consistency of the collaborative editing framework WOOT. Distributed Computing, 35(2):145–164, April 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00414-6.

Ketchpel:1999:SCA

[KGM99] Steven P. Ketchpel and Hector Garcia-Molina. A sound and complete algorithm for distributed commerce transactions. Distributed Computing, 12(1):13–29, March 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s004460050052.

Kant:1996:DPS

[KHvB96] Christian Kant, Teruo Higashino, and Gregor von Bochmann. Deriving protocol specifications from service specifications written in LOTOS. Distributed Computing, 10(1):29–47, July 1996. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050022.

Kapus-Kolar:1999:CDP

[KK99] Monika Kapus-Kolar. Comments on deriving protocol specifications from service specifications written in LOTOS. Distributed Computing, 12(4):175–177, September 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s004460050064.

Korman:2007:DVM

[KK07] Amos Korman and Shay Kutten. Distributed verification of minimum spanning trees. Distributed Computing, 20(4):253– 266, November 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-007-0025-1.

Kawamura:2015:FPM

[KK15] Akitoshi Kawamura and Yusuke Kobayashi. Fence patrolling by mobile agents with distinct speeds. Distributed Computing, 28 (2):147–154, April 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0226-3.

Kitamura:2021:LCS

[KKI21] Naoki Kitamura, Hirotaka Kitagawa, and Taisuke Izumi. Lowcongestion shortcut and graph parameters. Distributed Computing, 34(5):349–365, October 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00401-x.

Korman:2015:FCS

[KKM15] Amos Korman, Shay Kutten, and Toshimitsu Masuzawa. Fast and compact self-stabilizing verification, computation, and fault detection of an MST. Distributed Computing, 28(4):253–295, August 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-015-0242-y.

Klonowski:2019:ODA

[KKM19] Marek Klonowski, Dariusz R. Kowalski, and Jarosław Mirek. Ordered and delayed adversaries and how to work against them on a shared channel. Distributed Computing, 32(5):379–403, October 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-018-0341-7.

Kirkpatrick:2024:PBA

[KKN^{+}24] David Kirkpatrick, Irina Kostitsyna, Alfredo Navarra, Giuseppe Prencipe, and Nicola Santoro. On the power of bounded asynchrony: convergence by autonomous robots with limited visibility. Distributed Computing, 37(3):279–308, September 2024. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
024-00463-7.

Korman:2010:PLS

[KKP10] Amos Korman, Shay Kutten, and David Peleg. Proof labeling schemes. Distributed Computing, 22(4):215–233, May 2010. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
010-0095-3.

Kesselman:2010:PMQ

[KKS10] Alex Kesselman, Kirill Kogan, and Michael Segal. Packet mode and QoS algorithms for buffered crossbar switches with FIFO queuing. Distributed Computing, 23(3):163–175, November 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-010-0114-4.

Klasing:2017:MAR

[KKS17] Ralf Klasing, Adrian Kosowski, and Thomas Sauerwald. The multi-agent rotor-router on the ring: a deterministic alternative to parallel random walks. Distributed Computing, 30(2):127–148, April 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-016-0282-y.

Khan:2012:EDA

[KKT12] Maleq Khan, Fabian Kuhn, and Kunal Talwar. Efficient distributed approximation algorithms via probabilistic tree embeddings. Distributed Computing, 25(3):189–205, June 2012. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
012-0157-9.

Kuhn:2011:AML

[KLN11] Fabian Kuhn, Nancy Lynch, and Calvin Newport. The abstract MAC layer. Distributed Computing, 24(3–4):187–206, November 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-010-0118-0.

Kekkonen-Moneta:2002:TO

[KM02] Synnöve Kekkonen-Moneta. Torus orientation. Distributed Computing, 15(1):39–48, January 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s446-002-8029-y.

Korman:2007:GCL

[Kor07] Amos Korman. General compact labeling schemes for dynamic trees. Distributed Computing, 20(3):179–193, October 2007. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
007-0035-z.

Kozlov:2012:GIA

[Koz12] Dmitry N. Kozlov. Gathering identical autonomous systems on a circle using stigmergy. Distributed Computing, 25(6):461–472, December 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-012-0178-4.

Keller:1986:SDN

[KP86] Robert M. Keller and Prakash Panangaden. Semantics of digital networks containing indeterminate modules. Distributed Computing, 1(4):235–245, December 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01660035.

Katz:1992:VDP

[KP92] Shmuel Katz and Doron Peled. Verification of distributed programs using representative interleaving sequences. Distributed Computing, 6(2):107–120, September 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF02252682.

Katz:1993:SSE

[KP93] Shmuel Katz and Kenneth J. Perry. Self-stabilizing extensions for meassage-passing systems. Distributed Computing, 7 (1):17–26, November 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02278852.

Kowalski:2005:BUA

[KP05] Dariusz R. Kowalski and Andrzej Pelc. Broadcasting in undirected ad hoc radio networks. Distributed Computing, 18(1):43– 57, July 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-005-0126-7.

Kowalski:2007:ODB

[KP07] Dariusz R. Kowalski and Andrzej Pelc. Optimal deterministic broadcasting in known topology radio networks. Distributed Computing, 19(3):185–195, January 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-006-0007-8.

Khan:2008:FDA

[KP08a] Maleq Khan and Gopal Pandurangan. A fast distributed approximation algorithm for minimum spanning trees. Distributed Computing, 20(6):391–402, April 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-007-0047-8.

Korman:2008:CSD

[KP08b] Amos Korman and David Peleg. Compact separator decompositions in dynamic trees and applications to labeling schemes. Distributed Computing, 21(2):141–161, July 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-008-0061-5.

Kuznetsov:2023:PAA

[KPPT23] Petr Kuznetsov, Yvonne-Anne Pignolet, Pavel Ponomarev, and Andrei Tonkikh. Permissionless and asynchronous asset transfer. Distributed Computing, 36(3):349–371, September 2023. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
023-00449-x.

Kipnis:2010:CDS

[KPS10] Alex Kipnis and Boaz Patt-Shamir. On the complexity of distributed stable matching with small messages. Distributed Computing, 23(3):151–161, November 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0105-5.

Kleinberg:2011:LBR

[KPT11] Robert Kleinberg, Georgios Piliouras, and Éva Tardos. Load balancing without regret in the bulletin board model. Distributed Computing, 24(1):21–29, September 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-011-0129-5.

King:2018:RCJ

[KPY18] Valerie King, Seth Pettie, and Maxwell Young. A resourcecompetitive jamming defense. Distributed Computing, 31(6): 419–439, November 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0313-3.

Kessels:1990:DSD

[KR90] Joep L. W. Kessels and Martin Rem. Designing systolic, distributed buffers with bounded response time. Distributed Computing, 4(1):37–43, March 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01783664.

Kanellakis:1992:EPA

[KS92a] Paris C. Kanellakis and Alex A. Shvartsman. Efficient parallel algorithms can be made robust. Distributed Computing, 5(4):201– 217, April 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02277667.

Kurki-Suonio:1992:OSJ

[KS92b] Reino Kurki-Suonio. Operational specification with joint actions: Serializable databases. Distributed Computing, 6(1):19–37, July 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02276639.

Kshemkalyani:1998:NSC

[KS98] Ajay D. Kshemkalyani and Mukesh Singhal. Necessary and sufficient conditions on information for causal message ordering and their optimal implementation. Distributed Computing, 11(2):91– 111, April 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050044.

Kranakis:2001:DCO

[KS01] Evangelos Kranakis and Nicola Santoro. Distributed computing on oriented anonymous hypercubes with faulty components. Distributed Computing, 14(3):185–189, July 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/PL00008936.

Kurki-Suonio:2003:ASI

[KS03] R. Kurki-Suonio. Action systems in incremental and aspectoriented modeling. Distributed Computing, 16(2–3):201–217, September 2003. CODEN DICOEB. ISSN 0178-2770 (print),

1432-0452 (electronic). URL https://link.springer.com/
article/10.1007/s00446-003-0090-z.

Kshemkalyani:1998:CAD

[Ksh98] Ajay D. Kshemkalyani. Causality and atomicity in distributed computations. Distributed Computing, 11(4):169–189, October 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050048.

Kshemkalyani:2004:PLC

[Ksh04] Ajay D. Kshemkalyani. The power of logical clock abstractions. Distributed Computing, 17(2):131–150, August 2004. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
003-0105-9.

Korman:2013:TML

[KSV13] Amos Korman, Jean-Sébastien Sereni, and Laurent Viennot. Toward more localized local algorithms: removing assumptions concerning global knowledge. Distributed Computing, 26(5–6): 289–308, October 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0174-8.

Kuhn:2010:TWC

[KSW10] Fabian Kuhn, Stefan Schmid, and Roger Wattenhofer. Towards worst-case churn resistant peer-to-peer systems. Distributed Computing, 22(4):249–267, May 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0099-z.

Kohler:2012:FCS

[KT12] Sven Köhler and Volker Turau. Fault-containing self-stabilization in asynchronous systems with constant fault-gap. Distributed Computing, 25(3):207–224, June 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0155-3.

Kuznetsov:2022:ARB

[KT22] Petr Kuznetsov and Andrei Tonkikh. Asynchronous reconfiguration with Byzantine failures. Distributed Computing, 35(6): 477–502, December 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-022-00421-1.

Kaldewaij:1992:ROF

[KU92] Anne Kaldewaij and Jan Tijmen Udding. Rank order filters and priority queues. Distributed Computing, 6(2):99–105, September 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02252681.

Kutten:2006:ISI

[Kut06] Shay Kutten. Introduction to the special issue PODC’2004. Distributed Computing, 18(4):233, March 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-005-0152-5.

Kung:1986:MIP

[KW86] H. T. Kung and Jon A. Webb. Mapping image processing operations onto a linear systolic machine. Distributed Computing, 1 (4):246–257, December 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01660036.

Kuhn:2005:CTD

[KW05] Fabian Kuhn and Roger Wattenhofer. Constant-time distributed dominating set approximation. Distributed Computing, 17(4): 303–310, May 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-004-0112-5.

Kupavskii:2021:LBS

[KW21] Andrey Kupavskii and Emo Welzl. Lower bounds for searching robots, some faulty. Distributed Computing, 34(4):229–237, August 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-019-00358-y.

Koufogiannakis:2011:DAC

[KY11] Christos Koufogiannakis and Neal E. Young. Distributed algorithms for covering, packing and maximum weighted matching. Distributed Computing, 24(1):45–63, September 2011. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
011-0127-7.

Lynch:1992:UMP

[LA92] Nancy A. Lynch and Hagit Attiya. Using mappings to prove timing properties. Distributed Computing, 6(2):121–139, September 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02252683.

Lamport:1986:ICa

[Lam86a] Leslie Lamport. On interprocess communication. Distributed Computing, 1(2):77–85, June 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01786227.

Lamport:1986:ICb

[Lam86b] Leslie Lamport. On interprocess communication. Distributed Computing, 1(2):86–101, June 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01786228.

Lamport:1990:TAD

[Lam90] Leslie Lamport. A theorem on atomicity in distributed algorithms. Distributed Computing, 4(2):59–68, June 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01786631.

Lamport:1992:CLA

[Lam92] Leslie Lamport. Critique of the Lake Arrowhead Three. Distributed Computing, 6(1):65–71, July 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF02276641.

Lamport:2000:FH

[Lam00] Leslie Lamport. Fairness and hyperfairness. Distributed Computing, 13(4):239–245, November 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/PL00008921.

Lamport:2003:AFS

[Lam03] Leslie Lamport. Arbitration-free synchronization. Distributed Computing, 16(2–3):219–237, September 2003. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/s00446-002-
0076-2.

Lamport:2006:FP

[Lam06a] Leslie Lamport. Fast Paxos. Distributed Computing, 19(2): 79–103, October 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0005-x.

Lamport:2006:LBA

[Lam06b] Leslie Lamport. Lower bounds for asynchronous consensus. Distributed Computing, 19(2):104–125, October 2006. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
006-0155-x.

Lamport:2008:IDT

[Lam08] Leslie Lamport. Implementing dataflow with threads. Distributed Computing, 21(3):163–181, September 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-008-0065-1.

Lengauer:1991:TSC

[LBH91] Christian Lengauer, Michael Barnett, and Duncan G. Hudson. Towards systolizing compilation. Distributed Computing, 5(1):7– 24, June 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02311229.

Lowe:1999:UCV

[LD99] Gavin Lowe and Jim Davies. Using CSP to verify sequential consistency. Distributed Computing, 12(2–3):91–103, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050060.

Li:2005:OSB

[LDB05] Ninghui Li, Wenliang Du, and Dan Boneh. Oblivious signaturebased envelope. Distributed Computing, 17(4):293–302, May 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-004-0116-1.

Lo:2000:PSO

[LH00] Wai-Kau Lo and Vassos Hadzilacos. On the power of shared object types to implement one-resilient Consensus. Distributed Computing, 13(4):219–238, November 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/PL00008920.

Lin:1988:MCC

[Lin88] Huai-An Lin. A methodology for constructing communication protocols with multiple concurrent functions. Distributed Computing, 3(1):23–40, March 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01788565.

Lisper:1991:CTC

[Lis91] Björn Lisper. Computing transitive closure on systolic arrays of fixed size. Distributed Computing, 5(3):133–144, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02252956.

Liskov:1993:PUS

[Lis93] Barbara Liskov. Practical uses of synchronized clocks in distributed systems. Distributed Computing, 6(4):211–219, July 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02242709.

Livesey:1993:NCA

[Liv93] Mike Livesey. A note on consistency in asynchronous multicaches. Distributed Computing, 7(2):111–114, December 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02280840.

Lin:2019:AFT

[LKJ19] Yiyan Lin, Sandeep Kulkarni, and Arshad Jhumka. Automation of fault-tolerant graceful degradation. Distributed Computing, 32 (1):1–25, February 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0319-x.

Lewko:2015:CAA

[LL15] Allison Lewko and Mark Lewko. On the complexity of asynchronous agreement against powerful adversaries. Distributed Computing, 28(6):377–389, December 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-014-0224-5.

Ladkin:1999:LCT

[LLR99] Peter Ladkin, Leslie Lamport, and Denis Roegel. Lazy caching in TLA. Distributed Computing, 12(2–3):151–174, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050063.

Lenzen:2017:SCT

[LLR17] Christoph Lenzen, Nancy Lynch, and Tsvetomira Radeva. Searching without communicating: tradeoffs between performance and selection complexity. Distributed Computing, 30(3):169–191, June 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-016-0283-x.

Lundy:1991:SAD

[LM91] G. M. Lundy and Raymond E. Miller. Specification and analysis of a data transfer protocol using systems of communicating machines. Distributed Computing, 5(3):145–157, December 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02252957.

Lubitch:1995:CSN

[LM95] Ronit Lubitch and Shlomo Moran. Closed schedulers: a novel technique for analyzing asynchronous protocols. Distributed Computing, 8(4):203–210, June 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF02242738.

Lv:2016:FRI

[LM16] Yuezhou Lv and Thomas Moscibroda. Fair and resilient incentive tree mechanisms. Distributed Computing, 29(1):1–16, February 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-015-0250-y.

Livesey:2007:DDT

[LMM07] M. J. Livesey, R. Morrison, and D. S. Munro. The Doomsday distributed termination detection protocol. Distributed Computing, 19(5–6):419–431, April 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0013-x.

Li:2006:CMR

[LMP06] Xiaozhou Li, Jayadev Misra, and C. Greg. Plaxton. Concurrent maintenance of rings. Distributed Computing, 19(2):126–148, October 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-006-0008-7.

Levi:2021:PTP

[LMR21] Reut Levi, Moti Medina, and Dana Ron. Property testing of planarity in the CONGEST model. Distributed Computing, 34 (1):15–32, February 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-020-00382-3.

Ladan-Mozes:2008:OAL

[LMS08] Edya Ladan-Mozes and Nir Shavit. An optimistic approach to lock-free FIFO queues. Distributed Computing, 20(5):323–341, February 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-007-0050-0.

Lotker:2006:DMC

[LPSP06] Zvi Lotker, Boaz Patt-Shamir, and David Peleg. Distributed MST for constant diameter graphs. Distributed Computing, 18(6):453– 460, June 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-005-0127-6.

Lenzen:2019:DDC

[LPSP19] Christoph Lenzen, Boaz Patt-Shamir, and David Peleg. Distributed distance computation and routing with small messages. Distributed Computing, 32(2):133–157, April 2019. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
018-0326-6.

Lenzen:2013:DMD

[LPW13] Christoph Lenzen, Yvonne-Anne Pignolet, and Roger Wattenhofer. Distributed minimum dominating set approximations in restricted families of graphs. Distributed Computing, 26(2):119– 137, April 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-013-0186-z.

Lenzen:2019:NOS

[LR19] Christoph Lenzen and Joel Rybicki. Near-optimal self-stabilising counting and firing squads. Distributed Computing, 32(4): 339–360, August 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-018-0342-6.

Lam:1992:SMS

[LS92] Simon S. Lam and A. Udaya Shankar. Specifying modules to satisfy interfaces: A state transition system approach. Distributed Computing, 6(1):39–63, July 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF02276640.

Lynch:2013:SID

[LS13] Nancy A. Lynch and Alexander A. Shvartsman. Special issue on DISC 2010. Distributed Computing, 26(3):139, June 2013. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0191-2.

Lozes:2010:SEL

[LV10] Étienne Lozes and Jules Villard. A spatial equational logic for the applied π-calculus. Distributed Computing, 23(1):61–83, September 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452

(electronic). URL https://link.springer.com/article/10.
1007/s00446-010-0112-6.

Liskov:1986:SDP

[LW86] Barbara Liskov and William Weihl. Specifications of distributed programs. Distributed Computing, 1(2):102–118, June 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01786229.

Lee:2005:RRI

[LW05] Hyunyoung Lee and Jennifer L. Welch. Randomized registers and iterative algorithms. Distributed Computing, 17(3):209–221, March 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-004-0106-3.

Lenzen:2016:TBP

[LW16] Christoph Lenzen and Roger Wattenhofer. Tight bounds for parallel randomized load balancing. Distributed Computing, 29 (2):127–142, April 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0225-4.

Lynch:2003:SPP

[Lyn03] Nancy Lynch. Some perspectives on PODC. Distributed Computing, 16(2–3):71–74, September 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-002-0072-6.

Martin:1986:CCP

[Mar86] Alain J. Martin. Compiling communicating processes into delayinsensitive VLSI circuits. Distributed Computing, 1(4):226–234, December 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01660034.

Marzullo:2005:I

[Mar05] Keith Marzullo. Introduction. Distributed Computing, 18(1):59, July 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-005-0131-x.

Mattern:1987:ADT

[Mat87] Friedemann Mattern. Algorithms for distributed termination detection. Distributed Computing, 2(3):161–175, September 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01782776.

Moses:1986:CHO

[MDH86] Yoram Moses, Danny Dolev, and Joseph Y. Halpern. Cheating husbands and other stories: A case study of knowledge, action, and communication. Distributed Computing, 1(3):167– 176, September 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01661170.

Meyer:1989:SST

[MdV89] J.-J. C. Meyer and E. P. de Vink. Step semantics for “true” concurrency with recursion. Distributed Computing, 3(3):130– 145, September 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01784023.

Meldal:1991:CAS

[Mel91] Sigurd Meldal. A complete axiomatic semantics of spawning. Distributed Computing, 5(3):159–174, December 1991. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02252958.

Merritt:1993:I

[Mer93] Micheal Merritt. Introduction. Distributed Computing, 7(1):1, November 1993. CODEN DICOEB. ISSN 0178-2770 (print),

1432-0452 (electronic). URL https://link.springer.com/
article/10.1007/BF02278850.

Merritt:1999:I

[Mer99] Michael Merritt. Introduction. Distributed Computing, 12 (2–3):55–56, May 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050056.

Mittal:2004:FMS

[MG04] Neeraj Mittal and Vijay K. Garg. Finding missing synchronization in a distributed computation using controlled reexecution. Distributed Computing, 17(2):107–130, August 2004. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-003-0104-x.

Mittal:2005:TAC

[MG05] Neeraj Mittal and Vijay K. Garg. Techniques and applications of computation slicing. Distributed Computing, 17(3):251–277, March 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-004-0117-0.

Mendes:2015:MAB

[MHG15] Hammurabi Mendes, Maurice Herlihy, and Vijay K. Garg. Multidimensional agreement in Byzantine systems. Distributed Computing, 28(6):423–441, December 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0240-5.

Merayo:2018:PTA

[MHN18] Mercedes G. Merayo, Robert M. Hierons, and M. Núñez. Passive testing with asynchronous communications and timestamps. Distributed Computing, 31(5):327–342, October 2018. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
017-0308-0.

Milosevic:2014:TPT

[MHS14] Zarko Milosevic, Martin Hutle, and André Schiper. Tolerating permanent and transient value faults. Distributed Computing, 27 (1):55–77, February 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-013-0199-7.

Michail:2018:TDC

[Mic18] Othon Michail. Terminating distributed construction of shapes and patterns in a fair solution of automata. Distributed Computing, 31(5):343–365, October 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-017-0309-z.

Miller:1987:CSS

[Mil87] Raymond E. Miller. The construction of self-synchronizing finite state protocols. Distributed Computing, 2(2):104–112, June 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01667082.

Mitzenmacher:1997:CTP

[Mit97] Michael Mitzenmacher. Constant time per edge is optimal on rooted tree networks. Distributed Computing, 10(4):189–197, July 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050036.

Mashreghi:2021:BMS

[MK21] Ali Mashreghi and Valerie King. Broadcast and minimum spanning tree with o(m) messages in the asynchronous CONGEST model. Distributed Computing, 34(4):283–299, August 2021. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
020-00387-y.

Maheshwari:1997:CCD

[ML97] Umesh Maheshwari and Barbara Liskov. Collecting cyclic distributed garbage by controlled migration. Distributed Computing, 10(2):79–86, February 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050026.

Manduchi:2000:AVC

[MM00] G. Manduchi and M. Moro. Automatic verification for a class of distributed systems. Distributed Computing, 13(3):127–143, July 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/PL00008915.

Mizrahi:2008:CCC

[MM08] Tal Mizrahi and Yoram Moses. Continuous consensus via common knowledge. Distributed Computing, 20(5):305–321, February 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-007-0049-6.

Malkhi:2000:SRM

[MMR00] Dahlia Malkhi, Michael Merritt, and Ohad Rodeh. Secure reliable multicast protocols in a WAN. Distributed Computing, 13 (1):19–28, January 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050002.

Moser:1993:NSC

[MMSA93] Louise E. Moser, P. M. Melliar-Smith, and Vivek Agrawala. Necessary and sufficient conditions for broadcast consensus protocols. Distributed Computing, 7(2):75–85, December 1993. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02280837.

Malkhi:2003:OSB

[MMT03] Dahlia Malkhi, Michael Merritt, and Gadi Taubenfeld. Objects shared by Byzantine processes. Distributed Computing, 16 (1):37–48, February 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-002-0075-3.

Mavronicolas:2008:SCV

[MMT08] Marios Mavronicolas, Michael Merritt, and Gadi Taubenfeld. Sequentially consistent versus linearizable counting networks. Distributed Computing, 21(4):249–269, October 2008. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
008-0066-0.

Mertzios:2017:DMN

[MNS17] George B. Mertzios, Sotiris E. Nikoletseas, and Paul G. Spirakis. Determining majority in networks with local interactions and very small local memory. Distributed Computing, 30(1):1–16, February 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-016-0277-8.

Moniz:2011:RCH

[MNV11] Henrique Moniz, Nuno Ferreira Neves, and Paulo Veríssimo. Randomization can be a healer: consensus with dynamic omission failures. Distributed Computing, 24(3–4):165–175, November 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-010-0116-2.

Moir:2001:LPU

[Moi01] Mark Moir. Laziness pays! Using lazy synchronization mechanisms to improve non-blocking constructions. Distributed Computing, 14(4):193–204, December 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s004460100063.

Moses:2018:ISI

[Mos18] Yoram Moses. Introduction to the special issue of papers from DISC 2015. Distributed Computing, 31(4):255, August 2018. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
018-0337-3.

Mizoguchi:2012:SCS

[MOY12] Ryu Mizoguchi, Hirotaka Ono, and Masafumi Yamashita. On space complexity of self-stabilizing leader election in mediated population protocol. Distributed Computing, 25(6):451–460, December 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-012-0173-9.

Miller:2016:TVC

[MP16] Avery Miller and Andrzej Pelc. Time versus cost tradeoffs for deterministic rendezvous in networks. Distributed Computing, 29 (1):51–64, February 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-015-0253-8.

Mitzenmacher:2018:SMP

[MP18] Michael Mitzenmacher and Rasmus Pagh. Simple multi-party set reconciliation. Distributed Computing, 31(6):441–453, November 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-017-0316-0.

Mansour:2004:OSS

[MPSL04] Yishay Mansour, Boaz Patt-Shamir, and Ofer Lapid. Optimal smoothing schedules for real-time streams. Distributed Computing, 17(1):77–89, February 2004. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-003-0101-0.

Malkhi:1998:BQS

[MR98] Dahlia Malkhi and Michael Reiter. Byzantine quorum systems. Distributed Computing, 11(4):203–213, October 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050050.

MacKenzie:2003:DCS

[MR03] Philip MacKenzie and Michael K. Reiter. Delegation of cryptographic servers for capture-resilient devices. Distributed Computing, 16(4):307–327, December 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-003-0098-4.

Mostefaoui:2004:CBC

[MRR04] Achour Mostéfaoui, Sergio Rajsbaum, and Matthieu Roy. Condition-based consensus solvability: a hierarchy of conditions and efficient protocols. Distributed Computing, 17(1):1–20, February 2004. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452

(electronic). URL https://link.springer.com/article/10.
1007/s00446-003-0093-9.

Mostefaoui:2006:SCB

[MRR06] Achour Mostefaoui, Sergio Rajsbaum, and Michel Raynal. Synchronous condition-based consensus. Distributed Computing, 18 (5):325–343, April 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0148-1.

Malewicz:2006:DSD

[MRS06] Grzegorz Malewicz, Alexander Russell, and Alexander A. Shvartsman. Distributed scheduling for disconnected cooperation. Distributed Computing, 18(6):409–420, June 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
005-0149-0.

Mostefaoui:2008:CPR

[MRT08] Achour Mostefaoui, Sergio Rajsbaum, and Corentin Travers. On the computability power and the robustness of set agreementoriented failure detector classes. Distributed Computing, 21(3): 201–222, September 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0064-2.

Metivier:2011:OBC

[MRZ11] Y. Métivier, J. M. Robson, and A. Zemmari. An optimal bit complexity randomized distributed MIS algorithm. Distributed Computing, 23(5–6):331–340, April 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0121-5.

Maxemchuk:1989:PVC

[MS89] N. F. Maxemchuk and Krishan Sabnani. Probabilistic verification of communication protocols. Distributed Computing, 3(3): 118–129, September 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01784022.

Marzullo:1994:EDC

[MS94] Keith Marzullo and Laura S. Sabel. Efficient detection of a class of stable properties. Distributed Computing, 8(2):81–91, October 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02280830.

Mukund:1997:KTL

[MS97] Madhavan Mukund and Milind Sohoni. Keeping track of the latest gossip in a distributed system. Distributed Computing, 10 (3):137–148, March 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050031.

Minsky:2003:TMG

[MS03] Yaron M. Minsky and Fred B. Schneider. Tolerating malicious gossip. Distributed Computing, 16(1):49–68, February 2003. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
002-0082-4.

Mavronicolas:2010:IRS

[MS10] Marios Mavronicolas and Thomas Sauerwald. The impact of randomization in smoothing networks. Distributed Computing, 22 (5–6):381–411, August 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-009-0087-3.

Moser:2014:RFT

[MS14] Heinrich Moser and Ulrich Schmid. Reconciling fault-tolerant distributed algorithms and real-time computing. Distributed Computing, 27(3):203–230, June 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-013-0204-1.

Michail:2016:SEL

[MS16] Othon Michail and Paul G. Spirakis. Simple and efficient local codes for distributed stable network construction. Distributed Computing, 29(3):207–237, June 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-015-0257-4.

Michail:2018:HMC

[MS18] Othon Michail and Paul G. Spirakis. How many cooks spoil the soup? Distributed Computing, 31(6):455–469, November 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-017-0317-z.

Michail:2022:DCR

[MSS22] Othon Michail, George Skretas, and Paul G. Spirakis. Distributed computation and reconfiguration in actively dynamic networks. Distributed Computing, 35(2):185–206, April 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
021-00415-5.

Merritt:1993:KSM

[MT93] Michael Merritt and Gadi Taubenfeld. Knowledge in shared memory systems. Distributed Computing, 7(2):99–109, December 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02280839.

Merritt:1994:ARO

[MT94] Michael Merritt and Gadi Taubenfeld. Atomic m-register operations. Distributed Computing, 7(4):213–221, May 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02280835.

Malkhi:2007:CVV

[MT07] Dahlia Malkhi and Doug Terry. Concise version vectors in WinFS. Distributed Computing, 20(3):209–219, October 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-007-0044-y.

Maus:2022:LL

[MT22] Yannic Maus and Tigran Tonoyan. Linial for lists. Distributed Computing, 35(6):533–546, December 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-022-00424-y.

Mittal:2007:FOT

[MVP07] Neeraj Mittal, S. Venkatesan, and Sathya Peri. A family of optimal termination detection algorithms. Distributed Computing, 20(2):141–162, August 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-007-0031-3.

Moscibroda:2008:CUR

[MW08] Thomas Moscibroda and Roger Wattenhofer. Coloring unstructured radio networks. Distributed Computing, 21(4):271–284, October 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-008-0070-4.

Myoupo:1993:MDP

[Myo93] Jean Frédéric Myoupo. Mapping dynamic programming onto modular linear systolic arrays. Distributed Computing, 6(3):165– 179, April 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02242705.

Nesetril:2016:DLT

[NdM16] J. Nešetřil and P. Ossona de Mendez. A distributed low treedepth decomposition algorithm for bounded expansion classes. Distributed Computing, 29(1):39–49, February 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
015-0251-x.

Nguyen:1986:MTP

[NDO86] Van Nguyen, Alan Demers, and Susan Owicki. A model and temporal proof system for networks of processes. Distributed Computing, 1(1):7–25, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01843567.

Naor:2024:ELR

[NK24] Oded Naor and Idit Keidar. Expected linear round synchronization: the missing link for linear Byzantine SMR. Distributed Computing, 37(1):19–33, March 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-023-00459-9.

Newport:2011:MRN

[NL11] Calvin Newport and Nancy Lynch. Modeling radio networks. Distributed Computing, 24(2):101–118, October 2011. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
011-0135-7.

Nielsen:1990:BNE

[NRT90] M. Nielsen, G. Rozenberg, and P. S. Thiagarajan. Behavioural notions for elementary net systems. Distributed Computing, 4(1):45– 57, March 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01783665.

Nanongkai:2022:ECC

[NS22] Danupon Nanongkai and Michele Scquizzato. Equivalence classes and conditional hardness in massively parallel computations. Distributed Computing, 35(2):165–183, April 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
021-00418-2.

Neiger:1993:CKC

[NT93] Gil Neiger and Mark R. Tuttle. Common knowledge and consistent simultaneous coordination. Distributed Computing, 6(3):181– 192, April 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02242706.

Naor:2005:SDQ

[NW05] Moni Naor and Udi Wieder. Scalable and dynamic quorum systems. Distributed Computing, 17(4):311–322, May 2005. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
004-0114-3.

Nguyen:2021:PRS

[NYD21] Duong Nguyen, Sorrachai Yingchareonthawornchai, and Murat Demirbas. Precision, recall, and sensitivity of monitoring partially synchronous distributed programs. Distributed Computing, 34(5):319–348, October 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-021-00402-w.

Okun:2008:RSM

[OBG08] Michael Okun, Amnon Barak, and Eli Gafni. Renaming in synchronous message passing systems with Byzantine failures. Distributed Computing, 20(6):403–413, April 2008. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
007-0045-x.

Orda:1996:DSP

[OR96] Ariel Orda and Raphael Rom. Distributed shortest-path protocols for time-dependent networks. Distributed Computing, 10(1):49– 62, July 1996. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050023.

Ogierman:2018:SCM

[ORZ18] Adrian Ogierman, Andrea Richa, and Jin Zhang. Sade: competitive MAC under adversarial SINR. Distributed Computing, 31(3):241–254, June 2018. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-017-0307-1.

Pachl:1991:ATP

[Pac91] Jan Pachl. Analysis of toggle protocols. Distributed Computing, 5(1):25–35, June 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02311230.

Panconesi:2015:SIS

[Pan15] Alessandro Panconesi. Special issue with selected papers from PODC 2012. Distributed Computing, 28(1):1, February 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0239-y.

Parter:2017:VFT

[Par17] Merav Parter. Vertex fault tolerant additive spanners. Distributed Computing, 30(5):357–372, October 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-015-0252-9.

Patra:2014:ABA

[PCR14] Arpita Patra, Ashish Choudhury, and C. Pandu Rangan. Asynchronous Byzantine agreement with optimal resilience. Distributed Computing, 27(2):111–146, April 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0200-5.

Peleska:1991:DVF

[Pel91] Jan Peleska. Design and verification of fault tolerant systems with CSP. Distributed Computing, 5(2):95–106, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02259751.

Pelc:2007:AAA

[Pel07] Andrzej Pelc. Activating anonymous ad hoc radio networks. Distributed Computing, 19(5–6):361–371, April 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
007-0021-5.

Pelc:2008:SID

[Pel08] Andrzej Pelc. Special issue on DISC 07. Distributed Computing, 21(2):85, July 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0063-3.

Peleg:2013:SID

[Pel13] David Peleg. Special issue on DISC 2011. Distributed Computing, 26(4):193, August 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-013-0193-0.

Pettie:2010:DAU

[Pet10] Seth Pettie. Distributed algorithms for ultrasparse spanners and linear size skeletons. Distributed Computing, 22(3):147–166, March 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-009-0091-7.

Pixley:1988:IGC

[Pix88] Carl Pixley. An incremental garbage collection algorithm for multi-mutator systems. Distributed Computing, 3(1):41–50, March 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF01788566.

Pandya:1991:PLC

[PJ91] Partosh K. Pandya and Mathai Joseph. P — a logic — a compositional proof system for distributed programs. Distributed Computing, 5(1):37–54, June 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02311231.

Perrin:2022:EWF

[PMCP22] Matthieu Perrin, Achour Mostéfaoui, and Ludmila Courtillat-Piazza. Extending the wait-free hierarchy to multi-threaded systems. Distributed Computing, 35(4):375–398, August 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
022-00425-x.

Parter:2024:NOD

[PP24] Merav Parter and Asaf Petruschka. Near-optimal distributed computation of small vertex cuts. Distributed Computing, 37 (2):67–88, June 2024. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-023-00455-z.

Pagourtzis:2017:RBR

[PPS17] Aris Pagourtzis, Giorgos Panagiotakos, and Dimitris Sakavalas. Reliable broadcast with respect to topology knowledge. Distributed Computing, 30(2):87–102, April 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-016-0279-6.

Panconesi:1998:RNU

[PPV98] Alessandro Panconesi, Marina Papatriantafilou, and Paul Vitányi. Randomized naming using wait-free shared variables. Distributed Computing, 11(3):113–124, August 1998. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050045.

Pagli:2015:GCT

[PPV15] Linda Pagli, Giuseppe Prencipe, and Giovanni Viglietta. Getting close without touching: near-gathering for autonomous mobile robots. Distributed Computing, 28(5):333–349, October 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
015-0248-5.

Panconesi:2001:SSD

[PR01] Alessandro Panconesi and Romeo Rizzi. Some simple distributed algorithms for sparse networks. Distributed Computing, 14(2):97– 100, April 2001. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/PL00008932.

Pandurangan:2016:DSH

[PRT16] Gopal Pandurangan, Peter Robinson, and Amitabh Trehan. DEX: self-healing expanders. Distributed Computing, 29(3):163–185, June 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-015-0258-3.

Paliwoda:1991:ISS

[PS91] K. Paliwoda and J. W. Sanders. An incremental specification of the sliding-window protocol. Distributed Computing, 5(2): 83–94, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02259750.

Pedone:2002:HMS

[PS02] F. Pedone and A. Schiper. Handling message semantics with Generic Broadcast protocols. Distributed Computing, 15(2):97– 107, April 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460100061.

Panconesi:2010:FPD

[PS10a] Alessandro Panconesi and Mauro Sozio. Fast primal-dual distributed algorithms for scheduling and matching problems. Distributed Computing, 22(4):269–283, May 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-010-0100-x.

Patt-Shamir:2010:SIP

[PS10b] Boaz Patt-Shamir. Special issue on PODC 2008. Distributed Computing, 22(5–6):285, August 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0109-1.

Pogosyants:2000:VRC

[PSL00] Anna Pogosyants, Roberto Segala, and Nancy Lynch. Verification of the randomized consensus algorithm of Aspnes and Herlihy: a case study. Distributed Computing, 13(3):155–186, July 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
PL00008917.

Patt-Shamir:2008:ADT

[PSS08] Boaz Patt-Shamir and Allon Shafrir. Approximate distributed top-k queries. Distributed Computing, 21(1):1–22, June 2008. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
008-0055-3.

Pignolet:2013:MSW

[PST13] Yvonne-Anne Pignolet, Stefan Schmid, and Gilles Tredan. Misleading stars: what cannot be measured in the Internet? Distributed Computing, 26(4):209–222, August 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
013-0183-2.

Pignolet:2015:ATD

[PST15] Yvonne Anne Pignolet, Stefan Schmid, and Gilles Tredan. Adversarial topology discovery in network virtualization environments: a threat for ISPs? Distributed Computing, 28(2):91–109, April 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-014-0217-4.

Pignolet:2016:ULB

[PST16] Yvonne Anne Pignolet, Stefan Schmid, and Gilles Tredan. Upper and lower bounds for deterministic broadcast in powerline communication networks. Distributed Computing, 29(4):239–250, August 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-016-0263-1.

Pike:2012:FDE

[PSW12] Scott M. Pike, Srikanth Sastry, and Jennifer L. Welch. Failure detectors encapsulate fairness. Distributed Computing, 25 (4):313–333, August 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-012-0164-x.

Panangaden:1992:CCK

[PT92] Prakash Panangaden and Kim Taylor. Concurrent common knowledge: defining agreement for asynchronous systems. Distributed Computing, 6(2):73–93, September 1992. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02252679.

Pandurangan:2014:XLS

[PT14] Gopal Pandurangan and Amitabh Trehan. Xheal: a localized self-healing algorithm using expanders. Distributed Computing, 27(1):39–54, February 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-013-0192-1.

Peleg:1997:CWC

[PW97] David Peleg and Avishai Wool. Crumbling walls: a class of practical and efficient quorum systems. Distributed Computing, 10 (2):87–97, February 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s004460050027.

Pnueli:1986:VMP

[PZ86] Amir Pnueli and Lenore Zuck. Verification of multiprocess probabilistic protocols. Distributed Computing, 1(1):53–72, March 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01843570.

Rajopadhye:1989:SSA

[Raj89] Sanjay V. Rajopadhye. Synthesizing systolic arrays with control signals from recurrence equations. Distributed Computing, 3(2):88–105, June 1989. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01558666.

Rajsbaum:2005:I

[Raj05] Sergio Rajsbaum. Introduction. Distributed Computing, 17(4): 291, May 2005. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-004-0118-z.

Roussopoulos:2006:PLB

[RB06] Mema Roussopoulos and Mary Baker. Practical load balancing for content requests in peer-to-peer networks. Distributed Computing, 18(6):421–434, June 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0150-7.

Ramanathan:2007:RLE

[RFS07] Murali Krishna Ramanathan, Ronaldo A. Ferreira, and Wojciech Szpankowski. Randomized leader election. Distributed Computing, 19(5–6):403–418, April 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-007-0022-4.

Retvari:2013:CPR

[RGB13] Gábor Rétvári, András Gulyás, and József J. Bíró. Compact policy routing. Distributed Computing, 26(5–6):309–320, October 2013. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-012-0181-9.

Rhee:1998:MAR

[Rhe98] Injong Rhee. A modular algorithm for resource allocation. Distributed Computing, 11(3):157–168, August 1998. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050047.

Richa:2013:CTM

[RSZ13] Andrea Richa, Christian Scheideler, and Jin Zhang. Competitive throughput in multi-hop wireless networks despite adaptive jamming. Distributed Computing, 26(3):159–171, June 2013. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
012-0180-x.

Roditty:2016:CLS

[RT16] Liam Roditty and Roei Tov. Close to linear space routing schemes. Distributed Computing, 29(1):65–74, February 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-015-0256-5.

Rukoz:1991:HDD

[Ruk91] Marta Rukoz. Hierarchical deablock detection for nested transactions. Distributed Computing, 4(3):123–129, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01798959.

Shafer:1995:IHC

[SA95] Keith Shafer and Mohan Ahuja. Implementation of hierarchical F -channels for high-performance distributed computing. Distributed Computing, 8(4):211–218, June 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF02242739.

Sayeed:1995:OAA

[SAAAA95] Hasan M. Sayeed, Marwan Abu-Amara, and Hosame Abu-Amara. Optimal asynchronous agreement and leader election algorithm for complete networks with Byzantine faulty links. Distributed Computing, 9(3):147–156, December 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s004460050016.

Sangiorgi:2012:CTT

[San12] Davide Sangiorgi. Concurrency theory: timed automata, testing, program synthesis. Distributed Computing, 25(1):3–4, March 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-011-0156-2.

Singh:1995:HAM

[SB95] Gurdip Singh and Arthur J. Bernstein. A highly asynchronous minimum spanning tree protocol. Distributed Computing, 8(3): 151–161, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02242717.

Sharma:2012:WBG

[SB12] Gokarna Sharma and Costas Busch. Window-based greedy contention management for transactional memory: theory and practice. Distributed Computing, 25(3):225–248, June 2012. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
012-0159-7.

Sharma:2014:DTM

[SB14] Gokarna Sharma and Costas Busch. Distributed transactional memory for general networks. Distributed Computing, 27(5): 329–362, October 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0214-7.

Sanchez:2015:SEM

[SCA15] Daniel Sánchez, Juan M. Cebrián, and Juan L. Aragón. Softerror mitigation by means of decoupled transactional memory threads. Distributed Computing, 28(2):75–90, April 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
014-0215-6.

Schneider:1990:CG

[Sch90] Fred B. Schneider. Changing of the guard. Distributed Computing, 4(1):1, March 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01783660.

Schneider:1992:I

[Sch92] Fred B. Schneider. Introduction. Distributed Computing, 6(1): 1–3, July 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02276637.

Schneider:1995:CDC

[Sch95] Fred B. Schneider. Changes in distributed computing. Distributed Computing, 9(1):1, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01784238.

Schiper:1997:ECAa

[Sch97a] André Schiper. Early consensus in an asynchronous system with a weak failure detector. Distributed Computing, 10(3):149–157, March 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050032.

Schiper:1997:ECAb

[Sch97b] André Schiper. Early consensus in an asynchronous system with a weak failure detector. Distributed Computing, 10(4):198, July 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050037.

Schiper:2006:DGC

[Sch06] André Schiper. Dynamic group communication. Distributed Computing, 18(5):359–374, April 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-005-0129-4.

Sarma:2015:EDC

[SDP15] Atish Das Sarma, Michael Dinitz, and Gopal Pandurangan. Efficient distributed computation of distance sketches in networks. Distributed Computing, 28(5):309–320, October 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
015-0246-7.

Shavit:2002:I

[Sha02] Nir Shavit. Introduction. Distributed Computing, 15(4):191, December 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-002-0086-0.

Shafiei:2019:NBP

[Sha19] Niloufar Shafiei. Non-blocking Patricia tries with replace operations. Distributed Computing, 32(5):423–442, October 2019. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
019-00347-1.

Severson:2021:CCD

[SHD21] Eric E. Severson, David Haley, and David Doty. Composable computation in discrete chemical reaction networks. Distributed Computing, 34(6):437–461, December 2021. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-020-00378-z.

Singhal:1991:CDF

[Sin91] Mukesh Singhal. A class of deadlock-free Maekawa-type algorithms for mutual exclusion in distributed systems. Distributed Computing, 4(3):131–138, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01798960.

Singh:1997:ELE

[Sin97] Gurdip Singh. Efficient leader election using sense of direction. Distributed Computing, 10(3):159–165, March 1997. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050033.

Singh:1999:CBS

[Sin99] Gurdip Singh. Constraint-based structuring of network protocols. Distributed Computing, 12(1):1–12, March 1999. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050051.

Shankar:1987:TDD

[SL87] A. Udaya Shankar and Simon S. Lam. Time-dependent distributed systems: proving safety, liveness and real-time properties. Distributed Computing, 2(2):61–79, June 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF01667079.

Slivkins:2007:DEO

[Sli07] Aleksandrs Slivkins. Distance estimation and object location via rings of neighbors. Distributed Computing, 19(4):313–333, March 2007. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-006-0015-8.

Schwarz:1994:DCR

[SM94] Reinhard Schwarz and Friedemann Mattern. Detecting causal relationships in distributed computations: In search of the holy grail. Distributed Computing, 7(3):149–174, March 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/BF02277859.

Schiper:1994:SSP

[SS94] André Schiper and Alain Sandoz. Strong stable properties in distributed systems. Distributed Computing, 8(2):93–103, October 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02280831.

Schmid:2001:HRF

[SS01] Ulrich Schmid and Klaus Schossmaier. How to reconcile faulttolerant interval intersection with the Lipschitz condition. Distributed Computing, 14(2):101–111, April 2001. CODEN DI-COEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL

https://link.springer.com/article/10.1007/PL00008927.

Saad:2017:TEE

[SS17] George Saad and Jared Saia. A theoretical and empirical evaluation of an algorithm for self-healing computation. Distributed Computing, 30(6):391–412, December 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-016-0290-y.

Srikanth:1987:SAB

[ST87] T. K. Srikanth and Sam Toueg. Simulating authenticated broadcasts to derive simple fault-tolerant algorithms. Distributed Computing, 2(2):80–94, June 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01667080.

Shavit:1997:STM

[ST97] Nir Shavit and Dan Touitou. Software transactional memory. Distributed Computing, 10(2):99–116, February 1997. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050028.

Shavit:2016:CRD

[ST16] Nir Shavit and Gadi Taubenfeld. The computability of relaxed data structures: queues and stacks as examples. Distributed Computing, 29(5):395–407, October 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-016-0272-0.

Stoller:2000:DGP

[Sto00] Scott D. Stoller. Detecting global predicates in distributed systems with clocks. Distributed Computing, 13(2):85–98, April 2000. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050069.

Su:2019:DNB

[SV19] Lili Su and Nitin H. Vaidya. Defending non-Bayesian learning against adversarial attacks. Distributed Computing, 32(4): 277–289, August 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-018-0336-4.

Schneider:2010:OMI

[SW10] Johannes Schneider and Roger Wattenhofer. An optimal maximal independent set algorithm for bounded-independence graphs. Distributed Computing, 22(5–6):349–361, August 2010. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
010-0097-1.

Siek:2022:LUO

[SW22] Konrad Siek and Paweł T. Wojciechowski. Last-use opacity: a strong safety property for transactional memory with prerelease support. Distributed Computing, 35(3):265–301, June 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
022-00420-2.

Tang:1994:MV

[Tan94] Jian Tang. On multilevel voting. Distributed Computing, 8 (1):39–58, August 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02283569.

Taubenfeld:2010:SID

[Tau10] Gadi Taubenfeld. Special issue on DISC 2008. Distributed Computing, 23(2):85–86, October 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-010-0113-5.

Taubenfeld:2014:TSB

[Tau14] Gadi Taubenfeld. Tight space bounds for ℓ-exclusion. Distributed Computing, 27(3):165–179, June 2014. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0207-6.

Taubenfeld:2015:SIP

[Tau15] Gadi Taubenfeld. Special issue on PODC 2013. Distributed Computing, 28(6):375–376, December 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-015-0260-9.

Taubenfeld:2020:CSP

[Tau20] Gadi Taubenfeld. The computational structure of progress conditions and shared objects. Distributed Computing, 33(2):103–123, April 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-019-00356-0.

Tsay:1993:SIR

[TB93] Yih-Kuen Tsay and Rajive L. Bagrodia. Some impossibility results in interprocess synchronization. Distributed Computing, 6 (4):221–231, July 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF02242710.

Tapus:2009:DSE

[ŢH09] Cristian Ţăpuş and Jason Hickey. Distributed speculative execution for reliability and fault tolerance: an operational semantics. Distributed Computing, 21(6):433–455, March 2009. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
008-0073-1.

Tarkoma:2006:OCB

[TK06] Sasu Tarkoma and Jaakko Kangasharju. Optimizing contentbased routers: posets and forests. Distributed Computing, 19 (1):62–77, September 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-006-0002-0.

Tay:1995:DER

[TL95] Y. C. Tay and W. Tim Loke. On deadlocks of exclusive ANDrequests for resources. Distributed Computing, 9(2):77–94, October 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s004460050011.

Tseng:2024:IAB

[TLV24] Lewis Tseng, Guanfeng Liang, and Nitin H. Vaidya. Iterative approximate Byzantine consensus in arbitrary directed graphs. Distributed Computing, 37(3):225–246, September 2024. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
024-00468-2.

Torres-Rojas:1999:PCC

[TRA99] Francisco J. Torres-Rojas and Mustaque Ahamad. Plausible clocks: constant size logical clocks for distributed systems. Distributed Computing, 12(4):179–195, September 1999. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050065.

Tsuchiya:2011:VCA

[TS11] Tatsuhiro Tsuchiya and André Schiper. Verification of consensus algorithms using satisfiability solving. Distributed Computing, 23 (5–6):341–358, April 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-010-0123-3.

Tabareau:2019:CFD

[TT19] Nicolas Tabareau and Éric Tanter. Chemical foundations of distributed aspects. Distributed Computing, 32(3):193–216, June 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-018-0334-6.

Tromp:2002:RTP

[TV02] John Tromp and Paul Vitányi. Randomized two-process waitfree test-and-set. Distributed Computing, 15(3):127–135, July 2002. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460200071.

Udding:1986:FMD

[Udd86] Jan Tijmen Udding. A formal model for defining and classifying delay-insensitive circuits and systems. Distributed Computing, 1 (4):197–204, December 1986. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01660032.

Volk:2022:SOB

[VBA22] Matthias Volk, Borzoo Bonakdarpour, and Saba Aflaki. Synthesizing optimal bias in randomized self-stabilization. Distributed Computing, 35(1):37–57, February 2022. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-021-00408-4.

vandeSnepscheut:1987:FME

[vdS87] Jan L. A. van de Snepscheut. Fair mutual exclusion on a graph of processes. Distributed Computing, 2(2):113–115, June 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01667083.

Verhoeff:1988:DIC

[Ver88] Tom Verhoeff. Delay-insensitive codes — an overview. Distributed Computing, 3(1):1–8, March 1988. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/BF01788562.

vanGlabbeek:2023:CCP

[vGGT23] Rob van Glabbeek, Vincent Gramoli, and Pierre Tholoniat. Crosschain payment protocols with success guarantees. Distributed Computing, 36(2):137–157, June 2023. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-023-00446-0.

vanGlabbeek:2016:MVA

[vGHT16] Rob van Glabbeek, Peter Höfner, and Wee Lum Tan. Modelling and verifying the AODV routing protocol. Distributed Computing, 29(4):279–315, August 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-015-0262-7.

Vidyasankar:1990:CRW

[Vid90] K. Vidyasankar. Concurrent reading while writing revisited. Distributed Computing, 4(2):81–85, June 1990. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01786633.

Venkatesan:1994:EAO

[VJ94] S. Venkatesan and Tony T.-Y. Juang. Efficient algorithms for optimistic crash recovery. Distributed Computing, 8(2):105–114, October 1994. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02280832.

Valapil:2020:PSW

[VK20] Vidhya Tekken Valapil and Sandeep S. Kulkarni. Preserving stabilization while practically bounding state space using incorruptible partially synchronized clocks. Distributed Computing, 33(5): 423–443, October 2020. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-019-00365-z.

vanLeeuwen:1987:IUD

[vLT87] J. van Leeuwen and R. B. Tan. An improved upperbound for distributed election in bidirectional rings of processors. Distributed Computing, 2(3):149–160, September 1987. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/BF01782775.

Vogler:1991:FSB

[Vog91] Walter Vogler. Failures semantics based on interval semiwords is a congruence for refinement. Distributed Computing, 4(3): 139–162, September 1991. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/BF01798961.

Wattenhofer:2009:SIP

[Wat09] R. Wattenhofer. Special issue on PODC 2007. Distributed Computing, 21(5):315, February 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-008-0077-x.

Wawrzyniak:2015:LAA

[Waw15] Wojciech Wawrzyniak. A local approximation algorithm for minimum dominating set problem in anonymous planar networks. Distributed Computing, 28(5):321–331, October 2015. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
015-0247-6.

Widder:2012:CPM

[WBB12] Josef Widder, Martin Biely, and Jean-Paul Blanquart. Consensus in the presence of mortal Byzantine faulty processes. Distributed Computing, 24(6):299–321, January 2012. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-011-0147-3.

Welch:1993:MDP

[WL93] Jennifer L. Welch and Nancy A. Lynch. A modular drinking philosophers algorithm. Distributed Computing, 6(4):233–244, July 1993. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/BF02242711.

Widder:2007:BCS

[WS07] Josef Widder and Ulrich Schmid. Booting clock synchronization in partially synchronous systems with hybrid process and link failures. Distributed Computing, 20(2):115–140, August 2007. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic).

URL https://link.springer.com/article/10.1007/s00446-
007-0026-0.

Widder:2009:TMA

[WS09] Josef Widder and Ulrich Schmid. The theta-model: achieving synchrony without clocks. Distributed Computing, 22(1):29–47, April 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-009-0080-x.

Winkler:2019:CRD

[WSS19] Kyrill Winkler, Manfred Schwarz, and Ulrich Schmid. Consensus in rooted dynamic networks with short-lived stability. Distributed Computing, 32(5):443–458, October 2019. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-019-00348-0.

Walter:2004:DRM

[WWA04] Jennifer E. Walter, Jennifer L. Welch, and Nancy M. Amato. Distributed reconfiguration of metamorphic robot chains. Distributed Computing, 17(2):171–189, August 2004. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https:/

/link.springer.com/article/10.1007/s00446-003-0103-y.

Woodruff:2017:WDC

[WZ17] David P. Woodruff and Qin Zhang. When distributed computation is communication expensive. Distributed Computing, 30 (5):309–323, October 2017. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-014-0218-3.

Xu:2008:SAD

[XTB08] Bojian Xu, Srikanta Tirthapura, and Costas Busch. Sketching asynchronous data streams over sliding windows. Distributed Computing, 20(5):359–374, February 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://

link.springer.com/article/10.1007/s00446-007-0048-7.

Yang:1995:FSM

[YA95] Jae-Heon Yang and Jams H. Anderson. A fast, scalable mutual exclusion algorithm. Distributed Computing, 9(1):51–60, March 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01784242.

Yamaguchi:2003:PSR

[YEFH03] Hirozumi Yamaguchi, Khaled El-Fakih, and Teruo Higashino. Protocol synthesis and re-synthesis with optimal allocation of resources based on extended Petri nets. Distributed Computing, 16 (1):21–35, February 2003. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-002-0074-4.

Yu:2009:OIO

[YG09] Haifeng Yu and Phillip B. Gibbons. Optimal inter-object correlation when replicating for availability. Distributed Computing, 21 (5):367–384, February 2009. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-008-0072-2.

Yu:2006:SQS

[Yu06] Haifeng Yu. Signed quorum systems. Distributed Computing, 18 (4):307–323, March 2006. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.springer.

com/article/10.1007/s00446-005-0133-8.

Yu:2011:SHA

[Yu11] Haifeng Yu. Secure and highly-available aggregation queries in large-scale sensor networks via set sampling. Distributed Computing, 23(5–6):373–394, April 2011. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-011-0130-z.

Zhao:1992:SAA

[ZB92] Yi Zhao and Farokh B. Bastani. A self-adjusting algorithm for Byzantine agreement. Distributed Computing, 5(4):219–226, April 1992. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF02277668.

Zielinski:2008:LLA

[Zie08] Piotr Zieliński. Low-latency atomic broadcast in the presence of contention. Distributed Computing, 20(6):435–450, April 2008. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-007-0052-y.

Zielinski:2010:AWF

[Zie10] Piotr Zieliński. Anti-Ω: the weakest failure detector for set agreement. Distributed Computing, 22(5–6):335–348, August 2010. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-010-0101-9.

Zhu:2015:SSU

[ZIM15] Ji Zhu, Stratis Ioannidis, and Laurent Massoulié. Stable and scalable universal swarms. Distributed Computing, 28(6):391–406, December 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-

0452 (electronic). URL https://link.springer.com/article/
10.1007/s00446-014-0228-1.

Zamsky:1995:OTB

[ZIP95] Arkady Zamsky, Amos Israeli, and Shlomit S. Pinter. Optimal time Byzantine agreement for t < n/8 with linearmessages. Distributed Computing, 9(2):95–108, October 1995. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s004460050012.

Zhao:2016:NOC

[ZYC16] Yuda Zhao, Haifeng Yu, and Binbin Chen. Near-optimal communication-time tradeoff in fault-tolerant computation of aggregate functions. Distributed Computing, 29(1):17–38, February 2016. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00446-015-0254-7.